### v1

# Tropical Gradient Descent

Abstract
We propose a gradient descent method for solving optimisation problems arising in settings of tropical geometry—a variant of algebraic geometry that has become increasingly studied in applications such as computational biology, economics, and computer science. Our approach takes advantage of the polyhedral and combinatorial structures arising in tropical geometry to propose a versatile approach for approximating local minima in tropical statistical optimisation problems—a rapidly growing body of work in recent years. Theoretical results establish global solvability for 1-sample problems and a convergence rate of $O(1/\sqrt{k})$. Numerical experiments demonstrate the method’s superior performance over classical descent for tropical optimisation problems which exhibit tropical convexity but not classical convexity. Notably, tropical descent seamlessly integrates into advanced optimisation methods, such as Adam, offering improved overall performance.
Keywords:
tropical quasi-convexity; tropical projective torus; gradient descent; phylogenetics.
1 Introduction
Where algebraic geometry studies geometric properties that can be expressed algebraically of solution sets of systems of multivariate polynomials, tropical geometry restricts to the case of polynomials defined by linearising operations, where the sum of two elements is their maximum and the product of two elements is their sum. Evaluating polynomials with these operations results in piecewise linear functions. Tropical geometry is a relatively young field of pure mathematics established in the 1990s (with roots dating back further) and has recently become an area of active interest in computational and applied mathematics for its relevance to the problem of dynamic programming in computer science (Maclagan & Sturmfels, (2021)); in mechanism design for two-player games in game theory (Lin & Tran, (2019)); and statistical analyses of phylogenetic trees (Monod et al., (2022)). Given the connection of tropical geometry to applied sciences, developing and implementing computational techniques for tropical geometric settings is becoming increasingly important.
A central task to a vast majority of computational tasks and training machine learning models entails solving an optimisation problem. In this paper, we are motivated by the optimisation problems in statistics in tropical geometric settings. We focus on the classical gradient descent approach for optimisation and adapt it to the setting of tropical geometry. We thus provide a unifying gradient descent framework to solve the optimisation problems entailed in the computation of a wide range of statistical quantities and methods in tropical spaces which are an area of active interest in applied tropical geometry.
The remainder of this paper is organised as follows. The following section contains the necessary foundations of tropical geometry to present, an introduction to tropical location problems functions (Comăneci, (2023)), and the array of statistical optimisation tasks studied in this paper. In Section 3 we formulate steepest descent with respect to the tropical norm and present the theoretical guarantees for tropical descent, most notably its convergence for a wide class of tropically quasi-convex functions and a convergence rate of $O(1/\sqrt{k})$. In Section 4, we perform a comprehensive numerical study of our proposed tropical gradient descent method and its performance in the statistical optimisation problems of interest. We conclude with a discussion of possible future work on tropical gradient methods in Section 5.
2 Background and Preliminaries
In this section, we overview the background on tropical geometry and present the statistical concepts and methodologies which are defined by an optimisation problem that we will study in the tropical setting.
2.1 Tropical Geometry
Here, we review the algebraic and geometric structure of the tropical projective torus, the state space for phylogenetic data and the domain of our tropical statistical optimisation problems.
Definition 1 (Tropical Algebra).
The tropical algebra is the semiring $\overline{\mathbb{R}}=\mathbb{R}\cup\{-\infty\}$ with the addition and multiplication operators—tropical addition and tropical multiplication, respectively—given by
$a\boxplus b=\max\{a,b\},\quad a\odot b=a+b.$
The additive identity is $\infty$- ∞ and the multiplicative identity is 00. Tropical subtraction is not defined; tropical division is given by classical subtraction.
The tropical algebra defined above is sometimes referred to as the max-plus algebra; we can define the min-plus algebra analogously, taking $\min\{a,b\}=:a\oplus b$ as the additive operation on ${}\mathbb{R}\cup\{\infty\}$. These are algebraically equivalent through negation. Unless specified, we use the max-plus convention as this is better suited for phylogenetic statistics on the tropical projective torus (Speyer & Sturmfels,, 2004).
Definition 2 (Tropical Projective Torus).
The $n$-dimensional tropical projective torus is a quotient space constructed by endowing $\mathbb{R}^{n+1}$ with the equivalence relation
$\mathbf{x}\sim\mathbf{y}\Leftrightarrow\exists\,a:\>\mathbf{x}=a\odot\mathbf{y};$ (1)
it is denoted by $\mathbb{R}^{n+1}/\mathbb{R}\mathbf{1}$. The generalised Hilbert projective metric, also referred to as the tropical metric, is given by
$d_{{\rm tr}}(\mathbf{x},\mathbf{y})=\max_{i}\{x_{i}-y_{i}\}-\min_{i}\{x_{i}-y_{i}\}=\max_{i,j}\{x_{i}-y_{i}-x_{j}+y_{j}\}.$
This metric is induced by the tropical norm, which is given by
$\|\mathbf{x}\|_{{\rm tr}}=\max_{i}x_{i}-\min_{i}x_{i}.$
The tropical projective torus is the ambient space containing the tropical Grassmannian, which is equivalent to the space of phylogenetic trees (Speyer & Sturmfels, (2004)). This has sparked the study of tropical data science, formalising statistical techniques which respect the tropical geometry of the state space (Monod et al., (2018); Yoshida, (2020)).
The metric above is widely accepted and preferred for the development of geometric statistical tools on the tropical projective torus (Yoshida, (2020)). However, recent work has shown theoretical advantages to the use of an asymmetric metric (Comăneci & Joswig, (2023)).
Definition 3 (Tropical Asymmetric Distance).
The tropical asymmetric distance on $\mathbb{R}^{n}/\mathbb{R}\mathbf{1}$ is given by
$d_{\triangle_{\min}}(\mathbf{a},\mathbf{b})\coloneqq\sum_{i}(b_{i}-a_{i})-n\min_{j}(b_{j}-a_{j}),$
$d_{\triangle_{\max}}(\mathbf{a},\mathbf{b})\coloneqq n\max_{j}(b_{j}-a_{j})-\sum_{i}(b_{i}-a_{i}).$
Remark 4.
The asymmetric distances act as an $\ell_{1}$ norm on the tropical projective torus, which the symmetric metric $d_{\mathrm{tr}}$ acts as an $\ell_{\infty}$ norm. We note that $nd_{{\rm tr}}=d_{\triangle_{\min}}+d_{\triangle_{\max}}$. Figure 1 shows the balls for each tropical metric of interest.
$\mathbf{0}$$\mathbf{0}$$\mathbf{0}$
Figure 1: The $d_{\mathrm{tr}}$, $d_{\triangle_{\min}}$ and $d_{\triangle_{\max}}$ tropical unit balls in $\mathbb{R}^{3}/\mathbb{R}\mathbf{1}\cong\mathcal{H}=\{\sum x_{i}=0\}$.
We now define the primary geometric objects of interest in the tropical projective torus; hyperplanes, lines, and convex hulls. We use the $\max$-convention for these definitions, though we note each has a $\min$-convention equivalent.
Definition 5 (Tropical Hyperplanes).
The tropical hyperplane $\mathcal{H}_{\mathbf{a}}$ defined by the linear form $a_{1}\odot x_{1}\boxplus\dots\boxplus a_{n}\odot x_{n}$ is given by
$\mathcal{H}_{\mathbf{a}}=\{\mathbf{x}:\exists i\neq j\text{ s.t. }a_{i}+x_{i}=a_{j}+x_{j}=\max\{a_{k}+x_{k}\}\}.$
Definition 6 (Tropical Line Segments).
For any two points $\mathbf{a},\mathbf{b}\in\mathbb{R}^{n}/\mathbb{R}\mathbf{1}$, the tropical line segment between $\mathbf{a}$ and $\mathbf{b}$ is the set
$\gamma_{\mathbf{a}\mathbf{b}}=\{\alpha\odot\mathbf{a}\boxplus\beta\odot\mathbf{b}\mid\alpha,\beta\in\mathbb{R}\}$
with tropical addition taken coordinate-wise.
Tropical line segments define a unique geodesic path between any two points, but general geodesics are not uniquely defined on the tropical projective torus.
$A$
$B$$A$
$A$$B$$C$
Figure 2: A tropical hyperplane, tropical line segment and tropical convex hull in $\mathbb{R}^{3}/\mathbb{R}\mathbf{1}\cong\mathcal{H}=\{\mathbf{x}:\sum x_{i}=0\}$. The dashed lines show coordinate directions.
Definition 7 (Tropical Convex Hulls (Develin & Sturmfels,, 2004)).
A set $S$ is $\max$-tropically convex if it contains $\alpha\odot\mathbf{x}\boxplus\beta\odot\mathbf{y}$ for all $\mathbf{x},\mathbf{y}\in S$ and $\alpha,\beta\in\mathbb{R}$. For a finite subset $X=\{\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\}\subset\mathbb{R}^{n}/\mathbb{R}\mathbf{1}$, the max-plus tropical convex hull of $X$ is the set of all max-tropical linear combinations of points in $X$,
$\operatorname{tconv}_{\max}(X)=\{\alpha_{1}\odot\mathbf{x}_{1}\boxplus\alpha_{2}\odot\mathbf{x}_{2}\boxplus\cdots\boxplus\alpha_{n}\odot\mathbf{x}_{n}\mid\alpha_{1},\ldots,\alpha_{n}\in\mathbb{R}\}.$
Similarly, the min-tropical convex hull of $X$, $\operatorname{tconv}_{\min}(X)$ is the set of all min-tropical linear combinations of points in $X$.
A study of tropically convex functions requires an understanding of both max-tropical and min-tropical convex hulls, as we will see in our following consideration of tropical location problems.
2.2 Tropical Location Problems
Tropical location problems are a new avenue of research, motivated by the optimisation of statistical loss functions; they involve the minimisation over some dataset of some loss function which heuristically increases with distance from some kernel. Here we outline the relevant definitions of tropical location problems and their theoretical behaviour as presented by Comăneci, (2023). We introduce the problem of tropical linear regression (Akian et al., (2023)) as motivation.
2.2.1 Motivating Example
The problem of tropical linear regression looks to find the best-fit tropical hyperplane which minimises the maximal distance to a set of data points. While this is an inherently statistical optimisation problem, Akian et al., (2023) prove it to be polynomial-time equivalent to solving mean payoff games.
A tropical hyperplane is uniquely defined by a cone point $\mathbf{t}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$, allowing us to formulate the best-fit tropical hyperplane as an optimisation problem over $\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$.
Definition 8 (Tropical Linear Regression (Akian et al., (2023))).
Let $X=\{\mathbf{x}_{1},\dots,\mathbf{x}_{K}\}\subset\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ be our dataset. The tropical linear regression problem finds the vertex $\mathbf{t}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ of the best-fit hyperplane:
$\min_{\mathbf{t}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}}f(\mathbf{t})=\max_{k\leq K}d_{\mathrm{tr}}(\mathbf{x},\mathcal{H}_{\mathbf{t}})=\max_{k\leq K}(\mathbf{x}_{k}-\mathbf{t})_{i}-(\mathbf{x}_{k}-\mathbf{t})_{j}$
where $i,j$ are such that $\forall\,\ell\neq i,j:(\mathbf{x}_{k}-\mathbf{t})_{i}\geq(\mathbf{x}_{k}-\mathbf{t})_{j}\geq(\mathbf{x}_{k}-\mathbf{t})_{\ell}$.
In Figure 3, we see a heatmap of this objective function over the tropical projective plane. We see immediately that the objective is not convex. The linear regression objective has sparse gradient almost everywhere, which produces valleys. These properties are considered unfavourable extreme cases for gradient methods, but are particularly common in tropical optimisation problems. While the tropical linear regression problem is not convex, it is quasi-convex along tropical line segments; that is, its sub-level sets are tropically convex.
Figure 3: A heat map of a 3-sample linear regression objective on $\mathbb{R}^{3}/\mathbb{R}\mathbf{1}\cong\mathcal{H}=\{\mathbf{x}:\sum x_{i}=0\}$.
2.2.2 $\triangle$-Star Quasi-Convexity
In formalising tropical location problems, Comăneci, (2023) first considers functions which increase along tropical geodesics from a kernel $\mathbf{v}$; that is, for any point $\mathbf{t}$ on a geodesic from $\mathbf{v}$ to $\mathbf{w}$, we have that $f(\mathbf{t})\leq f(\mathbf{w})$. This property is equivalent to star-convexity of sublevel sets. To formally state this, we define (oriented) geodesic segments and $\triangle_{\min}$-star convexity of sets.
Definition 9 (Oriented Geodesic Segment).
The (oriented) geodesic segment between $\mathbf{a},\mathbf{b}\in\mathbb{R}^{n}/\mathbb{R}\mathbf{1}$, is given by $[\mathbf{a},\mathbf{b}]_{\triangle_{\max}}\coloneqq\{\mathbf{x}\in\mathbb{R}^{n}/\mathbb{R}\mathbf{1}:d_{\triangle_{\max}}(\mathbf{a},\mathbf{x})+d_{\triangle_{\max}}(\mathbf{x},\mathbf{b})=d_{\triangle_{\max}}(\mathbf{a},\mathbf{b})\}$.
Definition 10 ($\triangle_{\min}$-Star Convex Set).
A set $K\subseteq\mathbb{R}^{n}/\mathbb{R}\mathbf{1}$ is a $\triangle_{\max}$-star-convex set with kernel $\mathbf{v}$ if, for every $\mathbf{w}\in K$, we have $[\mathbf{v},\mathbf{w}]_{\triangle_{\max}}\subseteq K$.
Using the above definitions, we can characterise a loss function which is increasing along tropical geodesics as a function with $\triangle_{\max}$-star-convex sublevel sets. The following definition gives a closed form expression for such functions, while Theorem 12 proves their equivalence.
Definition 11 (Comăneci, (2023)).
A function $f:\mathbb{R}^{n}/\mathbb{R}\mathbf{1}\rightarrow\mathbb{R}$ is $\triangle_{\max}$-star-quasi-convex with kernel $\mathbf{v}$ if $f(\mathbf{x})=\hat{\gamma}(\mathbf{x}-\mathbf{v})$ for some increasing $\gamma:\mathbb{R}^{n}_{\geq 0}\rightarrow\mathbb{R}$, where $\hat{\gamma}(\mathbf{x})=\gamma((\max_{i}x_{i})\mathbf{1}-\mathbf{x})$.
Theorem 12 (Theorem 17, Comăneci, (2023)).
A continuous function $f:\mathbb{R}^{n}/\mathbb{R}\mathbf{1}\rightarrow\mathbb{R}$ is $\triangle_{\max}$-star-quasi-convex with kernel $\mathbf{v}$ if and only if all of its non-empty sub-level sets are $\triangle_{\max}$-star convex with kernel $\mathbf{v}$.
These $\triangle_{\max}$-star-convex functions act as a loss function with respect to a single datapoint. For example, the tropical metric $d_{\mathrm{tr}}$ is both $\triangle_{\min}$ and $\triangle_{\max}$-star-quasi-convex. The distance to a hyperplane $d_{\mathrm{tr}}(\mathbf{x},\mathcal{H}_{\mathbf{t}})$ is $\triangle_{\max}$-star-quasi-convex in $\mathbf{x}$, but $\triangle_{\min}$-star-quasi-convex in $\mathbf{t}$.
As a final remark, we note that $\triangle_{\max}$-star-quasi-convex functions are a special case of tropically quasi-convex functions by the following lemma. We limit our theoretical considerations to $\triangle_{\max}$-star-quasi-convex functions to utilise their closed form expression given in Definition 11, however numerical experiments demonstrate strong performance for more general tropical quasi-convex functions.
Lemma 13 (Comăneci, (2023)).
Any $\triangle_{\max}$-star-convex set is $\max$-tropically convex. Hence, any $\triangle_{\max}$-star-quasi-convex function is tropically quasiconvex in that it has tropically convex sublevel sets.
2.2.3 Tropical Location Problems
The $\triangle$-star-convex functions defined above are generally a measure of closeness to the kernel $\mathbf{v}$, but in minimising a statistical loss function we look at closeness to a sample of points. A location problem (Laporte et al., (2019)) measures the closeness to a sample rather than a single kernel.
Definition 14 (Max(/min)-tropical location problem (Comăneci, (2023))).
Consider a dataset $X=\{\mathbf{x}_{1},\dots,\mathbf{x}_{K}\}\subset\mathbb{R}^{n}/\mathbb{R}\mathbf{1}$. For $k\leq K$, let $h_{k}$ be a $\triangle_{\max}$-star-quasi-convex function with kernel $\mathbf{x}_{k}$, and let $g:\mathbb{R}^{m}\rightarrow\mathbb{R}$ be an increasing function. A max-tropical location problem is the minimisation of an objective function $f:\mathbb{R}^{n}/\mathbb{R}\mathbf{1}\rightarrow\mathbb{R}$ given by $h=g(h_{1},\dots,h_{m})$. A min-tropical location problem is defined similarly, where the $h_{k}$ are $\triangle_{\min}$-star-quasi-convex.
The main result by Comăneci, (2023) is the following; tropical location problems contain a minimum in the tropical convex hull of its dataset $X$.
Theorem 15 (Theorem 20, Comăneci, (2023)).
Let $f$ be the objective function of a max-tropical location problem. Then there is a minimum of $f$ belonging to $\operatorname{tconv}_{\min}(X)$. Conversely, min-tropical location problems have minima belonging to $\operatorname{tconv}_{\max}(X)$.
We can now revisit our motivating example (Definition 8), showing that tropical linear regression is not just a min-tropical location problem, but also demonstrates tropical quasi-convexity.
Proposition 16.
The tropical linear regression problem is a tropically quasi-convex $\min$-tropical location problem.
Proof.
$f(\mathbf{t})$ $=\max_{k\leq K}\min_{j}[\max_{i}[x_{ki}-t_{i}]-x_{kj}+t_{j}]$
$=\max_{k\leq K}\min_{j}[t_{j}-x_{kj}-\min_{i}[t_{i}-x_{ki}]]$
These are min-tropical location problems with:
$g(\mathbf{h})$ $=\max_{k\leq K}h_{k},$
$h_{k}(\mathbf{t})$ $=\hat{\gamma}(\mathbf{t}-\mathbf{x}_{k}),$
$\gamma(\mathbf{t})$ $=\min_{j}t_{i}.$
This $g$ and $\gamma$ are increasing and each $h_{k}$ has kernel $x_{k}$, satisfying the conditions for a min-tropical location problem.
As the $h_{k}$ are $\triangle_{\min}$-star-convex, their sublevel sets are min-tropically convex. This is preserved when taking a maximum over samples, so $f$ also has $\min$-tropically convex sublevel sets. ∎
As mentioned in Lemma 13, $\triangle_{\min}$-star-quasi-convexity implies tropical quasi-convexity. However, the tropical linear regression problem is a case in which the reverse implication does not hold; for any sample of size $\geq 1$, the linear regression loss function is not $\triangle_{\min}$-star-quasi-convex.
2.3 Tropical Data Science
We now review various tropical statistical optimisation problems which have been introduced in the literature in recent years, and which are used to test our tropical descent methodology. We note the varying degrees of convexity demonstrated by each of them.
These statistical concepts and methods showcase the breadth of applicability of our proposed methodology: optimisation is a central task in computing important statistical quantities, tasks in both descriptive and inferential statistics, and distances between probability measures. Adapting these statistical tasks to the setting of tropical geometry is not straightforward and it is not immediate that a single tropical optimisation method can be used to find solutions to the optimisation problems.
2.3.1 Centrality Statistics
Our first problems of interest, Fermat–Weber Points and Fréchet means, are centrality statistics for data on a general metric space; Fermat–Weber points (Wesolowsky, (1993)) act as a generalised median, while Fréchet means (Fréchet, (1948)) are a generalisation of the mean.
Definition 17 (Tropical Fermat–Weber Points (Lin & Yoshida, (2018))).
Let $X=\{\mathbf{x}_{1},\dots,\mathbf{x}_{K}\}\subset\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ be our dataset. A solution to the following optimisation problem is a tropical Fermat–Weber point:
$\min_{\mathbf{t}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}}f(\mathbf{t})=\frac{1}{K}\sum_{k\leq K}d_{{\rm tr}}(\mathbf{x}_{k},\mathbf{t}).$
Definition 18 (Tropical Fréchet means (Jazlan et al., (2024+); Monod et al., (2022))).
Let $X=\{\mathbf{x}_{1},\dots,\mathbf{x}_{K}\}\subset\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ be our dataset. A solution to the following optimisation problem is a tropical Fréchet mean:
$\min_{\mathbf{t}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}}f(\mathbf{t})=\left[\frac{1}{K}\sum_{k\leq K}d_{{\rm tr}}(\mathbf{x}_{k},\mathbf{t})^{2}\right]^{1/2}.$
The Fermat–Weber and Fréchet mean problems are both max-tropical and min-tropical location problems, as the tropical metric $d_{\mathrm{tr}}$ is both $\triangle_{\min}$ and $\triangle_{\max}$-star-quasiconvex. In fact, they are both classically convex and hence act as a benchmark against which we will test our gradient methods.
2.3.2 Wasserstein Projections
The next statistical optimisation problem was motivated by Cai & Lim, (2022), and looks to identify a tropical projection which minimises the Wasserstein distance between samples in different tropical projective torii.
Definition 19 (Tropical Wasserstein Projections (Talbut et al., (2023))).
Let $X=\{\mathbf{x}_{1},\dots,\mathbf{x}_{K}\}\subset\mathbb{R}^{N}/\mathbb{R}\mathbf{1},Y=\{\mathbf{y}_{1},\dots,\mathbf{y}_{K}\}\subset\mathbb{R}^{M}/\mathbb{R}\mathbf{1}$ be our datasets. Let $J_{1},\dots,J_{M}$ be a non-empty partition of $[N]$. Then the tropical $p$-Wasserstein projection problem is the minimisation of the following objective function:
$\min_{\mathbf{t}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}}f_{p}(\mathbf{t})=\left(\frac{1}{K}\sum_{k\leq K}\|[\max_{i\in J_{j}}(\mathbf{x}_{k}+\mathbf{t})_{i}]_{j}-\mathbf{y}_{k}\|_{\mathrm{tr}}^{p}\right)^{1/p}$
The tropical $\infty$-Wasserstein projection problem is the minimisation of:
$\min_{\mathbf{t}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}}f_{\infty}(\mathbf{t})=\max_{k\leq K}\|[\max_{i\in J_{j}}(\mathbf{x}_{k}+\mathbf{t})_{i}]_{j}-\mathbf{y}_{k}\|_{\mathrm{tr}}$
In contrast to the work by Lee et al., (2021) which looks to solve the optimal transport problem for intrinsic measures on $\mathbb{R}^{n}/\mathbb{R}\mathbf{1}$, this $p$-Wasserstein projection problem is computing an optimal projection for empirical samples.
Proposition 20.
The tropical $p,\infty$-Wasserstein projections are max-tropical location problems, and the $\infty$-Wasserstein projection has $\max$-tropically convex sublevel sets.
Proof.
Let $\mathbf{z}_{k}=(x_{ki}-y_{kj})_{i\leq N}$ where $j$ is such that $i\in J_{j}$. We then have that
$\|[\max_{i\in J_{j}}(\mathbf{x}_{k}+\mathbf{t})_{i}]_{j}-\mathbf{y}_{k}\|_{\mathrm{tr}}$ $=\max_{l}[t_{l}+z_{kl}]+\max_{j}[\min_{i\in J_{j}}-t_{i}-z_{ki}]$
$=\max_{j}[\min_{i\in J_{j}}[\max_{l}[t_{l}+z_{kl}]-t_{i}-z_{ki}]]$
The $p,\infty$-Wasserstein projections can then be written as
$f_{p}(\mathbf{t})$ $=\left(\frac{1}{K}\sum_{k\leq K}\left(\max_{j}[\min_{i\in J_{j}}[\max_{l}[t_{l}+z_{kl}]-t_{i}-z_{ki}]]\right)^{p}\right)^{1/p}$
$f_{\infty}(\mathbf{t})$ $=\max_{k\leq K}\max_{j}[\min_{i\in J_{j}}[\max_{l}[t_{l}+z_{kl}]-t_{i}-z_{ki}]]$
These are max-tropical location problems with:
$g_{p}(\mathbf{h})$ $=\left(\frac{1}{K}\sum_{k\leq K}h_{k}^{p}\right)^{1/p},$
$g_{\infty}(\mathbf{h})$ $=\max_{k\leq K}h_{k},$
$h_{k}(\mathbf{t})$ $=\hat{\gamma}(\mathbf{t}+\mathbf{z}_{k}),$
$\gamma(\mathbf{t})$ $=\max_{j}\min_{i\in J_{j}}t_{i}.$
The $g_{p},g_{\infty}$ and $\gamma$ are increasing and each $h_{k}$ has kernel $-z_{k}$, satisfying the conditions for a max-tropical location problem.
As the $h_{k}$ are $\triangle_{\max}$-star-convex, their sublevel sets are max-tropically convex. This is preserved when taking a maximum over samples, so $f_{\infty}$ also has $\max$-tropically convex sublevel sets. ∎
Figure 4: A heat map of a 3-sample $\infty$-Wasserstein projection objective on $\mathbb{R}^{3}/\mathbb{R}\mathbf{1}\cong\mathcal{H}=\{\mathbf{x}:\sum x_{i}=0\}$.
3 Steepest Descent with the Tropical Norm
This section comprises the theoretical contributions of this paper; we outline our proposed method of tropical descent, and the theoretical convergence guarantees for tropical descent when applied to tropical location problems.
We treat optimisation over $\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ as optimisation over $\mathbb{R}^{N}$ with the knowledge that $f$ is constant along $\mathbf{1}$-rays. We assume that for all $\mathbf{t}$, we can compute some gradient $\nabla f(\mathbf{t})$ which is contained in the generalised Clarke derivative of $\mathbf{t}$.
Definition 21 (Generalised Clarke Derivative (2.13, I.I of Kiwiel, (2006))).
We define the set $M_{f}(\mathbf{t})$ by
$M_{f}(\mathbf{t})=\{\mathbf{z}\in\mathbb{R}^{N}:\nabla f(\mathbf{y}_{l})\rightarrow\mathbf{z},\text{ where }\mathbf{y}_{l}\rightarrow\mathbf{t},f\text{ differentiable at each }\mathbf{y}_{l}\}.$
Then the generalised Clarke derivative is given by
$\partial f(\mathbf{t})=\operatorname{conv}M_{f}(\mathbf{t}).$
3.1 Proposed Method
In defining tropical descent, we refer back to the motivation behind classical gradient descent; that is, the steepest descent direction.
Definition 22 (Steepest Descent Direction (Boyd & Vandenberghe, (2004))).
[…]
$\mathbf{d}=\operatorname{argmax}\{\nabla f(\mathbf{t})^{\top}\mathbf{v}:\|\mathbf{v}\|=1\}.$
With respect to the Euclidean norm, the normalised gradient is a steepest descent direction. When we use a the tropical norm, the steepest descent directions are characterised by the following lemma.
Lemma 23.
Suppose $\nabla f\neq 0$. Then direction $\mathbf{v}$ is a normalised tropical steepest descent direction iff
$v_{i}=1$ $\Leftrightarrow\nabla f_{i}<0$
$v_{i}=0$ $\Leftrightarrow\nabla f_{i}>0$
Proof.
WLOG, we assume $0\leq v_{i}\leq 1$. Then
$\nabla f^{\top}\mathbf{v}\leq\sum_{\nabla f_{i}>0}\nabla f_{i}$
This is achieved if and only if $v_{i}=1$ for all $i$ such that $\nabla f_{i}>0$, and $v_{i}=0$ for all $i$ such that $\nabla f_{i}<0$. We note that any such $v_{i}$ satisfies $\|\mathbf{v}\|_{{\rm tr}}=1$ as $\sum\nabla f_{i}=0,\nabla f_{i}\neq 0$, and hence $\nabla f_{i}$ has positive and negative coordinates. ∎
This lemma uniquely defines $\mathbf{d}$ if and only if, for every coordinate $i$, $\nabla f_{i}\neq 0$. To uniquely define a descent direction for any $\mathbf{x}$, we take the direction $\mathbf{d}$ given by
$d_{i}=\begin{cases}1&\text{ if }\nabla f_{i}>0,\\ 0&\text{ otherwise.}\end{cases}$
We refer to this as a positive descent direction. In contrast,
$d_{i}=\begin{cases}-1&\text{ if }\nabla f_{i}<0,\\ 0&\text{ otherwise.}\end{cases}$
defines a negative descent direction.
Throughout the rest of this paper, we use the positive descent direction as convention, and will specify when the negative descent direction is used instead. Tropical descent therefore refers to the method outlined by Algorithm 1, for some specified step size rule, while we refer to classical gradient descent as classical descent.
Algorithm 1 Tropical Descent
for $n<$ max_steps do
     $\mathbf{g}_{n}\leftarrow\nabla f(\mathbf{t}_{n-1})$;
     $\mathbf{d}_{n}\leftarrow[\mathbf{g}_{n}>0]_{i\leq N}$;
     $\mathbf{t}_{n}\leftarrow\mathbf{t}_{n-1}-a_{n}\mathbf{d}_{n}$;
3.2 Convergence Guarantees
In this section we prove the theoretical guarantees of tropical descent for max-tropical location problems. We show that tropical descent must converge to the tropical convex hull of the data, and hence will necessarily find a minimum of a $\triangle_{\max}$-star-convex problem. Throughout this section, we assume the step sizes $a_{n}$ satisfy $a_{n}\rightarrow 0,\sum a_{n}=\infty$.
We first highlight the key difference in stability between classical descent and tropical descent; if some component of the gradient is locally zero, tropical descent cannot converge in that neighbourhood.
Proposition 24.
Assume that for all $n$, $\nabla f(\mathbf{t}_{n})\neq\mathbf{0}$. Suppose in some bounded open $U$, $\nabla f_{i}$ is identically 0. Then no point in $\operatorname{int}U$ is a convergence point for $\mathbf{t}_{n}$; in particular, if $\mathbf{t}_{m}\in U$ we will eventually leave $U$.
Proof of Proposition 24.
Assume otherwise. Then for all $n\geq m$, we have $\nabla f(\mathbf{t}_{n})_{i}=0$, so $t_{ni}$ is constant. At each step, we decrease at least one coordinate by $a_{n}$, so as $n\rightarrow\infty$:
$\sum_{j}t_{nj}\leq\sum_{j}t_{mj}-\sum_{n\geq m}a_{n}\rightarrow-\infty.$
Therefore $\|\mathbf{t}_{n}\|_{{\rm tr}}$ will become arbitrarily large, as $\min_{j}t_{nj}\rightarrow-\infty$ while $\max_{j}t_{nj}\geq t_{ni}$. Then we must have left $U$; contradiction. ∎
As a result, local minima such as the valleys in figures 4 and 3 can be stable with respect to classical descent but unstable with respect to tropical descent. We next prove an opposing result; tropical gradient descent must converge to the tropical convex hull of the data points. This result stems from the following observation on the gradient of $\triangle_{\max}$-star-quasi-convex functions.
Remark 25.
We note that for a $\triangle_{\max}$-star-quasi-convex function with kernel $\mathbf{v}$, we have that
$\nabla f(\mathbf{x})_{j}>0\Rightarrow j\in\operatorname{argmax}_{i}(x_{i}-v_{i}).$
Theorem 26.
Suppose $f$ is a max-tropical location problem with respect to the dataset $X=\{\mathbf{x}_{1},\dots,\mathbf{x}_{K}\}$. Let $(\mathbf{t}_{n})_{n\geq 1}$ be a sequence of points defined by tropical descent, such that for all $n$: $\nabla f(\mathbf{t}_{n})\neq 0$. Let $V=\operatorname{tconv}_{\min}(\mathbf{x}_{1},\dots,\mathbf{x}_{K})$, and define
$\Delta_{i}(\mathbf{t})=\sum_{k\in[K]}\sum_{j\in[N]}[t_{j}-x_{kj}-t_{i}+x_{ki}]^{+}.$
For any $\epsilon>0$, let $M_{1}$ be such that for all $m\geq M_{1}$, $a_{m}\leq\epsilon/N$. Let $M_{2}$ be such that for all $m\geq M_{2}$ $s_{m}\geq\max_{i}\Delta_{i}(\mathbf{t}_{M_{1}})+s_{M_{1}}-K\epsilon$. Then almost surely, for all $m\geq M_{2}$ we have:
$d_{{\rm tr}}(\mathbf{t}_{n},V)\leq\epsilon.$
Proof.
For each $i\leq N$, we define the functions
$\delta_{i}(\mathbf{t})=\min_{k\in[K]}[\sum_{j}[t_{j}-v_{kj}-t_{i}+v_{ki}]^{+}].$
Claim 27.
$d_{{\rm tr}}(\mathbf{t}_{m},V)\leq\max_{j}\delta_{j}(\mathbf{t}_{m})$.
Proof.
As discussed in Maclagan & Sturmfels, (2015), we have the projection map onto $V$ given by:
$\pi_{V}(\mathbf{t})=\lambda_{1}\odot\mathbf{x}_{1}\oplus\dots\oplus\lambda_{K}\odot\mathbf{x}_{K},\qquad\text{ where }\lambda_{k}=\max_{i}(t_{i}-x_{ki})$
So the distance to $V$ is given by $d(\mathbf{t},\pi_{V}(\mathbf{t}))$, and we have:
$d(\mathbf{t},V)$ $=d(\mathbf{t},\pi_{V}(t))$
$=\|(\min_{k}(\max_{i}(t_{i}-x_{ki})+x_{kj}-t_{j}))_{j}\|$
$\leq\max_{j}(\min_{k}(\max_{i}(t_{i}-x_{ki})+x_{kj}-t_{j}))$
$\leq\max_{j}\delta_{j}(\mathbf{t})$
Claim 28.
If $\delta_{i}(\mathbf{t}_{m})=0$ and $m\geq M_{1}$, then $\delta_{i}(\mathbf{t}_{m+1})<\epsilon$.
Proof.
For all $k\in[K],j\in[N]$, we have that $t_{j}-v_{kj}-t_{i}+v_{ki}$ is $1$-Lipschitz in $\mathbf{t}$. Therefore $\delta_{i}$ is $N$- Lipschitz in $\mathbf{t}$, and if $\delta_{i}(\mathbf{t}_{m})=0,\delta_{i}(\mathbf{t}_{m+1})\leq Na_{m}<\epsilon$. ∎
Claim 29.
If $\delta_{i}(\mathbf{t}_{m})>0$ then $\delta_{i}(\mathbf{t}_{m+1})\leq\delta_{i}(\mathbf{t}_{m})$.
Proof.
Suppose $\delta_{i}(\mathbf{t}_{m})>0$. Then $\forall k$, $\max_{j}(t_{j}-v_{kj})>t_{i}-v_{ki}$. By remark 25, $\nabla f_{i}\leq 0$ and $d_{mi}=0$. Therefore $\delta_{i}(\mathbf{t}_{m+1})\leq\delta_{i}(\mathbf{t}_{m})$. ∎
Claim 30.
Let $M$ be such that $s_{M}\geq\Delta_{i}(\mathbf{t}_{M_{1}})+s_{M_{1}}-K\epsilon$. Then for some $M_{1}\leq m\leq M$, $\delta_{i}(\mathbf{t}_{m})<\epsilon$.
Proof.
We note that $K\delta_{i}(\mathbf{t})\leq\Delta_{i}(\mathbf{t})$; hence if $\Delta_{i}(\mathbf{t}_{M_{1}})<K\epsilon$ then we are done.
We now assume $\Delta_{i}(\mathbf{t}_{M_{1}})\geq K\epsilon$ and prove the result by contradiction; suppose $\delta_{i}(\mathbf{t}_{m})>\epsilon$ for all $M_{1}\leq m\leq M$. Therefore at each step we have $\nabla f_{i}\leq 0$, $d_{mi}=0$ and $t_{mi}$ is fixed for $n\leq M$.
As $\nabla f\neq\mathbf{0}$, there is some $j$ such that $\nabla f_{j}>0$, and let $k$ be such that $j=\operatorname{argmax}_{l}t_{ml}-v_{kl}$. As $\delta_{i}(\mathbf{t}_{m})>\epsilon$:
$t_{j}-v_{kj}-t_{i}+v_{ki}>\epsilon/N\geq a_{m}.$
Hence this term in the summation of $\Delta_{i}$ decreases by $a_{m}$, while other terms are non-increasing. We conclude that $\Delta_{i}(\mathbf{t}_{M})\leq\Delta_{i}(\mathbf{t}_{M_{1}})-s_{M}+s_{M_{1}}\leq K\epsilon$. Contradiction. ∎
Finally, from claims 28 to 30, we conclude that $\forall m\geq M_{2},i\leq N$, we have $\delta_{i}(\mathbf{t}_{m})<\epsilon$, and by claim 27, we have $d_{{\rm tr}}(\mathbf{t}_{m},V)\leq\epsilon$. ∎
As proven in Comăneci, (2023), tropical location problems have minima in the tropical convex hull of the data. This result says that according to tropical gradient descent, only such minima can be stable. As a corollary, we can achieve the same result for min-tropical location problems by using negative tropical descent directions.
Corollary 31.
Suppose $f$ is a min-tropical location problem with respect to the dataset $X=\{\mathbf{x}_{1},\dots,\mathbf{x}_{K}\}$. Let $(\mathbf{t}_{n})_{n\geq 1}$ be a sequence of points defined via negative tropical descent directions, such that for all $n$: $\nabla f(\mathbf{t}_{n})\neq 0$. Let $V=\operatorname{tconv}_{\max}(\mathbf{x}_{1},\dots,\mathbf{x}_{K})$, and define
$\Delta^{\prime}_{j}(\mathbf{t})=\sum_{k\in[K]}\sum_{i\in[N]}[t_{j}-x_{kj}-t_{i}+x_{ki}]^{+}.$
For any initial $\mathbf{t}_{0}$ and $\epsilon>0$, let $M_{1}$ be such that for all $m\geq M_{1}$, $a_{m}\leq\epsilon/N$. Let $M_{2}$ be such that for all $m\geq M_{2}$ $s_{m}\geq\max_{i}\Delta^{\prime}_{j}(\mathbf{t}_{M_{1}})+s_{M_{1}}-K\epsilon$. Then almost surely, for all $m\geq M_{2}$ we have:
$d_{{\rm tr}}(\mathbf{t}_{n},V)\leq\epsilon.$
Despite following from the previous results, the following corollary may be considered the most powerful result of this paper; tropical descent will necessarily minimise any max-tropically quasi-convex function.
Corollary 32.
Let $f(\mathbf{t})$ be a max-tropically quasiconvex function, and let $(\mathbf{t}_{n})_{n\geq 1}$ be a sequence of points defined by tropical descent such that for all $n$: $\nabla f(\mathbf{t}_{n})\neq 0$. Then there is some global minimum $\mathbf{t}^{*}$ such that for all $m\leq M_{2}$, $d_{{\rm tr}}(\mathbf{t}_{m},\mathbf{t}^{*})\leq\epsilon$.
We conclude by establishing the convergence rates of tropical descent in practice; we take step sizes given by $a_{m}=\alpha m^{-1/2}\|\nabla f(\mathbf{t}_{m})\|_{{\rm tr}}$, and assume gradients have bounded norm as in the the case of piecewise linear objective functions. In this case, we prove convergence at a rate of $O(\sqrt{m})$.
Proposition 33.
Suppose $f$ is a max-tropical location problem as in Theorem 26. Assume that for all $\mathbf{t}_{m}$, $1/L\leq\|\nabla f(\mathbf{t}_{m})\|_{{\rm tr}}\leq 2$, and assume further there is some $D$ such that $\forall k\leq K,n\in\mathbb{N}:\,d_{{\rm tr}}(\mathbf{t}_{n},\mathbf{x}_{k})\leq D$. We define $a_{m}=\alpha m^{-1/2}\|\nabla f(\mathbf{t}_{m})\|_{{\rm tr}}$. Then for all $m$ satisfying
$\sqrt{m+1}/2-KDNL(N-1)/8\alpha+\sqrt{(\sqrt{m+1}/2-KDNL(N-1)/8\alpha)^{2}+LKN-1}>\sqrt{2},$
we have that:
$d_{{\rm tr}}(\mathbf{t}_{m},V)\leq\frac{2\alpha N}{\sqrt{m+1}-KDNL(N-1)/4\alpha-1}=\epsilon_{m}.$
In particular, if $f$ is max-tropically quasiconvex, then there is some global minimum $\mathbf{t}^{*}$ such that
$d_{{\rm tr}}(\mathbf{t}_{m},\mathbf{t}^{*})\leq\frac{2\alpha N}{\sqrt{m+1}-DNL(N-1)/4\alpha-1}.$
Proof.
We fix $m$ and define:
$M_{1}=\left\lfloor\left(\sqrt{m+1}/2-KDNL(N-1)/8\alpha+\sqrt{(\sqrt{m+1}/2-KDNL(N-1)/8\alpha)^{2}+LKN-1}\right)^{2}\right\rfloor.$
We note that:
$\sqrt{M_{1}}$ $\leq\sqrt{m+1}/2-KDNL(N-1)/8\alpha+\sqrt{(\sqrt{m+1}/2-KDNL(N-1)/8\alpha)^{2}+LKN-1},$
$0\leq\sqrt{M_{1}}$ $\leq\frac{L}{2N}\left(\frac{N\sqrt{m+1}}{L}-\frac{KDN^{2}(N-1)}{4\alpha}+\sqrt{\left(\frac{N\sqrt{m+1}}{L}-\frac{KDN^{2}(N-1)}{4\alpha}\right)^{2}+\frac{4N^{2}(LKN-1)}{L^{2}}}\right).$
So $\sqrt{M_{1}}$ lies between the roots of the polynomial
$p(t)=\frac{N}{L}t^{2}+\left(\frac{KDN^{2}(N-1)}{4\alpha}-\frac{N\sqrt{m+1}}{L}\right)t+\frac{N-LKN^{2}}{L}.$
Therefore:
$0\geq\frac{N}{L}M_{1}+\left(\frac{KDN^{2}(N-1)}{4\alpha}-\frac{N\sqrt{m+1}}{L}\right)\sqrt{M_{1}}+\frac{N-LKN^{2}}{L}$
$\frac{N\sqrt{M_{1}}\sqrt{M_{1}+1}}{L}\leq\left(\frac{N\sqrt{m+1}}{L}-\frac{KDN^{2}(N-1)}{4\alpha}\right)\sqrt{M_{1}}+KN^{2}$
$KDN(N-1)/2+2\alpha L^{-1}(\sqrt{M_{1}+1}-\sqrt{m+1})\leq\frac{2\alpha KN}{\sqrt{M_{1}}}$
Claim 34.
For all $n\geq M_{1}$, we have $a_{n}\leq\epsilon_{m}/N$.
Proof.
We have assumed that
$\sqrt{m+1}/2-KDNL(N-1)/8\alpha+\sqrt{(\sqrt{m+1}/2-KDNL(N-1)/8\alpha)^{2}+LKN-1}>\sqrt{2}$
Then:
$M_{1}$ $\geq\left(\sqrt{m+1}/2-KDNL(N-1)/8\alpha+\sqrt{(\sqrt{m+1}/2-KDNL(N-1)/8\alpha)^{2}+LKN-1}\right)^{2}-1$
$\geq\left(\sqrt{m+1}/2-KDNL(N-1)/8\alpha+\sqrt{(\sqrt{m+1}/2-KDNL(N-1)/8\alpha)^{2}+LKN-1}-1\right)^{2}$
$\sqrt{M_{1}}$ $\geq\sqrt{m+1}/2-KDNL(N-1)/8\alpha+\sqrt{(\sqrt{m+1}/2-KDNL(N-1)/8\alpha)^{2}+LKN-1}-1$
$\geq(\sqrt{m+1}-KDNL(N-1)/4\alpha)-1$
Hence:
$a_{n}$ $\leq\frac{2\alpha}{M_{1}^{1/2}}$
$\leq\frac{2\alpha}{\sqrt{m+1}-KDNL(N-1)/4\alpha-1}=\epsilon_{m}/N$
Claim 35.
For all $n\geq m$, we have $\max_{i}\Delta_{i}(\mathbf{t}_{M_{1}})+s_{M_{1}}-s_{m}\leq K\epsilon_{m}$.
Proof.
We note that
$\max_{i}\Delta_{i}(\mathbf{t})$ $\leq\sum_{i}\Delta_{i}(\mathbf{t})$
$=\sum_{k\leq K}\sum_{i\leq j\leq N}|t_{j}-x_{kj}-t_{i}+x_{ki}|$
$=\sum_{k\leq K}\sum_{i\leq j\leq N}d_{{\rm tr}}(\mathbf{t},\mathbf{x}_{k}),$
and hence
$\max_{i}\Delta_{i}(\mathbf{t}_{M_{1}})$ $\leq\frac{KDN(N-1)}{2}.$
The partial sums are bounded by:
$s_{M_{1}}-s_{m}$ $\leq\sum_{n=M_{1}+1}^{m}-a_{n}$
$\leq\sum_{n=M_{1}+1}^{m}\frac{-\alpha}{Ln^{1/2}}$
$\leq-2\alpha L^{-1}(\sqrt{m+1}-\sqrt{M_{1}+1})$
Therefore:
$\max_{i}\Delta_{i}(\mathbf{t}_{M_{1}})+s_{M_{1}}-s_{m}$ $\leq KDN(N-1)/2-2\alpha L^{-1}(\sqrt{m+1}-\sqrt{M_{1}+1})$
$\leq\frac{2\alpha KN}{\sqrt{M_{1}}}$
$\leq K\epsilon_{m}$
Hence by Theorem 26, we have almost surely that
$d_{{\rm tr}}(\mathbf{t}_{m},V)\leq\epsilon_{m}$
The bound for the tropically quasiconvex case follows from Corollary 32.
4 Numerical Experiments
In this section we compare practical performance of classical descent, tropical descent, SGD, tropical SGD, Adam, Adamax, and Trop Adamax, with a focus on their stability behaviour. The implementation for all these experiments is available in an anonymous GitHub repository at https://github.com/Roroast/TropicalGradDescent.
4.1 Methods
Descent
We will perform classical descent (CD) and tropical descent (TD) using equivalent same step size sequences $a_{n}$ for both methods, which we fix as
$a_{n}=\frac{\gamma\|\nabla f(\mathbf{t}_{n})\|}{\sqrt{n}},$
[…]
Stochastic Gradient Descent
Stochastic gradient descent is a variant of gradient descent which looks to improve computational speed by approximating $\nabla f(t)$; rather than computing the loss and its gradient at $\mathbf{t}$ with respect to the entire dataset $f(\mathbf{t};X)$, we evaluate the loss and its gradient with respect to a random sample $\mathbf{x}_{n}\in X$, $f(\mathbf{t};\mathbf{x}_{n})$. This is simplified with the notation $f_{n}(\mathbf{t})$. As with gradient descent, we use the step size sequence give by
$a_{n}=\frac{\gamma\|\nabla f_{n}(\mathbf{t}_{n})\|}{\sqrt{n}},$
[…]
Adam Variants
The Adam algorithm (Kingma & Ba, (2017)) uses gradient momentum to compound consistent but small effects. This is done by taking first and second moments of the gradient over all steps using exponentially decaying weights. The Adamax algorithm is a variant of the Adam algorithm motivated by the use of the $\infty$-moments rather than second moments.
To adapt the Adam algorithm for the tropical setting, we take moments of an un-normalized tropical descent direction. When looking to take higher moments of $\mathbf{m}_{n}$ over $n$, the $L_{\infty}$ is the natural option for the tropical setting. We therefore look to use the Adamax recursion relation on the $\infty$-moment estimate $\mathbf{u}_{n}$:
$\mathbf{u}_{n+1}=\max(\beta_{2}\mathbf{u}_{n},\mathbf{d}_{n}).$
We therefore define TropAdamax as in Algorithm 2.
Algorithm 2 TropAdamax
$\mathbf{m},\mathbf{u}\leftarrow\mathbf{0}$;
for $n<$ max_steps do
     $\mathbf{g}_{n}\leftarrow\nabla f(\mathbf{t}_{n-1})$;
     $\mathbf{d}_{n}\leftarrow[\mathbf{g}_{n}>0]*(\max(\mathbf{g}_{n})-\min(\mathbf{g}_{n}))$;
     $\mathbf{m}_{n}\leftarrow\beta_{1}*\mathbf{m}_{n-1}+(1-\beta_{1})*\mathbf{d}_{n}$;
     $\mathbf{u}_{n}\leftarrow\max(\beta_{2}*\mathbf{u}_{n-1},|\mathbf{d}_{n}|)$;
     $\mathbf{t}_{n}\leftarrow\mathbf{t}_{n}-\alpha/(1-\beta_{1}^{n})*\mathbf{m}_{n}/\mathbf{u}_{n}$;
4.2 Computation
We test the gradient methods discussed on 12 datasets of different sizes, shapes and dimensions; we consider datasets generated by a branching process, a coalescent process, or a Gaussian distribution on $\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$, for $N=6,28$ and samples of size 10 or 100. See Appendix A for further details on the simulated datasets. The learning rates used have been tuned for each gradient method and dataset size and dimension (Appendix C). In all experiments, random initialisations are taken from a Gaussian distribution on $\mathbb{R}^{N}$. We provide an overview of the general behaviour of each method for the various optimisation problems introduced in section 2.3.
Rather than investigating the relative convergence rates of each method, we are interested in the relative stability of local minima. We therefore plot the CDF of the relative log error after 1000 steps over 50 random initialisations, which can highlight certain error values at which several initialisations have accumulated. When the CDF of one method dominates another at log error $=\epsilon$, we expect it has a higher probability of achieving a log error of at most $\epsilon$.
4.2.1 Centrality Statistics
We first consider the behavior of tropical descent methods for the centrality statistics, Fermat–Weber points and Fréchet means. These optimisation problems are classically convex problems, as well as being both $\max$-tropical and $\min$-tropical location problems, and therefore lend themselves well to gradient methods. The experimental behaviour of these problems is similar, so in this section we include the results for the Fermat–Weber problem while the results for Fréchet means can be found in Appendix E.
Figure 5 shows the CDF of the final log relative error of the non-stochastic methods for the 12 datasets. We note that the stochastic methods achieves the worst error for all 12 datasets, and TropSGD is often performing worse than SGD. The disparity between SGD and TropSGD is greatest when the dimensionality of the data is greater than size of the dataset, while the disparity is smallest when the dimensionality is much smaller than the dataset size. We see similar behaviour in the log error of Tropical Descent and TropAdamax; they perform on par with their classical counterparts for each dataset other than the 28 by 10 datasets, but when the data dimensionality exceeds the number of data points, the tropical gradient methods are not achieving the same accuracy. We also note that convergence of all methods is less consistent for coalescent data than branching or Gaussian data; while it is unclear why this is, we note that coalescent data is supported on the space of ultrametric trees, which is a tropical hyperplane (Ardila & Klivans, (2006); Page et al., (2020)) and may influence the uniqueness of minima.
Figure 5: The CDF of the log relative error after 1000 steps for each gradient method across 50 initialisations when minimising the Fermat–Weber objective. Each subfigure corresponds to a dataset of different dimensionality (columns) and distribution (rows).
Data CD TD SGD TropSGD Adam Adamax TropAdamax
6$\times$10 Branching Data -4.60 -4.60 -4.48 -2.95 -4.60 -4.60 -4.60
6$\times$10 Coalescent Data -4.58 -4.55 -4.38 -1.34 -4.58 -4.58 -4.55
6$\times$10 Gaussian Data -4.60 -4.60 -4.44 -2.63 -4.58 -4.58 -4.60
6$\times$100 Branching Data -4.60 -4.60 -4.45 -3.95 -4.59 -4.59 -4.59
6$\times$100 Coalescent Data -4.55 -4.52 -4.37 -2.76 -4.54 -4.57 -4.54
6$\times$100 Gaussian Data -4.60 -4.60 -4.45 -4.00 -4.59 -4.60 -4.59
28$\times$10 Branching Data -4.60 -4.60 -4.15 -1.90 -4.60 -4.60 -4.56
28$\times$10 Coalescent Data -4.52 -3.76 -3.99 -2.24 -4.52 -4.56 -3.84
28$\times$10 Gaussian Data -4.59 -4.59 -4.09 -2.47 -4.60 -4.60 -4.47
28$\times$100 Branching Data -4.59 -4.59 -4.05 -3.13 -4.59 -4.59 -4.59
28$\times$100 Coalescent Data -4.49 -4.47 -3.86 -2.54 -4.38 -4.56 -4.48
28$\times$100 Gaussian Data -4.59 -4.59 -4.09 -3.02 -4.59 -4.59 -4.56
Table 1: The mean log relative error of each gradient method after 1000 steps for each dataset when minimising the Fermat–Weber objective function. The minimal mean errors for each dataset are in bold.
4.2.2 Wasserstein Projections
The Wasserstein projection problems highlight the differences in convergent behaviour with and without the tropical gradient, as they are $\max$-tropical location problems, but not classically convex. The CDFs of the final log relative errors can be found in Figures 6 and 7.
Figure 6: The CDF of the log relative error after 1000 steps for each gradient method across 50 initialisations when minimising the 2-Wasserstein projection objective. Each subfigure corresponds to a dataset of different dimensionality (columns) and distribution (rows).
Figure 7: The CDF of the log relative error after 1000 steps for each gradient method across 50 initialisations when minimising the $\infty$-Wasserstein projection objective. Each subfigure corresponds to a dataset of different dimensionality (columns) and distribution (rows).
In Figures 6 and 7 we see large steps at various intervals in the CDF plots, which is indicative of stable local minima. This is particularly prevalent in low dimensional space. Classical descent, Adam, and Adamax are particularly susceptible to this, while tropical descent and TropAdamax are often more likely to pass these local minima. This difference is most pronounced when solving the $\infty$-Wasserstein problem in low dimensions—we are almost certain to reach the minimum with small error, while classical methods have a 50-80% chance of achieving the same error.
Data CD TD SGD TropSGD Adam Adamax TropAdamax
6$\times$10 Branching Data -3.52 -4.13 -1.98 -2.17 -3.49 -3.51 -3.99
6$\times$10 Coalescent Data -3.53 -4.31 -2.38 -3.17 -3.77 -3.85 -4.16
6$\times$10 Gaussian Data -3.66 -3.72 -3.21 -2.76 -3.84 -3.85 -3.90
6$\times$100 Branching Data -3.57 -4.38 -3.28 -1.64 -3.94 -3.94 -4.39
6$\times$100 Coalescent Data -4.35 -4.54 -3.31 -2.54 -4.43 -4.44 -4.26
6$\times$100 Gaussian Data -3.78 -4.15 -3.39 -2.69 -3.77 -3.79 -4.42
28$\times$10 Branching Data -3.59 -4.29 -3.18 -2.03 -3.63 -3.64 -4.34
28$\times$10 Coalescent Data -4.07 -3.80 -3.45 -2.50 -4.12 -4.09 -3.82
28$\times$10 Gaussian Data -3.53 -3.89 -3.22 -2.18 -3.61 -3.58 -3.85
28$\times$100 Branching Data -3.65 -4.54 -3.42 -2.64 -3.86 -3.87 -4.53
28$\times$100 Coalescent Data -4.12 -4.57 -3.89 -3.44 -4.20 -4.21 -4.55
28$\times$100 Gaussian Data -3.65 -4.50 -3.39 -2.67 -3.85 -3.86 -4.48
Table 2: The mean log relative error of each gradient method after 1000 steps for each dataset when minimising the $2$-Wasserstein projection objective function. The minimal mean errors for each dataset are in bold.
Data CD TD SGD TropSGD Adam Adamax TropAdamax
6$\times$10 Branching Data -2.56 -4.29 -0.58 -0.94 -2.58 -2.59 -4.28
6$\times$10 Coalescent Data -4.07 -4.48 -1.13 -1.38 -4.13 -4.09 -4.48
6$\times$10 Gaussian Data -4.00 -4.44 -2.08 -3.10 -4.04 -4.02 -4.38
6$\times$100 Branching Data -3.63 -4.46 -2.14 -1.17 -3.60 -3.65 -4.46
6$\times$100 Coalescent Data -4.11 -4.51 -2.19 -2.30 -4.18 -4.14 -4.51
6$\times$100 Gaussian Data -3.02 -4.36 -1.48 -1.23 -3.06 -3.09 -4.31
28$\times$10 Branching Data -3.72 -3.66 -2.19 -1.14 -3.81 -3.81 -3.55
28$\times$10 Coalescent Data -4.31 -4.13 -2.83 -1.99 -4.39 -4.39 -4.01
28$\times$10 Gaussian Data -4.16 -3.47 -2.08 -0.96 -4.43 -4.36 -3.12
28$\times$100 Branching Data -3.55 -3.86 -1.38 -1.16 -3.59 -3.49 -3.71
28$\times$100 Coalescent Data -4.45 -4.12 -2.11 -2.05 -4.54 -4.54 -3.95
28$\times$100 Gaussian Data -3.23 -3.61 -1.74 -1.61 -3.27 -3.22 -2.94
Table 3: The mean log relative error of each gradient method after 1000 steps for each dataset when minimising the $\infty$-Wasserstein projection objective function. The minimal mean errors for each dataset are in bold.
From tables 2 and 3, we see that when solving the 2-Wasserstein projection problem, tropical methods are optimal in all cases other than $28\times 10$ coalescent data. For the $\infty$-Wasserstein projection problem in small dimensions, tropical descent outperforms other methods by some margin, while the comparative performances are less consistent for the $\infty$-Wasserstein projection problem in high dimensions.
4.2.3 Linear Regression
The linear regression problem is a $\min$ tropical location problem, so we run the same code with input $-\mathbf{t}$. As the coalescent data lies on a tropical hyperplane (Ardila & Klivans, (2006); Page et al., (2020)), the minimum of the linear regression function for such data is exactly 0. Figure 8 therefore shows the CDF of the log error rather than the CDF of the log relative error.
As we can see in table 4, the tropical methods outperform their classical counterparts for every dataset, often by a factor of two. The tropical descent method in particular is the most effective for all but one dataset. In Figure 8 we see that the CDF of log errors for tropical descent and TropAdamax consistently dominates those of classical descent, Adam and Adamax, to the extent that the support of their CDFs does not intersect that of their classical counterparts for some datasets. We also note that unlike the previous problems, TropSGD is outperforming SGD when solving the linear regression problem, often to the extent that TropSGD outperforms all classical gradient methods.
Figure 8: The CDF of the log relative error after 1000 steps for each gradient method across 50 initialisations when minimising the linear regression objective. Each subfigure corresponds to a dataset of different dimensionality (columns) and distribution (rows).
Data CD TD SGD TropSGD Adam Adamax TropAdamax
6$\times$10 Branching Data -0.96 -5.96 -0.92 -2.10 -1.09 -1.07 -5.57
6$\times$10 Coalescent Data -4.27 -6.25 -3.80 -5.78 -4.44 -4.54 -6.22
6$\times$10 Gaussian Data -1.13 -5.42 -0.86 -1.33 -1.15 -1.14 -4.79
6$\times$100 Branching Data -1.08 -5.34 -0.86 -1.54 -1.10 -1.11 -3.73
6$\times$100 Coalescent Data -6.42 -6.35 -3.29 -5.69 -6.81 -6.88 -6.22
6$\times$100 Gaussian Data -0.93 -5.43 -1.09 -1.93 -0.92 -0.92 -3.97
28$\times$10 Branching Data -1.60 -5.83 -1.13 -4.47 -1.41 -1.48 -5.53
28$\times$10 Coalescent Data -2.56 -6.52 -2.22 -5.50 -2.68 -2.61 -6.44
28$\times$10 Gaussian Data -1.35 -5.68 -0.91 -3.99 -1.20 -1.23 -5.56
28$\times$100 Branching Data -0.79 -4.33 -0.54 -1.46 -0.69 -0.70 -3.44
28$\times$100 Coalescent Data -3.35 -5.34 -2.10 -6.04 -3.79 -3.94 -4.95
28$\times$100 Gaussian Data -0.93 -3.88 -0.64 -1.22 -0.84 -0.85 -3.44
Table 4: The mean log relative error of each gradient method after 1000 steps for each dataset when minimising the linear regression objective function. The minimal mean errors for each dataset are in bold.
5 Discussion
We have introduced the concept of tropical descent as the steepest descent method with respect to the tropical norm. We have demonstrated, both theoretically and experimentally, that it possesses well-behaved properties with respect to a wide class of optimisation problems on the tropical projective torus, including several key statistical problems for the analysis of phylogenetic data. However, our experimental results are stronger than our theoretical results suggest; the optimisation problems studied are location problems, rather than $\triangle$-star-quasi-convex problems, and yet tropical descent appears to consistently converge to a global minimum, not just the tropical convex hull of our dataset. It would therefore be a natural extension of this work to identify how much we can relax the $\triangle$-star-quasi-convexity condition while guaranteeing convergence.
In this paper, we consider tree data with 4 to 8 leaves, and samples of size 10 to 100—at this scale, we have observed (Trop/)SGD to be a significant sacrifice in accuracy with little benefit in computational time. Furthermore, at this scale, the relative merits of SGD versus its tropical counterpart are inconsistent; TropSGD out-performs SGD when solving the linear regression problem, but performs worse when finding centrality statistics. However, for applications to larger datasets, we would expect the stochastic approaches to be more suitable. It would then be important to understand the theoretical differences between SGD and TropSGD in terms of their convergence guarantees in the tropical setting and their comparative performance in practice.
Finally, as shown by Akian et al., (2023), the tropical linear regression problem is polynomial-time equivalent to mean payoff games - two-player perfect information games on a directed graph (Ehrenfeucht & Mycielski, (1979); Gurvich et al., (1988)). This family of problems of particular interest in game theory as they are known to be NP $\cap$ co-NP. While our methods are designed to approximate the minimum rather than compute it exactly, tropical descent may prove to be an efficient tool when solving large-scale mean payoff games.
Acknowledgments
R.T. receives partial funding from a Technical University of Munich–-Imperial College London Joint Academy of Doctoral Studies (JADS) award (2021 cohort, PIs Drton/Monod). A.M. is supported by the Engineering and Physical Sciences Research Council under grant reference [EP/Y028872/1] and a London Mathematical Society Emmy Noether Fellowship.
Appendix A Data Generation
The same 12 datasets are used for each experiment in Section 4. These datasets live in $\mathbb{R}^{6}/\mathbb{R}\mathbf{1}$ or $\mathbb{R}^{28}/\mathbb{R}\mathbf{1}$; these dimensions correspond to the spaces of trees with 4 and 8 leaves respectively. Branching process datasets are generated using the R package ape (Paradis & Schliep, (2019)) with exponentially distributed branch lengths—this produces data supported on the space of metric trees, i.e., the tropical Grassmannian. Coalescent data has also been generated using ape and is supported on the space of ultrametric trees, which is a tropical linear space. Gaussian data is generated using NumPy (Harris et al., (2020)), and is supported on the general tropical projective torus.
In the computation of the Wasserstein projection objectives, we require a secondary dataset of different dimensionality; for each of the 12 datasets, we generate datasets of the same size and distribution, but in $\mathbb{R}^{3}/\mathbb{R}\mathbf{1}$ or $\mathbb{R}^{21}/\mathbb{R}\mathbf{1}$ which contain the spaces of trees with 3 and 7 leaves respectively.
Appendix B PCA and Logistic Regression
Figure 9: The CDF of the log relative error after 1000 steps for each gradient method across 50 initialisations when minimising the PCA objective. Each subfigure corresponds to a dataset of different dimensionality (columns) and distribution (rows).
Figure 10: The CDF of the log relative error after 1000 steps for each gradient method across 50 initialisations when minimising the logistic regression objective. Each subfigure corresponds to a dataset of different dimensionality (columns) and distribution (rows).
As well as the location problems outlined in this paper, we performed the same numerical experiments with the tropical polytope PCA loss function (Yoshida et al., (2019)) and tropical logistic regression loss function (Aliatimis et al., (2023)). These loss functions take multiple points in $\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ as argument, and we execute our gradient methods on each argument simultaneously. Figures 9 and 10 show the resulting CDF of our gradient methods for these problems. These problems do not lend themselves well to gradient methods, as the relative log error is rather large for all of our proposed methods. However, for both problems the tropical descent and TropAdamax methods are outperforming the classical methods quite consistently.
Appendix C Learning Rate Tuning
The Adam algorithms—Adam, Adamax and TropAdamax—are run using the default $\beta=(0.9,0.999)$ parameters. We then tune the learning rates used for each method; Gradient Descent, Tropical Descent, SGD, Tropical SGD, Adam, Adamax, and TropAdamax. We perform an initial test of learning rates $\gamma=0.0001,0.0003,0.001,0.003,0.01,0.03,0.1,0.3,1,3,10,30$ over 1000 steps for each algorithm, with 5 random initialisations for each dataset. From this, we narrowed our search to take 18 initializations at 0.5 intervals on the log scale in near optimal regions for each problem, gradient method, and data dimensionality. We then selected the learning rate with minimal mean log relative error in each case; tables 5 - 11 show the selected learning rates.
Data Dim Data Count CD TD SGD TropSGD Adam Adamax TropAdamax
6 10 0.135 0.135 0.135 0.0821 0.00409 0.00409 0.0111
100 0.135 0.135 0.0821 0.135 0.00409 0.00409 0.00674
28 10 0.368 0.368 0.368 0.607 0.00674 0.00674 0.0183
100 0.368 0.223 0.368 0.607 0.00674 0.00674 0.0111
Table 5: Learning rates for each gradient method when solving the Fermat–Weber problem.
Data Dim Data Count CD TD SGD TropSGD Adam Adamax TropAdamax
6 10 0.135 0.135 0.135 0.135 0.00674 0.00674 0.00674
100 0.135 0.0821 0.0821 0.223 0.00674 0.00674 0.0111
28 10 0.368 0.223 0.223 0.607 0.00674 0.00674 0.0183
100 0.368 0.135 0.223 0.607 0.00674 0.00674 0.0111
Table 6: Learning rates for each gradient method when solving the Fréchet mean problem.
Data Dim Data Count CD TD SGD TropSGD Adam Adamax TropAdamax
6 10 1.65 7.39 0.0821 0.223 0.135 0.223 0.0821
100 1.65 4.48 0.368 0.135 0.0821 0.135 1.00
28 10 1.65 0.368 0.607 0.607 0.00674 0.0111 0.0183
100 2.72 2.72 0.368 0.607 0.0302 0.0821 0.0498
Table 7: Learning rates for each gradient method when solving the 2-Wasserstein projection problem.
Data Dim Data Count CD TD SGD TropSGD Adam Adamax TropAdamax
6 10 0.0821 0.135 0.0302 0.223 0.00674 0.00674 0.0183
100 0.368 0.135 0.607 0.0498 0.00248 0.00409 0.0183
28 10 0.223 0.607 0.223 0.607 0.00409 0.0111 0.0821
100 0.223 0.607 0.223 0.607 0.00409 0.0111 0.0821
Table 8: Learning rates for each gradient method when solving the $\infty$-Wasserstein projection problem.
Data Dim Data Count CD TD SGD TropSGD Adam Adamax TropAdamax
6 10 2.72 4.48 0.607 1.65 0.135 0.135 0.368
100 2.72 4.48 1.00 0.607 0.0821 0.135 0.607
28 10 7.39 4.48 0.368 1.65 0.0183 0.0183 0.223
100 4.48 7.39 0.368 0.607 0.0302 0.0821 0.0821
Table 9: Learning rates for each gradient method when solving the polytope PCA problem.
Data Dim Data Count CD TD SGD TropSGD Adam Adamax TropAdamax
6 10 7.39 12.2 1.00 2.72 1.00 1.00 0.607
100 7.39 12.2 1.00 0.368 0.607 1.00 1.00
28 10 20.1 245 7.39 90.0 4.48 4.48 2.72
100 20.1 90.0 4.48 90.0 1.65 7.39 2.72
Table 10: Learning rates for each gradient method when solving the logistic regression problem.
Data Dim Data Count CD TD SGD TropSGD Adam Adamax TropAdamax
6 10 0.0302 0.0821 0.223 0.0821 0.00248 0.00248 0.0111
100 0.0498 0.0821 0.368 0.0821 0.00248 0.00248 0.0111
28 10 0.368 0.0821 0.0821 0.135 0.0111 0.0302 0.0111
100 0.223 0.607 1.00 12.2 0.0015 0.00409 0.0498
Table 11: Learning rates for each gradient method when solving the linear regression problem.
Appendix D Runtimes
Experiments were implemented on a AMD EPYC 7742 node using a single core, 8 GB. Table 12 shows the average time taken per initalisation for each gradient method and optimisation problem.
Stats Prob CD TD SGD TropSGD Adam Adamax TropAdamax
Fermat–Weber 0.18s 0.19s 0.26s 0.26s 0.22s 0.23s 0.26s
Fréchet Mean 0.23s 0.23s 0.30s 0.30s 0.28s 0.28s 0.31s
2-Wasserstein Projection 1.18s 1.18s 1.75s 1.77s 1.23s 1.24s 1.27s
$\infty$-Wasserstein Projection 1.17s 1.17s 1.71s 1.71s 1.21s 1.21s 1.24s
Polytope PCA 0.60s 0.61s 0.78s 0.78s 0.64s 0.64s 0.68s
Logistic Regression 0.45s 0.45s 0.62s 0.62s 0.68s 0.50s 0.53s
Linear Regression 0.22s 0.22s 0.27s 0.28s 0.26s 0.27s 0.30s
Table 12: Average runtimes for 1000 steps of each gradient method when minimising each objective function.
Appendix E Further Figures
Here we include the results for experiments run for the computation of Fréchet means.
Data CD TD SGD TropSGD Adam Adamax TropAdamax
6$\times$10 Branching Data -4.59 -4.56 -4.01 -2.93 -4.58 -4.59 -4.54
6$\times$10 Coalescent Data -4.58 -4.49 -4.06 -1.86 -4.57 -4.58 -4.50
6$\times$10 Gaussian Data -4.59 -4.35 -4.10 -2.63 -4.58 -4.59 -4.32
6$\times$100 Branching Data -4.60 -4.59 -4.36 -3.90 -4.59 -4.59 -4.59
6$\times$100 Coalescent Data -4.57 -4.57 -4.44 -3.14 -4.54 -4.56 -4.55
6$\times$100 Gaussian Data -4.59 -4.59 -4.44 -4.03 -4.59 -4.60 -4.59
28$\times$10 Branching Data -4.57 -3.69 -3.66 -1.99 -4.54 -4.57 -3.79
28$\times$10 Coalescent Data -4.56 -3.72 -4.01 -2.57 -4.52 -4.57 -3.71
28$\times$10 Gaussian Data -4.56 -4.05 -3.71 -2.49 -4.54 -4.57 -4.02
28$\times$100 Branching Data -4.59 -4.58 -3.99 -3.23 -4.58 -4.59 -4.58
28$\times$100 Coalescent Data -4.54 -4.56 -4.14 -2.87 -4.46 -4.57 -4.52
28$\times$100 Gaussian Data -4.59 -4.55 -3.90 -3.01 -4.59 -4.59 -4.58
Table 13: The mean log relative error of each gradient method after 1000 steps for each dataset when minimising the Fréchet mean objective function. The minimal mean errors for each dataset are in bold.
Figure 11: The CDF of the log relative error after 1000 steps for each gradient method across 50 initialisations when minimising the Fréchet mean objective. Each subfigure corresponds to a dataset of different dimensionality (columns) and distribution (rows).

### v2

@@
 Abstract
-We propose a gradient descent method for solving optimisation problems arising in settings of tropical geometry—a variant of algebraic geometry that has become increasingly studied in applications such as computational biology, economics, and computer science.
-Our approach takes advantage of the polyhedral and combinatorial structures arising in tropical geometry to propose a versatile approach for approximating local minima in tropical statistical optimisation problems—a rapidly growing body of work in recent years.
-Theoretical results establish global solvability for 1-sample problems and a convergence rate of $O(1/\sqrt{k})$.
-Numerical experiments demonstrate the method’s superior performance over classical descent for tropical optimisation problems which exhibit tropical convexity but not classical convexity.
-Notably, tropical descent seamlessly integrates into advanced optimisation methods, such as Adam, offering improved overall performance.
+We propose a gradient descent method for solving optimization problems arising in settings of tropical geometry—a variant of algebraic geometry that has become increasingly studied in applications such as computational biology, economics, and computer science.
+Our approach takes advantage of the polyhedral and combinatorial structures arising in tropical geometry to propose a versatile approach for approximating local minima in tropical statistical optimization problems—a rapidly growing body of work in recent years.
+Theoretical results establish global solvability for 1-sample problems and a convergence rate matching classical gradient descent.
+Numerical experiments demonstrate the method’s superior performance over classical descent for tropical optimization problems which exhibit tropical convexity but not classical convexity.
+We also demonstrate the seamless integration of tropical descent into advanced optimization methods, such as Adam, offering improved overall performance.
 Keywords:
@@
 1 Introduction
-Where algebraic geometry studies geometric properties that can be expressed algebraically of solution sets of systems of multivariate polynomials, tropical geometry restricts to the case of polynomials defined by linearising operations, where the sum of two elements is their maximum and the product of two elements is their sum.
+Where algebraic geometry studies geometric properties of solution sets of systems of multivariate polynomials, tropical geometry restricts to the case of polynomials defined by linearizing operations, where the “sum” of two elements is their maximum and the “product” of two elements is their sum.
 Evaluating polynomials with these operations results in piecewise linear functions.
-Tropical geometry is a relatively young field of pure mathematics established in the 1990s (with roots dating back further) and has recently become an area of active interest in computational and applied mathematics for its relevance to the problem of dynamic programming in computer science (Maclagan & Sturmfels, (2021)); in mechanism design for two-player games in game theory (Lin & Tran, (2019)); and statistical analyses of phylogenetic trees (Monod et al., (2022)).
-Given the connection of tropical geometry to applied sciences, developing and implementing computational techniques for tropical geometric settings is becoming increasingly important.
-A central task to a vast majority of computational tasks and training machine learning models entails solving an optimisation problem.
-In this paper, we are motivated by the optimisation problems in statistics in tropical geometric settings.
-We focus on the classical gradient descent approach for optimisation and adapt it to the setting of tropical geometry.
-We thus provide a unifying gradient descent framework to solve the optimisation problems entailed in the computation of a wide range of statistical quantities and methods in tropical spaces which are an area of active interest in applied tropical geometry.
-The remainder of this paper is organised as follows.
-The following section contains the necessary foundations of tropical geometry to present, an introduction to tropical location problems functions (Comăneci, (2023)), and the array of statistical optimisation tasks studied in this paper.
-In Section 3 we formulate steepest descent with respect to the tropical norm and present the theoretical guarantees for tropical descent, most notably its convergence for a wide class of tropically quasi-convex functions and a convergence rate of $O(1/\sqrt{k})$.
-In Section 4, we perform a comprehensive numerical study of our proposed tropical gradient descent method and its performance in the statistical optimisation problems of interest.
+Tropical geometry is a relatively young field of pure mathematics established in the 1990s (with roots dating back further) and has recently become an area of active interest in computational and applied mathematics for its relevance to the statistical analyses of phylogenetic trees (Monod et al.,, 2018), the problem of dynamic programming in computer science (Maclagan & Sturmfels,, 2015), and in mechanism design for two-player games in game theory (Lin & Tran,, 2019).
+Following the identification of the tropical Grassmannian and the space of phylogenetic trees (Speyer & Sturmfels,, 2004) in particular, there has been surge of research (Barnhill et al.,, 2024b; Lin & Yoshida,, 2018; Comăneci & Joswig,, 2024; Monod et al.,, 2018) which necessitates the development of new computational techniques for optimisation in tropical geometric settings.
+The training of machine learning models and the vast majority of computational tasks in statistics entail solving optimization problems; this paper is motivated by statistical optimization problems in tropical geometric settings.
+Initial research into tropical statistics considered optimization problems such as the identification of Fermat–Weber points which, due to the piecewise linear nature of tropical geometry, can be re-framed as a linear program (Comăneci & Joswig,, 2024; Lin & Yoshida,, 2018).
+Unfortunately, the complexity of these linear programs often scales with the size of our dataset, so gradient descent is more commonly implemented in practice (Barnhill et al.,, 2024a, b).
+More recently—following the identification of ReLU neural networks and tropical rational functions (Zhang et al.,, 2018)—more sophisticated neural network optimization problems are being studied in the tropical setting (Pasque et al.,, 2024; Yoshida et al.,, 2024), which necessitates the use of gradient methods for optimization in the tropical setting.
+In this paper, we address the natural question—given the need for gradient methods in tropical data science, how can we tailor classical gradient methods to the tropical setting?
+We provide a foundational tropical gradient descent framework to solve the optimization problems which arise in the computation of a wide range of tropical statistics, an area of active interest in applied tropical geometry.
+The remainder of this paper is organized as follows.
+The following section contains the necessary foundations of tropical geometry, an introduction to tropical location problems (Comăneci,, 2024), and the array of statistical optimization tasks studied in this paper.
+Section 3 contains the theoretical contributions of this paper; we formulate steepest descent with respect to the tropical norm and present the theoretical guarantees for tropical descent, most notably its convergence for a wide class of tropically quasi-convex functions and a convergence rate of $O(1/\sqrt{k})$.
+In Section 4, we perform a comprehensive numerical study of our proposed tropical descent method and its performance in the statistical optimization problems of interest.
 We conclude with a discussion of possible future work on tropical gradient methods in Section 5.
 2 Background and Preliminaries
-In this section, we overview the background on tropical geometry and present the statistical concepts and methodologies which are defined by an optimisation problem that we will study in the tropical setting.
+In this section, we overview the background on tropical geometry and present the statistical optimization problems which we study in the tropical setting.
 2.1 Tropical Geometry
-Here, we review the algebraic and geometric structure of the tropical projective torus, the state space for phylogenetic data and the domain of our tropical statistical optimisation problems.
+Here, we review the algebraic and geometric structure of the tropical projective torus, the state space for phylogenetic data and the domain of our tropical statistical optimization problems.
 Definition 1 (Tropical Algebra).
@@
 Definition 2 (Tropical Projective Torus).
-The $n$-dimensional tropical projective torus is a quotient space constructed by endowing $\mathbb{R}^{n+1}$ with the equivalence relation
+The $n$-dimensional tropical projective torus is a quotient space constructed by endowing $\mathbb{R}^{N}$ with the equivalence relation
 $\mathbf{x}\sim\mathbf{y}\Leftrightarrow\exists\,a:\>\mathbf{x}=a\odot\mathbf{y};$ (1)
-it is denoted by $\mathbb{R}^{n+1}/\mathbb{R}\mathbf{1}$. The generalised Hilbert projective metric, also referred to as the tropical metric, is given by
+it is denoted by $\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$. The generalized Hilbert projective metric, also referred to as the tropical metric, is given by
 $d_{{\rm tr}}(\mathbf{x},\mathbf{y})=\max_{i}\{x_{i}-y_{i}\}-\min_{i}\{x_{i}-y_{i}\}=\max_{i,j}\{x_{i}-y_{i}-x_{j}+y_{j}\}.$
@@
 $\|\mathbf{x}\|_{{\rm tr}}=\max_{i}x_{i}-\min_{i}x_{i}.$
-The tropical projective torus is the ambient space containing the tropical Grassmannian, which is equivalent to the space of phylogenetic trees (Speyer & Sturmfels, (2004)).
-This has sparked the study of tropical data science, formalising statistical techniques which respect the tropical geometry of the state space (Monod et al., (2018); Yoshida, (2020)).
-The metric above is widely accepted and preferred for the development of geometric statistical tools on the tropical projective torus (Yoshida, (2020)).
-However, recent work has shown theoretical advantages to the use of an asymmetric metric (Comăneci & Joswig, (2023)).
+Remark 3.
+While a single point in the tropical projective torus is given by some ray $\mathbf{x}+\mathbb{R}\mathbf{1}$, we note that there is a unique representative of this equivalence class whose coordinates sum to zero.
+In taking such representatives, we can identify the tropical projective torus $\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ and the hyperplane =$\mathcal{H}=\{\mathbf{x}\in\mathbb{R}^{N}:\sum x_{i}=0\}$.
+Throughout this paper, we use this homeomorphism to visualize the tropical projective torus (e.g., Figures 1 and 2).
+The tropical projective torus is the ambient space containing the tropical Grassmannian, which is equivalent to the space of phylogenetic trees (Speyer & Sturmfels,, 2004).
+This has sparked the study of tropical data science, formalizing statistical techniques which respect the tropical geometry of the state space (Monod et al.,, 2018; Yoshida,, 2022).
+The generalized Hilbert projective metric defined above is widely accepted and preferred for the development of geometric statistical tools on the tropical projective torus (Yoshida,, 2022).
+However, recent work has shown theoretical advantages to the use of an asymmetric metric (Comăneci & Joswig,, 2024).
 Definition 4 (Tropical Asymmetric Distance).
-The tropical asymmetric distance on $\mathbb{R}^{n}/\mathbb{R}\mathbf{1}$ is given by
+The tropical asymmetric distances on $\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ are given by
 $d_{\triangle_{\min}}(\mathbf{a},\mathbf{b})\coloneqq\sum_{i}(b_{i}-a_{i})-N\min_{j}(b_{j}-a_{j}),$
@@
 Remark 5.
-The asymmetric distances act as an $\ell_{1}$ norm on the tropical projective torus, which the symmetric metric $d_{\mathrm{tr}}$ acts as an $\ell_{\infty}$ norm.
+The asymmetric distances act as an $\ell_{1}$ norm on the tropical projective torus, while the symmetric metric $d_{\mathrm{tr}}$ acts as an $\ell_{\infty}$ norm.
 We note that $Nd_{{\rm tr}}=d_{\triangle_{\min}}+d_{\triangle_{\max}}$.
@@
 We use the $\max$-convention for these definitions, though we note each has a $\min$-convention equivalent.
-Definition 5 (Tropical Hyperplanes).
-The tropical hyperplane $\mathcal{H}_{\mathbf{a}}$ defined by the linear form $a_{1}\odot x_{1}\boxplus\dots\boxplus a_{n}\odot x_{n}$ is given by
+Definition 6 (Tropical Hyperplane).
+The tropical hyperplane $\mathcal{H}_{\mathbf{a}}$ defined by the linear form $a_{1}\odot x_{1}\boxplus\cdots\boxplus a_{N}\odot x_{N}$ is given by
 $\mathcal{H}_{\mathbf{a}}=\{\mathbf{x}:\exists\,i\neq j\text{ s.t. }a_{i}+x_{i}=a_{j}+x_{j}=\max\{a_{k}+x_{k}\}\}.$
-Definition 6 (Tropical Line Segments).
+Definition 7 (Tropical Line Segment).
 For any two points $\mathbf{a},\mathbf{b}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$, the tropical line segment between $\mathbf{a}$ and $\mathbf{b}$ is the set
@@
 Tropical line segments define a unique geodesic path between any two points, but general geodesics are not uniquely defined on the tropical projective torus.
-$A$
-$B$$A$
-$A$$B$$C$
+$\mathbf{a}$
+$\mathbf{b}$$\mathbf{a}$
+$\mathbf{a}$$\mathbf{b}$$\mathbf{c}$
 Figure 2: A tropical hyperplane, tropical line segment and tropical convex hull in $\mathbb{R}^{3}/\mathbb{R}\mathbf{1}\cong\mathcal{H}=\{\mathbf{x}:\sum x_{i}=0\}$.
 The dashed lines show coordinate directions.
-Definition 7 (Tropical Convex Hulls (Develin & Sturmfels,, 2004)).
-A set $S$ is $\max$-tropically convex if it contains $\alpha\odot\mathbf{x}\boxplus\beta\odot\mathbf{y}$ for all $\mathbf{x},\mathbf{y}\in S$ and $\alpha,\beta\in\mathbb{R}$.
-For a finite subset $X=\{\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\}\subset\mathbb{R}^{n}/\mathbb{R}\mathbf{1}$, the max-plus tropical convex hull of $X$ is the set of all max-tropical linear combinations of points in $X$,
-$\operatorname{tconv}_{\max}(X)=\{\alpha_{1}\odot\mathbf{x}_{1}\boxplus\alpha_{2}\odot\mathbf{x}_{2}\boxplus\cdots\boxplus\alpha_{n}\odot\mathbf{x}_{n}\mid\alpha_{1},\ldots,\alpha_{n}\in\mathbb{R}\}.$
-Similarly, the min-tropical convex hull of $X$, $\operatorname{tconv}_{\min}(X)$ is the set of all min-tropical linear combinations of points in $X$.
-A study of tropically convex functions requires an understanding of both max-tropical and min-tropical convex hulls, as we will see in our following consideration of tropical location problems.
+Definition 8 (Tropical Convex Hull (Develin & Sturmfels,, 2004)).
+A set $S\subset\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ is $\max$-tropically convex if it contains $\alpha\odot\mathbf{x}\boxplus\beta\odot\mathbf{y}$ for all $\mathbf{x},\mathbf{y}\in S$ and $\alpha,\beta\in\mathbb{R}$.
+For a finite subset $X=\{\mathbf{x}_{1},\ldots,\mathbf{x}_{K}\}\subset\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$, the max-plus tropical convex hull of $X$ is the set of all max-tropical linear combinations of points in $X$,
+$\operatorname{tconv}_{\max}(X)\coloneqq\{\alpha_{1}\odot\mathbf{x}_{1}\boxplus\cdots\boxplus\alpha_{K}\odot\mathbf{x}_{K}\mid\alpha_{1},\ldots,\alpha_{K}\in\mathbb{R}\}.$
+Similarly, the min-tropical convex hull of $X$, $\operatorname{tconv}_{\min}(X)$, is the set of all min-tropical linear combinations of points in $X$.
+A study of tropically convex functions requires an understanding of both max-tropical and min-tropical convexity, as we will see in our following consideration of tropical location problems.
 2.2 Tropical Location Problems
-Tropical location problems are a new avenue of research, motivated by the optimisation of statistical loss functions; they involve the minimisation over some dataset of some loss function which heuristically increases with distance from some kernel.
-Here we outline the relevant definitions of tropical location problems and their theoretical behaviour as presented by Comăneci, (2023).
-We introduce the problem of tropical linear regression (Akian et al., (2023)) as motivation.
+Tropical location problems are a new avenue of research, motivated by the optimization of statistical loss functions; they involve the minimization over some dataset of some loss function which heuristically increases with distance from some kernel.
+Here we outline the relevant definitions of tropical location problems and their theoretical behavior as presented by Comăneci, (2024).
+We introduce the problem of tropical linear regression (Akian et al.,, 2023) as motivation.
 2.2.1 Motivating Example
-The problem of tropical linear regression looks to find the best-fit tropical hyperplane which minimises the maximal distance to a set of data points.
-While this is an inherently statistical optimisation problem, Akian et al., (2023) prove it to be polynomial-time equivalent to solving mean payoff games.
-A tropical hyperplane is uniquely defined by a cone point $\mathbf{t}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$, allowing us to formulate the best-fit tropical hyperplane as an optimisation problem over $\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$.
-Definition 8 (Tropical Linear Regression (Akian et al., (2023))).
+The problem of tropical linear regression looks to find the best-fit tropical hyperplane which minimizes the maximal distance to a set of data points.
+While this is an inherently statistical optimization problem, it has recently been proven to be polynomial-time equivalent to solving mean payoff games (Akian et al.,, 2023).
+A tropical hyperplane is uniquely defined by a cone point $\mathbf{t}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$, allowing us to formulate the best-fit tropical hyperplane as an optimization problem over $\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$.
+Definition 9 (Tropical Linear Regression (Akian et al.,, 2023)).
 Let $X=\{\mathbf{x}_{1},\dots,\mathbf{x}_{K}\}\subset\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ be our dataset.
@@
 where $i,j$ are such that $\forall\,\ell\neq i,j:(\mathbf{x}_{k}-\mathbf{t})_{i}\geq(\mathbf{x}_{k}-\mathbf{t})_{j}\geq(\mathbf{x}_{k}-\mathbf{t})_{\ell}$.
-In Figure 3, we see a heatmap of this objective function over the tropical projective plane.
+In Figure 3, we see a heat map of this objective function over the tropical projective plane for a sample of size 3.
 We see immediately that the objective is not convex.
 The linear regression objective has sparse gradient almost everywhere, which produces valleys.
-These properties are considered unfavourable extreme cases for gradient methods, but are particularly common in tropical optimisation problems.
+These properties are considered unfavourable extreme cases for gradient methods, but are particularly common in tropical optimization problems.
 While the tropical linear regression problem is not convex, it is quasi-convex along tropical line segments; that is, its sub-level sets are tropically convex.
 Figure 3: A heat map of a 3-sample linear regression objective on $\mathbb{R}^{3}/\mathbb{R}\mathbf{1}\cong\mathcal{H}=\{\mathbf{x}:\sum x_{i}=0\}$.
-2.2.2 $\triangle$-Star Quasi-Convexity
-In formalising tropical location problems, Comăneci, (2023) first considers functions which increase along tropical geodesics from a kernel $\mathbf{v}$; that is, for any point $\mathbf{t}$ on a geodesic from $\mathbf{v}$ to $\mathbf{w}$, we have that $f(\mathbf{t})\leq f(\mathbf{w})$.
+2.2.2 Quasi-Convexity of Tropical Functions
+In formalizing tropical location problems, the work of Comăneci, (2024) first considers functions which increase along tropical geodesics from a kernel $\mathbf{v}$; that is, for any point $\mathbf{t}$ on a geodesic from $\mathbf{v}$ to $\mathbf{w}$, we have that $f(\mathbf{t})\leq f(\mathbf{w})$.
 This property is equivalent to star-convexity of sublevel sets.
@@
 Definition 10 (Oriented Geodesic Segment).
-The (oriented) geodesic segment between $\mathbf{a},\mathbf{b}\in\mathbb{R}^{n}/\mathbb{R}\mathbf{1}$, is given by $[\mathbf{a},\mathbf{b}]_{\triangle_{\max}}\coloneqq\{\mathbf{x}\in\mathbb{R}^{n}/\mathbb{R}\mathbf{1}:d_{\triangle_{\max}}(\mathbf{a},\mathbf{x})+d_{\triangle_{\max}}(\mathbf{x},\mathbf{b})=d_{\triangle_{\max}}(\mathbf{a},\mathbf{b})\}$.
+The (oriented) geodesic segment between $\mathbf{a},\mathbf{b}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$, is given by $[\mathbf{a},\mathbf{b}]_{\triangle_{\min}}\coloneqq\{\mathbf{x}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}:d_{\triangle_{\min}}(\mathbf{a},\mathbf{x})+d_{\triangle_{\min}}(\mathbf{x},\mathbf{b})=d_{\triangle_{\min}}(\mathbf{a},\mathbf{b})\}$.
+The oriented geodesic segment is the union of all geodesics from $\mathbf{a}$ to $\mathbf{b}$ with respect to $d_{\triangle_{\min}}$.
 Definition 11 ($\triangle_{\min}$-Star Convex Set).
-A set $K\subseteq\mathbb{R}^{n}/\mathbb{R}\mathbf{1}$ is a $\triangle_{\max}$-star-convex set with kernel $\mathbf{v}$ if, for every $\mathbf{w}\in K$, we have $[\mathbf{v},\mathbf{w}]_{\triangle_{\max}}\subseteq K$.
-Using the above definitions, we can characterise a loss function which is increasing along tropical geodesics as a function with $\triangle_{\max}$-star-convex sublevel sets.
+A set $K\subseteq\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ is a $\triangle_{\min}$-star-convex set with kernel $\mathbf{v}$ if, for every $\mathbf{w}\in K$, we have $[\mathbf{v},\mathbf{w}]_{\triangle_{\min}}\subseteq K$.
+Using the above definitions, we can characterize a loss function which is increasing along tropical geodesics as a function with $\triangle_{\min}$-star-convex sublevel sets.
 The following definition gives a closed form expression for such functions, while Theorem 13 proves their equivalence.
-Definition 11 (Comăneci, (2023)).
-A function $f:\mathbb{R}^{n}/\mathbb{R}\mathbf{1}\rightarrow\mathbb{R}$ is $\triangle_{\max}$-star-quasi-convex with kernel $\mathbf{v}$ if $f(\mathbf{x})=\hat{\gamma}(\mathbf{x}-\mathbf{v})$ for some increasing $\gamma:\mathbb{R}^{n}_{\geq 0}\rightarrow\mathbb{R}$, where $\hat{\gamma}(\mathbf{x})=\gamma((\max_{i}x_{i})\mathbf{1}-\mathbf{x})$.
-Theorem 12 (Theorem 17, Comăneci, (2023)).
-A continuous function $f:\mathbb{R}^{n}/\mathbb{R}\mathbf{1}\rightarrow\mathbb{R}$ is $\triangle_{\max}$-star-quasi-convex with kernel $\mathbf{v}$ if and only if all of its non-empty sub-level sets are $\triangle_{\max}$-star convex with kernel $\mathbf{v}$.
-These $\triangle_{\max}$-star-convex functions act as a loss function with respect to a single datapoint.
+Definition 12 ($\triangle_{\min}$-Star-Quasi-Convex Functions (Comăneci,, 2024)).
+A function $f:\mathbb{R}^{N}/\mathbb{R}\mathbf{1}\rightarrow\mathbb{R}$ is $\triangle_{\min}$-star-quasi-convex with kernel $\mathbf{v}$ if $f(\mathbf{x})=\hat{\gamma}(\mathbf{x}-\mathbf{v})$ for some $\gamma:\mathbb{R}^{N}_{\geq 0}\rightarrow\mathbb{R}$ which is increasing in every coordinate, where we define $\hat{\gamma}(\mathbf{x})\coloneqq\gamma(\mathbf{x}-(\min_{i}x_{i})\mathbf{1})$.
+Theorem 13 (Theorem 17 of (Comăneci,, 2024)).
+A continuous function $f:\mathbb{R}^{N}/\mathbb{R}\mathbf{1}\rightarrow\mathbb{R}$ is $\triangle_{\min}$-star-quasi-convex with kernel $\mathbf{v}$ if and only if all of its non-empty sub-level sets are $\triangle_{\min}$-star convex with kernel $\mathbf{v}$.
+Remark 14.
+We can define $\triangle_{\max}$-star-quasi-convex functions similarly, but we also note that a function $f(\mathbf{t})$ is a $\triangle_{\max}$-star-quasi-convex function if and only if $f(-\mathbf{t})$ is a $\triangle_{\min}$-star-quasi-convex function.
+These $\triangle_{\min}$-star-convex functions act as a loss function with respect to a single data point.
 For example, the tropical metric $d_{\mathrm{tr}}$ is both $\triangle_{\min}$ and $\triangle_{\max}$-star-quasi-convex.
 The distance to a hyperplane $d_{\mathrm{tr}}(\mathbf{x},\mathcal{H}_{\mathbf{t}})$ is $\triangle_{\max}$-star-quasi-convex in $\mathbf{x}$, but $\triangle_{\min}$-star-quasi-convex in $\mathbf{t}$.
-As a final remark, we note that $\triangle_{\max}$-star-quasi-convex functions are a special case of tropically quasi-convex functions by the following lemma.
-We limit our theoretical considerations to $\triangle_{\max}$-star-quasi-convex functions to utilise their closed form expression given in Definition 11, however numerical experiments demonstrate strong performance for more general tropical quasi-convex functions.
-Lemma 13 (Comăneci, (2023)).
-Any $\triangle_{\max}$-star-convex set is $\max$-tropically convex. Hence, any $\triangle_{\max}$-star-quasi-convex function is tropically quasiconvex in that it has tropically convex sublevel sets.
+As a final remark, we note that $\triangle_{\min}$-star-quasi-convex functions are a special case of tropically quasi-convex functions by the following lemma.
+We limit our theoretical considerations to $\triangle_{\min}$-star-quasi-convex functions to utilize their closed form expression given in Definition 12, however numerical experiments demonstrate strong performance of our methodology for more general tropical quasi-convex functions.
+Proposition 15 (Proposition 8 of (Comăneci,, 2024)).
+Any $\triangle_{\min}$-star-convex set is $\min$-tropically convex. Hence, any $\triangle_{\min}$-star-quasi-convex function is tropically quasi-convex in that it has tropically convex sublevel sets.
 2.2.3 Tropical Location Problems
-The $\triangle$-star-convex functions defined above are generally a measure of closeness to the kernel $\mathbf{v}$, but in minimising a statistical loss function we look at closeness to a sample of points.
-A location problem (Laporte et al., (2019)) measures the closeness to a sample rather than a single kernel.
-Definition 14 (Max(/min)-tropical location problem (Comăneci, (2023))).
-Consider a dataset $X=\{\mathbf{x}_{1},\dots,\mathbf{x}_{K}\}\subset\mathbb{R}^{n}/\mathbb{R}\mathbf{1}$.
-For $k\leq K$, let $h_{k}$ be a $\triangle_{\max}$-star-quasi-convex function with kernel $\mathbf{x}_{k}$, and let $g:\mathbb{R}^{m}\rightarrow\mathbb{R}$ be an increasing function.
-A max-tropical location problem is the minimisation of an objective function $f:\mathbb{R}^{n}/\mathbb{R}\mathbf{1}\rightarrow\mathbb{R}$ given by $h=g(h_{1},\dots,h_{m})$.
-A min-tropical location problem is defined similarly, where the $h_{k}$ are $\triangle_{\min}$-star-quasi-convex.
-The main result by Comăneci, (2023) is the following; tropical location problems contain a minimum in the tropical convex hull of its dataset $X$.
-Theorem 15 (Theorem 20, Comăneci, (2023)).
-Let $f$ be the objective function of a max-tropical location problem.
-Then there is a minimum of $f$ belonging to $\operatorname{tconv}_{\min}(X)$.
-Conversely, min-tropical location problems have minima belonging to $\operatorname{tconv}_{\max}(X)$.
-We can now revisit our motivating example (Definition 8), showing that tropical linear regression is not just a min-tropical location problem, but also demonstrates tropical quasi-convexity.
+The $\triangle_{\min}$-star-convex functions defined above are generally a measure of closeness to the kernel $\mathbf{v}$, but in minimizing a statistical loss function we look at closeness to a sample of points.
+A location problem (Laporte et al.,, 2019) measures the closeness to a sample rather than a single kernel.
+Definition 16 (Min(/max)-Tropical Location Problem (Comăneci,, 2024)).
+Let $X=\{\mathbf{x}_{1},\dots,\mathbf{x}_{K}\}\subset\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ be our dataset.
+For $k\leq K$, let $h_{k}$ be a $\triangle_{\min}$-star-quasi-convex function with kernel $\mathbf{x}_{k}$, and let $g:\mathbb{R}^{K}\rightarrow\mathbb{R}$ be increasing in every coordinate.
+A min-tropical location problem is the minimization of an objective function $f:\mathbb{R}^{N}/\mathbb{R}\mathbf{1}\rightarrow\mathbb{R}$ given by $f=g(h_{1},\dots,h_{K})$.
+A max-tropical location problem is defined similarly, where the $h_{k}$ are $\triangle_{\max}$-star-quasi-convex.
+The main result of (Comăneci,, 2024) is the following: tropical location problems contain a minimum in the tropical convex hull of its dataset $X$.
+Theorem 17 (Theorem 20 of (Comăneci,, 2024)).
+Let $f$ be the objective function of a min-tropical location problem.
+Then there is a minimum of $f$ belonging to $\operatorname{tconv}_{\max}(X)$.
+Conversely, max-tropical location problems have minima belonging to $\operatorname{tconv}_{\min}(X)$.
+We can now revisit our motivating example (Definition 9), showing that tropical linear regression is not just a min-tropical location problem, but also demonstrates tropical quasi-convexity in that it has tropically convex sublevel sets.
 Proposition 18.
-The tropical linear regression problem is a tropically quasi-convex $\min$-tropical location problem.
-Proof.
+The tropical linear regression problem is a $\min$-tropical location problem with tropically convex sublevel sets.
+Proof
 $f(\mathbf{t})$ $=\max_{k\leq K}\min_{j}[\max_{i}[x_{ki}-t_{i}]-x_{kj}+t_{j}]$
 $=\max_{k\leq K}\min_{j}[t_{j}-x_{kj}-\min_{i}[t_{i}-x_{ki}]]$
-These are min-tropical location problems with:
+This is a min-tropical location problem with:
 $g(\mathbf{h})$ $=\max_{k\leq K}h_{k},$
@@
 $\gamma(\mathbf{t})$ $=\min_{j}t_{i}.$
-This $g$ and $\gamma$ are increasing and each $h_{k}$ has kernel $x_{k}$, satisfying the conditions for a min-tropical location problem.
-As the $h_{k}$ are $\triangle_{\min}$-star-convex, their sublevel sets are min-tropically convex.
+This $g$ and $\gamma$ are increasing in each coordinate and each $h_{k}$ has kernel $\mathbf{x}_{k}$, satisfying the conditions for a min-tropical location problem.
+As the $h_{k}$ are $\triangle_{\min}$-star-convex, their sublevel sets are min-tropically convex by Proposition 15.
 This is preserved when taking a maximum over samples, so $f$ also has $\min$-tropically convex sublevel sets.
 ∎
-As mentioned in Lemma 13, $\triangle_{\min}$-star-quasi-convexity implies tropical quasi-convexity.
-However, the tropical linear regression problem is a case in which the reverse implication does not hold; for any sample of size $\geq 1$, the linear regression loss function is not $\triangle_{\min}$-star-quasi-convex.
+As mentioned in Proposition 15, $\triangle_{\min}$-star-quasi-convexity implies tropical quasi-convexity.
+However, the tropical linear regression problem is a case in which the reverse implication does not hold; in general, the linear regression loss function is not $\triangle_{\min}$-star-quasi-convex.
 2.3 Tropical Data Science
-We now review various tropical statistical optimisation problems which have been introduced in the literature in recent years, and which are used to test our tropical descent methodology.
+While young, the field of tropical data science is motivated by the crucial question of statistical learning from the increasing volume of phylogenetic tree data (Kahn,, 2011).
+In contrast to the BHV (Billera et al.,, 2001) or Robinson–Foulds metrics (Robinson & Foulds,, 1981), the tropical interpretation of tree space provides both interpretable geometry and computational efficiency (Steel & Penny,, 1993; Monod et al.,, 2018).
+Many of the statistical methods defined for the tropical setting are framed as optimisation problems (Barnhill et al.,, 2024b), and in this subsection we review some such tropical statistical optimization problems which we use to test our tropical descent methodology.
 We note the varying degrees of convexity demonstrated by each of them.
-These statistical concepts and methods showcase the breadth of applicability of our proposed methodology: optimisation is a central task in computing important statistical quantities, tasks in both descriptive and inferential statistics, and distances between probability measures.
-Adapting these statistical tasks to the setting of tropical geometry is not straightforward and it is not immediate that a single tropical optimisation method can be used to find solutions to the optimisation problems.
+While we restrict our considerations through the rest of this paper to statistics which arise from tropical location problems, in Appendix A we discuss other tropical statistical optimization problems.
 2.3.1 Centrality Statistics
-Our first problems of interest, Fermat–Weber Points and Fréchet means, are centrality statistics for data on a general metric space; Fermat–Weber points (Wesolowsky, (1993)) act as a generalised median, while Fréchet means (Fréchet, (1948)) are a generalisation of the mean.
-Definition 17 (Tropical Fermat–Weber Points (Lin & Yoshida, (2018))).
-Let $X=\{\mathbf{x}_{1},\dots,\mathbf{x}_{K}\}\subset\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ be our dataset. A solution to the following optimisation problem is a tropical Fermat–Weber point:
+Our first problems of interest, Fermat–Weber points and Fréchet means, are centrality statistics for data on a general metric space; Fermat–Weber points (Drezner et al.,, 2004) act as a generalized median, while Fréchet means (Fréchet,, 1948) are a generalization of the mean.
+Definition 19 (Tropical Fermat–Weber Points (Lin & Yoshida,, 2018)).
+Let $X=\{\mathbf{x}_{1},\dots,\mathbf{x}_{K}\}\subset\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ be our dataset. A solution to the following optimization problem is a tropical Fermat–Weber point:
 $\min_{\mathbf{t}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}}f(\mathbf{t})=\frac{1}{K}\sum_{k\leq K}d_{{\rm tr}}(\mathbf{x}_{k},\mathbf{t}).$
-Definition 18 (Tropical Fréchet means (Jazlan et al., (2024+); Monod et al., (2022))).
-Let $X=\{\mathbf{x}_{1},\dots,\mathbf{x}_{K}\}\subset\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ be our dataset. A solution to the following optimisation problem is a tropical Fréchet mean:
+Definition 20 (Tropical Fréchet means (Monod et al.,, 2018)).
+Let $X=\{\mathbf{x}_{1},\dots,\mathbf{x}_{K}\}\subset\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ be our dataset. A solution to the following optimization problem is a tropical Fréchet mean:
 $\min_{\mathbf{t}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}}f(\mathbf{t})=\left[\frac{1}{K}\sum_{k\leq K}d_{{\rm tr}}(\mathbf{x}_{k},\mathbf{t})^{2}\right]^{1/2}.$
-The Fermat–Weber and Fréchet mean problems are both max-tropical and min-tropical location problems, as the tropical metric $d_{\mathrm{tr}}$ is both $\triangle_{\min}$ and $\triangle_{\max}$-star-quasiconvex.
-In fact, they are both classically convex and hence act as a benchmark against which we will test our gradient methods.
+The Fermat–Weber and Fréchet mean problems are both max-tropical and min-tropical location problems, as the tropical metric $d_{\mathrm{tr}}$ is both $\triangle_{\min}$ and $\triangle_{\max}$-star-quasi-convex.
+In fact, they are also both classically convex and hence act as a benchmark against which we can test our gradient methods.
 2.3.2 Wasserstein Projections
-The next statistical optimisation problem was motivated by Cai & Lim, (2022), and looks to identify a tropical projection which minimises the Wasserstein distance between samples in different tropical projective torii.
-Definition 19 (Tropical Wasserstein Projections (Talbut et al., (2023))).
-Let $X=\{\mathbf{x}_{1},\dots,\mathbf{x}_{K}\}\subset\mathbb{R}^{N}/\mathbb{R}\mathbf{1},Y=\{\mathbf{y}_{1},\dots,\mathbf{y}_{K}\}\subset\mathbb{R}^{M}/\mathbb{R}\mathbf{1}$ be our datasets.
+The next statistical optimization problem was motivated by Cai & Lim, (2022), and looks to identify a tropical projection which minimizes the Wasserstein distance between samples in different tropical projective tori.
+Definition 21 (Tropical Wasserstein Projections (Talbut et al.,, 2023)).
+Let $X=\{\mathbf{x}_{1},\dots,\mathbf{x}_{K}\}\subset\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ be a dataset in one space, and $Y=\{\mathbf{y}_{1},\dots,\mathbf{y}_{K}\}\subset\mathbb{R}^{M}/\mathbb{R}\mathbf{1}$ be a dataset in a different space.
 Let $J_{1},\dots,J_{M}$ be a non-empty partition of $[N]$.
-Then the tropical $p$-Wasserstein projection problem is the minimisation of the following objective function:
-$\min_{\mathbf{t}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}}f_{p}(\mathbf{t})=\left(\frac{1}{K}\sum_{k\leq K}\|[\max_{i\in J_{j}}(\mathbf{x}_{k}+\mathbf{t})_{i}]_{j}-\mathbf{y}_{k}\|_{\mathrm{tr}}^{p}\right)^{1/p}$
-The tropical $\infty$-Wasserstein projection problem is the minimisation of:
-$\min_{\mathbf{t}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}}f_{\infty}(\mathbf{t})=\max_{k\leq K}\|[\max_{i\in J_{j}}(\mathbf{x}_{k}+\mathbf{t})_{i}]_{j}-\mathbf{y}_{k}\|_{\mathrm{tr}}$
+Then the tropical $p$-Wasserstein projection problem is the minimization of the following objective function:
+$\min_{\mathbf{t}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}}f_{p}(\mathbf{t})=\left(\frac{1}{K}\sum_{k\leq K}\|[\max_{i\in J_{j}}(\mathbf{x}_{k}-\mathbf{t})_{i}]_{j}-\mathbf{y}_{k}\|_{\mathrm{tr}}^{p}\right)^{1/p}$
+The tropical $\infty$-Wasserstein projection problem is the minimization of:
+$\min_{\mathbf{t}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}}f_{\infty}(\mathbf{t})=\max_{k\leq K}\|[\max_{i\in J_{j}}(\mathbf{x}_{k}-\mathbf{t})_{i}]_{j}-\mathbf{y}_{k}\|_{\mathrm{tr}}$
 In contrast to the work by Lee et al., (2022) which looks to solve the optimal transport problem for intrinsic measures on $\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$, this $p$-Wasserstein projection problem is computing an optimal projection for empirical samples.
 Proposition 22.
-The tropical $p,\infty$-Wasserstein projections are max-tropical location problems, and the $\infty$-Wasserstein projection has $\max$-tropically convex sublevel sets.
-Proof.
-Let $\mathbf{z}_{k}=(x_{ki}-y_{kj})_{i\leq N}$ where $j$ is such that $i\in J_{j}$. We then have that
-$\|[\max_{i\in J_{j}}(\mathbf{x}_{k}+\mathbf{t})_{i}]_{j}-\mathbf{y}_{k}\|_{\mathrm{tr}}$ $=\max_{l}[t_{l}+z_{kl}]+\max_{j}[\min_{i\in J_{j}}-t_{i}-z_{ki}]$
-$=\max_{j}[\min_{i\in J_{j}}[\max_{l}[t_{l}+z_{kl}]-t_{i}-z_{ki}]]$
+The tropical $p,\infty$-Wasserstein projections are min-tropical location problems, and the $\infty$-Wasserstein projection has $\min$-tropically convex sublevel sets.
+Proof Let $\mathbf{z}_{k}=(x_{ki}-y_{kj})_{i\leq N}$ where $j$ is such that $J_{j}$ is the unique partition set containing $i$. We then have that
+$\|[\max_{i\in J_{j}}(\mathbf{x}_{k}-\mathbf{t})_{i}]_{j}-\mathbf{y}_{j}\|_{\mathrm{tr}}$ $=\max_{b}\max_{a\in J_{b}}[z_{ka}-t_{a}]-\min_{j}\max_{i\in J_{j}}[z_{ki}-t_{i}]$
+$=\max_{a}[z_{ka}-t_{a}]+\max_{j}\min_{i\in J_{j}}[t_{i}-z_{ki}]$
+$=\max_{j}[\min_{i\in J_{j}}[t_{i}-z_{ki}-\min_{a}[t_{a}-z_{ka}]]]$
 The $p,\infty$-Wasserstein projections can then be written as
-$f_{p}(\mathbf{t})$ $=\left(\frac{1}{K}\sum_{k\leq K}\left(\max_{j}[\min_{i\in J_{j}}[\max_{l}[t_{l}+z_{kl}]-t_{i}-z_{ki}]]\right)^{p}\right)^{1/p}$
-$f_{\infty}(\mathbf{t})$ $=\max_{k\leq K}\max_{j}[\min_{i\in J_{j}}[\max_{l}[t_{l}+z_{kl}]-t_{i}-z_{ki}]]$
+$f_{p}(\mathbf{t})$ $=\left(\frac{1}{K}\sum_{k\leq K}\left(\max_{j}[\min_{i\in J_{j}}[t_{i}-z_{ki}-\min_{a}[t_{a}-z_{ka}]]]\right)^{p}\right)^{1/p}$
+$f_{\infty}(\mathbf{t})$ $=\max_{k\leq K}\max_{j}[\min_{i\in J_{j}}[t_{i}-z_{ki}-\min_{a}[t_{a}-z_{ka}]]]$
 These are max-tropical location problems with:
@@
 $g_{\infty}(\mathbf{h})$ $=\max_{k\leq K}h_{k},$
-$h_{k}(\mathbf{t})$ $=\hat{\gamma}(\mathbf{t}+\mathbf{z}_{k}),$
+$h_{k}(\mathbf{t})$ $=\hat{\gamma}(\mathbf{t}-\mathbf{z}_{k}),$
 $\gamma(\mathbf{t})$ $=\max_{j}\min_{i\in J_{j}}t_{i}.$
-The $g_{p},g_{\infty}$ and $\gamma$ are increasing and each $h_{k}$ has kernel $-z_{k}$, satisfying the conditions for a max-tropical location problem.
-As the $h_{k}$ are $\triangle_{\max}$-star-convex, their sublevel sets are max-tropically convex.
-This is preserved when taking a maximum over samples, so $f_{\infty}$ also has $\max$-tropically convex sublevel sets.
+The $g_{p},g_{\infty}$ and $\gamma$ are increasing in each coordinate and each $h_{k}$ has kernel $\mathbf{z}_{k}$, satisfying the conditions for a min-tropical location problem.
+As the $h_{k}$ are $\triangle_{\min}$-star-convex, their sublevel sets are min-tropically convex by Proposition 15.
+This is preserved when taking a maximum over samples, so $f_{\infty}$ also has $\min$-tropically convex sublevel sets.
 ∎
 Figure 4: A heat map of a 3-sample $\infty$-Wasserstein projection objective on $\mathbb{R}^{3}/\mathbb{R}\mathbf{1}\cong\mathcal{H}=\{\mathbf{x}:\sum x_{i}=0\}$.
+Figure 4 shows a heat map for the $\infty$-Wasserstein projection objectives. As in the tropical linear regression problem, it is piecewise linear, non-convex and has sparse gradient.
 3 Steepest Descent with the Tropical Norm
 This section comprises the theoretical contributions of this paper; we outline our proposed method of tropical descent, and the theoretical convergence guarantees for tropical descent when applied to tropical location problems.
-We treat optimisation over $\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ as optimisation over $\mathbb{R}^{N}$ with the knowledge that $f$ is constant along $\mathbf{1}$-rays.
-We assume that for all $\mathbf{t}$, we can compute some gradient $\nabla f(\mathbf{t})$ which is contained in the generalised Clarke derivative of $\mathbf{t}$.
-Definition 21 (Generalised Clarke Derivative (2.13, I.I of Kiwiel, (2006))).
+We treat optimization over $\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ as optimization over $\mathbb{R}^{N}$ with the knowledge that $f$ is constant along $\mathbf{1}$-rays.
+We assume we can compute some gradient-like function $\nabla f:\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}$ which is satisfying Assumption 23.
+Assumption 23.
+Let $f=g(h_{1},\dots,h_{K})$ be a min-tropical location problem. Then we have that
+$\nabla f(\mathbf{t})_{j}<0\Rightarrow\exists\,k\text{ s.t. }j\in\text{argmin}_{i}(t_{i}-x_{ki}).$
+This is a very general assumption, which holds for some of the most natural gradient-like functions.
+For example, if $\nabla f(\mathbf{t})$ is given by some Clarke gradient, that is, some element of the generalized Clarke derivative of $f$, then Lemma 25 shows that Assumption 23 holds.
+Definition 24 (Generalized Clarke Derivative (2.13, I.I of Kiwiel, (1985))).
 We define the set $M_{f}(\mathbf{t})$ by
 $M_{f}(\mathbf{t})=\{\mathbf{z}\in\mathbb{R}^{N}:\nabla f(\mathbf{y}_{l})\rightarrow\mathbf{z},\text{ where }\mathbf{y}_{l}\rightarrow\mathbf{t},f\text{ differentiable at each }\mathbf{y}_{l}\}.$
-Then the generalised Clarke derivative is given by
+Then the generalized Clarke derivative is given by
 $\partial f(\mathbf{t})=\operatorname{conv}M_{f}(\mathbf{t}).$
+Lemma 25.
+Suppose $f=g(h_{1},\dots,h_{K})$ is min-tropical location problem, and for all $\mathbf{t}$, $\nabla f(\mathbf{t})$ is some element of $\partial f(\mathbf{t})$. Then Assumption 23 holds.
+Proof Suppose that $\nabla f(\mathbf{t})_{j}<0$.
+Then there is a sequence of differentiable $\mathbf{y}_{\ell}$ satisfying $\nabla f(\mathbf{y}_{\ell})_{j}<0$, $\mathbf{y}_{\ell}\rightarrow\mathbf{t}$.
+Hence by the chain rule, as $g$ is increasing in every $h_{k}$, there is some $h_{k}$ which is strictly decreasing in the $j^{th}$ coordinate at $\mathbf{y}_{\ell}$.
+WLOG, the same $h_{k}$ is strictly decreasing in the $j^{th}$ coordinate for all $\mathbf{y}_{\ell}$.
+As $\gamma$ is increasing in every coordinate, $h_{k}=\hat{\gamma}(\mathbf{y}_{\ell}-\mathbf{x}_{k})$ is strictly decreasing in the $j^{th}$ coordinate if and only if $j\in\text{argmin}_{i}(y_{\ell i}-x_{ki})$.
+This is preserved through the limit as $\ell\rightarrow\infty$, so we conclude that $j\in\text{argmin}_{i}[t_{i}-x_{i}]$.
+∎
+In practice, rather than computing elements of the generalized Clarke derivative, we use PyTorch automatic differentiation (Paszke et al.,, 2019).
+The derivatives computed by PyTorch are dependent on the computational representation of a function, but are designed to respect the chain rule.
+We must therefore ensure that our computational representation of $g$ has non-negative derivatives in each $h_{k}$, while each $h_{k}$ is computed such that its $t_{j}$ derivative is negative only if $i\in\text{argmin}_{i}(t_{i}-x_{ki})$.
+For such a computational representation of $f$, Assumption 23 will hold by the chain rule.
+For all the tropical location problems we consider in this paper, we find their most natural computational representation gives rise to derivatives $\nabla f$ satisfying Assumption 23.
 3.1 Proposed Method
-In defining tropical descent, we refer back to the motivation behind classical gradient descent; that is, the steepest descent direction.
-Definition 22 (Steepest Descent Direction (Boyd & Vandenberghe, (2004))).
+In defining tropical descent, we refer back to the motivation behind classical gradient descent—steepest descent directions.
+Definition 26 (Steepest Descent Direction (Boyd & Vandenberghe,, 2004)).
 […]
-$\mathbf{d}=\operatorname{argmax}\{\nabla f(\mathbf{t})^{\top}\mathbf{v}:\|\mathbf{v}\|=1\}.$
-With respect to the Euclidean norm, the normalised gradient is a steepest descent direction. When we use a the tropical norm, the steepest descent directions are characterised by the following lemma.
+$\mathbf{d}=\text{argmin}\{\nabla f(\mathbf{t})^{\top}\mathbf{v}:\|\mathbf{v}\|=1\}.$
+With respect to the Euclidean norm, the normalized derivative is a steepest descent direction.
+When we use a the tropical norm, the steepest descent directions are characterized by the following lemma.
 Lemma 27.
-Suppose $\nabla f\neq 0$. Then direction $\mathbf{v}$ is a normalised tropical steepest descent direction iff
+Suppose $\nabla f\neq\mathbf{0}$. Then the direction $\mathbf{v}$ is a normalized tropical steepest descent direction iff
 $v_{i}=1$ $\Leftrightarrow\nabla f_{i}<0$
 $v_{i}=0$ $\Leftrightarrow\nabla f_{i}>0$
-Proof.
-WLOG, we assume $0\leq v_{i}\leq 1$. Then
-$\nabla f^{\top}\mathbf{v}\leq\sum_{\nabla f_{i}>0}\nabla f_{i}$
-This is achieved if and only if $v_{i}=1$ for all $i$ such that $\nabla f_{i}>0$, and $v_{i}=0$ for all $i$ such that $\nabla f_{i}<0$.
-We note that any such $v_{i}$ satisfies $\|\mathbf{v}\|_{{\rm tr}}=1$ as $\sum\nabla f_{i}=0,\nabla f_{i}\neq 0$, and hence $\nabla f_{i}$ has positive and negative coordinates.
+Proof WLOG, we assume $0\leq v_{i}\leq 1$. Then
+$\nabla f^{\top}\mathbf{v}\geq\sum_{i:\,\nabla f_{i}<0}\nabla f_{i}$
+This is achieved if and only if $v_{i}=1$ for all $i$ such that $\nabla f_{i}<0$, and $v_{i}=0$ for all $i$ such that $\nabla f_{i}>0$.
+We note that any such $v_{i}$ satisfies $\|\mathbf{v}\|_{{\rm tr}}=1$ as $\sum_{i}\nabla f_{i}=0$ by the equivalence relation of the tropical projective torus, and $\nabla f_{i}\neq\mathbf{0}$ so $\nabla f$ has positive and negative coordinates.
 ∎
-This lemma uniquely defines $\mathbf{d}$ if and only if, for every coordinate $i$, $\nabla f_{i}\neq 0$.
-To uniquely define a descent direction for any $\mathbf{x}$, we take the direction $\mathbf{d}$ given by
-$d_{i}=\begin{cases}1&\text{ if }\nabla f_{i}>0,\\ 0&\text{ otherwise.}\end{cases}$
-We refer to this as a positive descent direction. In contrast,
-$d_{i}=\begin{cases}-1&\text{ if }\nabla f_{i}<0,\\ 0&\text{ otherwise.}\end{cases}$
-defines a negative descent direction.
-Throughout the rest of this paper, we use the positive descent direction as convention, and will specify when the negative descent direction is used instead.
-Tropical descent therefore refers to the method outlined by Algorithm 1, for some specified step size rule, while we refer to classical gradient descent as classical descent.
+Lemma 27 uniquely defines a descent direction $\mathbf{d}$ if and only if, for every coordinate $i$, $\nabla f_{i}\neq 0$.
+To uniquely define a descent direction for any $\mathbf{t}$, we take the direction $\mathbf{d}$ given by
+$d_{i}=\begin{cases}1&\text{ if }\nabla f_{i}<0,\\ 0&\text{ otherwise.}\end{cases}$
+We refer to this as a min-tropical descent direction. In contrast,
+$d_{i}=\begin{cases}-1&\text{ if }\nabla f_{i}>0,\\ 0&\text{ otherwise.}\end{cases}$
+defines a max-tropical descent direction.
+Throughout the rest of this paper, we use the min-tropical descent direction as convention, and will specify when the max-tropical descent direction is used instead.
+Tropical descent therefore refers to the method outlined by Algorithm 1, for some specified step size sequence $(a_{n})_{n\geq 1}$, while we refer to classical gradient descent as classical descent.
 Algorithm 1 Tropical Descent
 for $n<$ max_steps do
-     $\mathbf{g}_{n}\leftarrow\nabla f(\mathbf{t}_{n-1})$;
-     $\mathbf{d}_{n}\leftarrow[\mathbf{g}_{n}>0]_{i\leq N}$;
-     $\mathbf{t}_{n}\leftarrow\mathbf{t}_{n-1}-a_{n}\mathbf{d}_{n}$;
+     $\mathbf{g}_{ni}\leftarrow\nabla f_{i}(\mathbf{t}_{n-1})$;
+     $\mathbf{d}_{ni}\leftarrow\mathbf{1}_{\mathbf{g}_{ni}<0}$;
+     $\mathbf{t}_{n}\leftarrow\mathbf{t}_{n-1}+a_{n}\mathbf{d}_{n}$;
+We note that over a large number of steps, Algorithm 1 can produce large $t_{ni}$ values as each coordinate is non-decreasing in $n$.
+If this results in overflow or precision errors, a normalisation operation can be included in each step, such as subtracting $\frac{1}{N}\sum_{i}t_{ni}$ from each $t_{ni}$.
+The function $f$ is invariant under such a translation due to the equivalence relation on the tropical projective torus.
+In (Barnhill et al.,, 2024a, b), classical descent is implemented for the computation of Fermat–Weber points; this is the most natural choice as the Fermat–Weber problem is classically convex.
+However, the linear regression and Wasserstein projection problems lack classical convexity, and hence do not consistently produce accurate solutions via classical descent.
+It is these tropically convex problems for which tropical descent is necessary.
 3.2 Convergence Guarantees
-In this section we prove the theoretical guarantees of tropical descent for max-tropical location problems.
-We show that tropical descent must converge to the tropical convex hull of the data, and hence will necessarily find a minimum of a $\triangle_{\max}$-star-convex problem.
+In this section we prove the theoretical guarantees of tropical descent for min-tropical location problems.
+We show that tropical descent must converge to the tropical convex hull of the data, and hence will necessarily find a minimum of a $\triangle_{\min}$-star-convex problem.
 Throughout this section, we assume the step sizes $a_{n}$ satisfy $a_{n}\rightarrow 0,\sum a_{n}=\infty$.
-We first highlight the key difference in stability between classical descent and tropical descent; if some component of the gradient is locally zero, tropical descent cannot converge in that neighbourhood.
+We first highlight the key difference in stability between classical descent and tropical descent; if some component of the derivative is locally zero, tropical descent cannot converge in that neighbourhood.
 Proposition 28.
+Let $\mathbf{t}_{n}$ be some sequence defined by tropical descent.
 Assume that for all $n$, $\nabla f(\mathbf{t}_{n})\neq\mathbf{0}$.
 Suppose in some bounded open $U$, $\nabla f_{i}$ is identically 0.
-Then no point in $\operatorname{int}U$ is a convergence point for $\mathbf{t}_{n}$; in particular, if $\mathbf{t}_{m}\in U$ we will eventually leave $U$.
-Proof of Proposition 24.
-Assume otherwise. Then for all $n\geq m$, we have $\nabla f(\mathbf{t}_{n})_{i}=0$, so $t_{ni}$ is constant. At each step, we decrease at least one coordinate by $a_{n}$, so as $n\rightarrow\infty$:
-$\sum_{j}t_{nj}\leq\sum_{j}t_{mj}-\sum_{n\geq m}a_{n}\rightarrow-\infty.$
-Therefore $\|\mathbf{t}_{n}\|_{{\rm tr}}$ will become arbitrarily large, as $\min_{j}t_{nj}\rightarrow-\infty$ while $\max_{j}t_{nj}\geq t_{ni}$. Then we must have left $U$; contradiction. ∎
-As a result, local minima such as the valleys in figures 4 and 3 can be stable with respect to classical descent but unstable with respect to tropical descent.
-We next prove an opposing result; tropical gradient descent must converge to the tropical convex hull of the data points.
-This result stems from the following observation on the gradient of $\triangle_{\max}$-star-quasi-convex functions.
-Remark 25.
-We note that for a $\triangle_{\max}$-star-quasi-convex function with kernel $\mathbf{v}$, we have that
-$\nabla f(\mathbf{x})_{j}>0\Rightarrow j\in\operatorname{argmax}_{i}(x_{i}-v_{i}).$
+Then no point in $\operatorname{int}U$ is a convergence point for $\mathbf{t}_{n}$; in particular, if $\mathbf{t}_{m}\in U$ there is some $n>m$ such that $\mathbf{t}_{n}\notin U$.
+Proof Assume otherwise.
+Then for all $n\geq m$, we have $\nabla f(\mathbf{t}_{n})_{i}=0$, so $t_{ni}$ is constant.
+At each step, we increase at least one coordinate by $a_{n}$, so as $n\rightarrow\infty$:
+$\sum_{j}t_{nj}\geq\sum_{j}t_{mj}+\sum_{n\geq m}a_{n}\rightarrow\infty.$
+Therefore $\|\mathbf{t}_{n}\|_{{\rm tr}}$ will become arbitrarily large, as $\max_{j}t_{nj}\rightarrow\infty$ while $\min_{j}t_{nj}\leq t_{mi}$. Then we must have left $U$; contradiction. ∎
+As a result of the proposition above, local minima such as the valleys in Figures 3 and 4 can be stable with respect to classical descent but unstable with respect to tropical descent.
+We next prove the main convergence result of this work; tropical descent must converge to the tropical convex hull of the data points.
 Theorem 29.
-Suppose $f$ is a max-tropical location problem with respect to the dataset $X=\{\mathbf{x}_{1},\dots,\mathbf{x}_{K}\}$.
+Suppose $f$ is a min-tropical location problem with respect to the dataset $X=\{\mathbf{x}_{1},\dots,\mathbf{x}_{K}\}$.
 Let $(\mathbf{t}_{n})_{n\geq 1}$ be a sequence of points defined by tropical descent, such that for all $n$: $\nabla f(\mathbf{t}_{n})\neq 0$.
-Let $V=\operatorname{tconv}_{\min}(\mathbf{x}_{1},\dots,\mathbf{x}_{K})$, and define
-$\Delta_{i}(\mathbf{t})=\sum_{k\in[K]}\sum_{j\in[N]}[t_{j}-x_{kj}-t_{i}+x_{ki}]^{+}.$
-For any $\epsilon>0$, let $M_{1}$ be such that for all $m\geq M_{1}$, $a_{m}\leq\epsilon/N$.
-Let $M_{2}$ be such that for all $m\geq M_{2}$ $s_{m}\geq\max_{i}\Delta_{i}(\mathbf{t}_{M_{1}})+s_{M_{1}}-K\epsilon$.
-Then almost surely, for all $m\geq M_{2}$ we have:
+Let $V=\operatorname{tconv}_{\max}(\mathbf{x}_{1},\dots,\mathbf{x}_{K})$, and define
+$\Delta_{j}(\mathbf{t})=\sum_{k\in[K]}\sum_{i\in[N]}[t_{j}-x_{kj}-t_{i}+x_{ki}]^{+}.$
+For any $\epsilon>0$, let $M_{1}$ be such that for all $m\geq M_{1}$, $a_{m}<\epsilon/N$.
+Let $M_{2}$ be such that for all $m\geq M_{2}$ $s_{m}\geq\max_{i}\Delta_{i}(\mathbf{t}_{M_{1}})+s_{M_{1}}-K\epsilon$, where $s_{m}$ is the partial sum of the step size sequence $a_{m}$.
+Then for all $m\geq M_{2}$ we have:
 $d_{{\rm tr}}(\mathbf{t}_{n},V)\leq\epsilon.$
-Proof.
-For each $i\leq N$, we define the functions
-$\delta_{i}(\mathbf{t})=\min_{k\in[K]}[\sum_{j}[t_{j}-v_{kj}-t_{i}+v_{ki}]^{+}].$
+Proof For each $i\leq N$, we define the functions
+$\delta_{j}(\mathbf{t})=\min_{k\in[K]}\left[\sum_{i\in[N]}[t_{j}-x_{kj}-t_{i}+x_{ki}]^{+}\right].$
 Claim 30.
-$d_{{\rm tr}}(\mathbf{t}_{m},V)\leq\max_{j}\delta_{j}(\mathbf{t}_{m})$.
-Proof.
-As discussed in Maclagan & Sturmfels, (2015), we have the projection map onto $V$ given by:
-$\pi_{V}(\mathbf{t})=\lambda_{1}\odot\mathbf{x}_{1}\oplus\dots\oplus\lambda_{K}\odot\mathbf{x}_{K},\qquad\text{ where }\lambda_{k}=\max_{i}(t_{i}-x_{ki})$
-So the distance to $V$ is given by $d(\mathbf{t},\pi_{V}(\mathbf{t}))$, and we have:
+For all $\mathbf{t}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$, we have $d_{{\rm tr}}(\mathbf{t},V)\leq\max_{j}\delta_{j}(\mathbf{t})$.
+Proof As discussed by Maclagan & Sturmfels, (2015), there is a well-defined metric projection map $\pi_{V}$ onto $V$ given by:
+$\pi_{V}(\mathbf{t})=\lambda_{1}\odot\mathbf{x}_{1}\boxplus\dots\boxplus\lambda_{K}\odot\mathbf{x}_{K},$
+[…]
+The distance from $\mathbf{t}$ to $V$ is then given by $d(\mathbf{t},\pi_{V}(\mathbf{t}))$, and we have:
 $d(\mathbf{t},V)$ $=d(\mathbf{t},\pi_{V}(t))$
-$=\|(\min_{k}(\max_{i}(t_{i}-x_{ki})+x_{kj}-t_{j}))_{j}\|$
-$\leq\max_{j}(\min_{k}(\max_{i}(t_{i}-x_{ki})+x_{kj}-t_{j}))$
-$\leq\max_{j}\delta_{j}(\mathbf{t})$
+$=\|(\max_{k}(\min_{i}(t_{i}-x_{ki})-t_{j}+x_{kj})_{j\leq N}\|.$
+We note that for all $j,k$, the term $\min_{i}(t_{i}-x_{ki})-t_{j}+x_{kj}$ is negative. Hence:
+$d(\mathbf{t},V)$ $\leq 0-\min_{j}(\max_{k}(\min_{i}(t_{i}-x_{ki})-t_{j}+x_{kj}))$
+$\leq\max_{j}(\min_{k}(t_{j}-x_{kj}-\min_{i}(t_{i}-x_{ki})))$
+$\leq\max_{j}\delta_{j}(\mathbf{t}).$
+The claim is proved. ∎
+We will now prove that for all $m\geq M_{2}$, we have that $\delta_{j}(\mathbf{t}_{m})\leq\epsilon$.
 Claim 31.
 If $\delta_{i}(\mathbf{t}_{m})=0$ and $m\geq M_{1}$, then $\delta_{i}(\mathbf{t}_{m+1})<\epsilon$.
-Proof.
-For all $k\in[K],j\in[N]$, we have that $t_{j}-v_{kj}-t_{i}+v_{ki}$ is $1$-Lipschitz in $\mathbf{t}$.
-Therefore $\delta_{i}$ is $N$- Lipschitz in $\mathbf{t}$, and if $\delta_{i}(\mathbf{t}_{m})=0,\delta_{i}(\mathbf{t}_{m+1})\leq Na_{m}<\epsilon$.
+Proof For all $k\in[K],j\in[N]$, we have that $t_{mj}-x_{kj}-t_{mi}+x_{ki}$ is $1$-Lipschitz in $\mathbf{t}$ (with respect to $d_{\mathrm{tr}}$).
+Therefore $\delta_{j}$ is $N$-Lipschitz in $\mathbf{t}$.
+By the tropical descent rule:
+$\delta_{j}(t_{m+1})$ $\leq\delta_{j}(\mathbf{t}_{m})+Nd_{\mathrm{tr}}(\mathbf{t}_{m},\mathbf{t}_{m+1})$
+$=0+Na_{m}$
+$<\epsilon.$
+The claim is proved. ∎
+Claim 32.
+If $\delta_{j}(\mathbf{t}_{m})>0$ then $\delta_{j}(\mathbf{t}_{m+1})\leq\delta_{j}(\mathbf{t}_{m})$.
+Proof Suppose $\delta_{j}(\mathbf{t}_{m})>0$.
+Then $\forall k$, $t_{mj}-x_{kj}>\min_{i}(t_{mi}-x_{ki})$, so by Lemma 25 we have $\nabla f_{j}\geq 0$.
+The tropical steepest descent direction therefore has $d_{mj}=0$.
+We conclude that
+$\forall\,k,i:\quad$ $t_{(m+1)j}-x_{kj}-t_{(m+1)i}+x_{ki}\leq t_{mj}-x_{kj}-t_{mi}+x_{ki},$
+$\Rightarrow\quad\delta_{j}(\mathbf{t}_{m+1})\leq\delta_{j}(\mathbf{t}_{m}).$
+The claim is proved. ∎
+Claim 33.
+For some $M_{1}\leq m\leq M_{2}$, $\delta_{j}(\mathbf{t}_{m})<\epsilon$.
+Proof We note that $K\delta_{j}(\mathbf{t})\leq\Delta_{j}(\mathbf{t})$; hence if $\Delta_{j}(\mathbf{t}_{M_{1}})<K\epsilon$ then we are done.
+We assume $\delta_{j}(\mathbf{t}_{m})>\epsilon$ for all $M_{1}\leq m\leq M_{2}$, and prove the result by contradiction.
+Then as in the proof of the previous claim, at each step we have $\nabla f_{j}\geq 0$, $d_{mj}=0$ and $t_{mj}$ is fixed for $M_{1}\leq m\leq M_{2}$.
+We have assumed that for each $m$, $\nabla f\neq\mathbf{0}$ and so there is some $i$ such that $\nabla f_{i}<0$, and by Lemma 25, there is some $k$ such that $i\in\text{argmin}_{\ell}t_{m\ell}-v_{k\ell}$.
+Hence:
+$\forall\,\ell:\quad t_{mj}-v_{kj}-t_{mi}+v_{ki}$ $\geq t_{mj}-v_{kj}-t_{m\ell}+v_{k\ell},$
+$\Rightarrow\quad t_{mj}-v_{kj}-t_{mi}+v_{ki}$ $\geq\delta_{j}(\mathbf{t}_{m})/N,$
+$>\epsilon/N,$
+$\geq a_{m}.$
+Hence this term in the summation of $\Delta_{j}(\mathbf{t}_{m})$ decreases by $a_{m}$, while other terms are non-increasing as $t_{mj}$ is fixed.
+Noting that $K\delta_{j}(\mathbf{t})\leq\Delta_{j}(\mathbf{t})$, we conclude
+$K\delta_{j}(\mathbf{t}_{M_{2}})$ $\leq\Delta_{j}(\mathbf{t}_{M_{2}}),$
+$\leq\Delta_{j}(\mathbf{t}_{M_{1}})-s_{M_{2}}+s_{M_{1}},$
+$\leq K\epsilon.$
+Contradiction. So the claim is proved. ∎
+Finally, from the claims above, we conclude that $\forall m\geq M_{2},i\in[N]$, we have $\delta_{i}(\mathbf{t}_{m})<\epsilon$, and so by our first claim, we have $d_{{\rm tr}}(\mathbf{t}_{m},V)\leq\epsilon$.
 ∎
-Claim 29.
-If $\delta_{i}(\mathbf{t}_{m})>0$ then $\delta_{i}(\mathbf{t}_{m+1})\leq\delta_{i}(\mathbf{t}_{m})$.
-Proof.
-Suppose $\delta_{i}(\mathbf{t}_{m})>0$.
-Then $\forall k$, $\max_{j}(t_{j}-v_{kj})>t_{i}-v_{ki}$.
-By remark 25, $\nabla f_{i}\leq 0$ and $d_{mi}=0$.
-Therefore $\delta_{i}(\mathbf{t}_{m+1})\leq\delta_{i}(\mathbf{t}_{m})$.
-∎
-Claim 30.
-Let $M$ be such that $s_{M}\geq\Delta_{i}(\mathbf{t}_{M_{1}})+s_{M_{1}}-K\epsilon$. Then for some $M_{1}\leq m\leq M$, $\delta_{i}(\mathbf{t}_{m})<\epsilon$.
-Proof.
-We note that $K\delta_{i}(\mathbf{t})\leq\Delta_{i}(\mathbf{t})$; hence if $\Delta_{i}(\mathbf{t}_{M_{1}})<K\epsilon$ then we are done.
-We now assume $\Delta_{i}(\mathbf{t}_{M_{1}})\geq K\epsilon$ and prove the result by contradiction; suppose $\delta_{i}(\mathbf{t}_{m})>\epsilon$ for all $M_{1}\leq m\leq M$.
-Therefore at each step we have $\nabla f_{i}\leq 0$, $d_{mi}=0$ and $t_{mi}$ is fixed for $n\leq M$.
-As $\nabla f\neq\mathbf{0}$, there is some $j$ such that $\nabla f_{j}>0$, and let $k$ be such that $j=\operatorname{argmax}_{l}t_{ml}-v_{kl}$. As $\delta_{i}(\mathbf{t}_{m})>\epsilon$:
-$t_{j}-v_{kj}-t_{i}+v_{ki}>\epsilon/N\geq a_{m}.$
-Hence this term in the summation of $\Delta_{i}$ decreases by $a_{m}$, while other terms are non-increasing.
-We conclude that $\Delta_{i}(\mathbf{t}_{M})\leq\Delta_{i}(\mathbf{t}_{M_{1}})-s_{M}+s_{M_{1}}\leq K\epsilon$.
-Contradiction.
-∎
-Finally, from claims 28 to 30, we conclude that $\forall m\geq M_{2},i\leq N$, we have $\delta_{i}(\mathbf{t}_{m})<\epsilon$, and by claim 27, we have $d_{{\rm tr}}(\mathbf{t}_{m},V)\leq\epsilon$. ∎
-As proven in Comăneci, (2023), tropical location problems have minima in the tropical convex hull of the data.
-This result says that according to tropical gradient descent, only such minima can be stable.
-As a corollary, we can achieve the same result for min-tropical location problems by using negative tropical descent directions.
+As proven by Comăneci, (2024), tropical location problems have minima in the tropical convex hull of the data.
+The theorem above says that according to tropical descent, only such minima can be stable.
+We cannot do much better in terms of convergence results for general location problems; they can have disconnected sublevel sets so we would not expect gradient methods to find a global minimum.
+We would require some degree of convexity to guarantee global solvability, which is exactly what the following corollary gives us; tropical descent will necessarily converge to a minimum of any $\triangle_{\min}$-star-quasi-convex function.
 Corollary 34.
-Suppose $f$ is a min-tropical location problem with respect to the dataset $X=\{\mathbf{x}_{1},\dots,\mathbf{x}_{K}\}$.
-Let $(\mathbf{t}_{n})_{n\geq 1}$ be a sequence of points defined via negative tropical descent directions, such that for all $n$: $\nabla f(\mathbf{t}_{n})\neq 0$.
-Let $V=\operatorname{tconv}_{\max}(\mathbf{x}_{1},\dots,\mathbf{x}_{K})$, and define
-$\Delta^{\prime}_{j}(\mathbf{t})=\sum_{k\in[K]}\sum_{i\in[N]}[t_{j}-x_{kj}-t_{i}+x_{ki}]^{+}.$
+Let $f(\mathbf{t})$ be a $\triangle_{\min}$-star-quasi-convex function, and let $(\mathbf{t}_{n})_{n\geq 1}$ be a sequence of points defined by tropical descent such that for all $n$: $\nabla f(\mathbf{t}_{n})\neq 0$.
+Then there is some global minimum $\mathbf{t}^{*}$ such that for all $m\leq M_{2}$, $d_{{\rm tr}}(\mathbf{t}_{m},\mathbf{t}^{*})\leq\epsilon$ where $M_{2}$ is as in Theorem 29.
+As a further corollary, we note that we can achieve the same result for max-tropical location problems by using max-tropical descent directions.
+Corollary 35.
+Suppose $f$ is a max-tropical location problem with respect to the dataset $X=\{\mathbf{x}_{1},\dots,\mathbf{x}_{K}\}$.
+Let $(\mathbf{t}_{n})_{n\geq 1}$ be a sequence of points defined via max-tropical descent directions, such that for all $n$: $\nabla f(\mathbf{t}_{n})\neq 0$.
+Let $V=\operatorname{tconv}_{\min}(\mathbf{x}_{1},\dots,\mathbf{x}_{K})$, and define
+$\Delta^{\prime}_{i}(\mathbf{t})=\sum_{k\in[K]}\sum_{j\in[N]}[t_{j}-x_{kj}-t_{i}+x_{ki}]^{+}.$
 For any initial $\mathbf{t}_{0}$ and $\epsilon>0$, let $M_{1}$ be such that for all $m\geq M_{1}$, $a_{m}\leq\epsilon/N$.
 Let $M_{2}$ be such that for all $m\geq M_{2}$ $s_{m}\geq\max_{i}\Delta^{\prime}_{j}(\mathbf{t}_{M_{1}})+s_{M_{1}}-K\epsilon$.
-Then almost surely, for all $m\geq M_{2}$ we have:
+Then for all $m\geq M_{2}$ we have:
 $d_{{\rm tr}}(\mathbf{t}_{n},V)\leq\epsilon.$
-Despite following from the previous results, the following corollary may be considered the most powerful result of this paper; tropical descent will necessarily minimise any max-tropically quasi-convex function.
-Corollary 32.
-Let $f(\mathbf{t})$ be a max-tropically quasiconvex function, and let $(\mathbf{t}_{n})_{n\geq 1}$ be a sequence of points defined by tropical descent such that for all $n$: $\nabla f(\mathbf{t}_{n})\neq 0$.
-Then there is some global minimum $\mathbf{t}^{*}$ such that for all $m\leq M_{2}$, $d_{{\rm tr}}(\mathbf{t}_{m},\mathbf{t}^{*})\leq\epsilon$.
-We conclude by establishing the convergence rates of tropical descent in practice; we take step sizes given by $a_{m}=\alpha m^{-1/2}\|\nabla f(\mathbf{t}_{m})\|_{{\rm tr}}$, and assume gradients have bounded norm as in the the case of piecewise linear objective functions.
+Remark 36.
+The results above indicate convergence to $V$ or $\mathbf{t}^{*}$ in the tropical topology of $\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$, but $\mathbf{t}_{n}$ cannot not converge in the Euclidean topology of $\mathbb{R}^{N}$ due to the divergent partial sums of $\alpha_{n}$.
+We conclude by establishing the convergence rate of tropical descent in practice; we take step sizes given by $a_{m}=\alpha m^{-1/2}\|\nabla f(\mathbf{t}_{m})\|_{{\rm tr}}$, and assume derivatives have bounded tropical norm as is the case for piecewise linear 1-Lipschitz objective functions $h_{k}$.
 In this case, we prove convergence at a rate of $O(\sqrt{m})$.
 Proposition 37.
-Suppose $f$ is a max-tropical location problem as in Theorem 26.
-Assume that for all $\mathbf{t}_{m}$, $1/L\leq\|\nabla f(\mathbf{t}_{m})\|_{{\rm tr}}\leq 2$, and assume further there is some $D$ such that $\forall k\leq K,n\in\mathbb{N}:\,d_{{\rm tr}}(\mathbf{t}_{n},\mathbf{x}_{k})\leq D$.
+Suppose $f$ is a min-tropical location problem as in Theorem 29.
+Assume that there is some $L\in\mathbb{N}$ such that for all $\mathbf{t}_{m}$, $1/L\leq\|\nabla f(\mathbf{t}_{m})\|_{{\rm tr}}\leq 2$, and assume further there is some $D$ such that $\forall k\leq K,n\in\mathbb{N}:\,d_{{\rm tr}}(\mathbf{t}_{n},\mathbf{x}_{k})\leq D$.
 We define $a_{m}=\alpha m^{-1/2}\|\nabla f(\mathbf{t}_{m})\|_{{\rm tr}}$.
 Then for all $m$ satisfying
-$\sqrt{m+1}/2-KDNL(N-1)/8\alpha+\sqrt{(\sqrt{m+1}/2-KDNL(N-1)/8\alpha)^{2}+LKN-1}>\sqrt{2},$
+$1<\xi(m)\coloneqq\frac{\sqrt{m+1}}{2}-\frac{KDNL(N-1)}{8\alpha}+\sqrt{\left(\frac{\sqrt{m+1}}{2}-\frac{KDNL(N-1)}{8\alpha}\right)^{2}+LKN-1},$
 we have that:
 $d_{{\rm tr}}(\mathbf{t}_{m},V)\leq\frac{2\alpha N}{\sqrt{m+1}-KDNL(N-1)/4\alpha-1}=\epsilon_{m}.$
-In particular, if $f$ is max-tropically quasiconvex, then there is some global minimum $\mathbf{t}^{*}$ such that
+In particular, if $f$ is a $\triangle_{\min}$-star-quasi-convex function, then there is some global minimum $\mathbf{t}^{*}$ such that
 $d_{{\rm tr}}(\mathbf{t}_{m},\mathbf{t}^{*})\leq\frac{2\alpha N}{\sqrt{m+1}-DNL(N-1)/4\alpha-1}.$
-Proof.
-We fix $m$ and define:
-$M_{1}=\left\lfloor\left(\sqrt{m+1}/2-KDNL(N-1)/8\alpha+\sqrt{(\sqrt{m+1}/2-KDNL(N-1)/8\alpha)^{2}+LKN-1}\right)^{2}\right\rfloor.$
+Proof We fix $m$ and define:
+$M_{1}=\left\lfloor\xi(m)^{2}\right\rfloor.$
 We note that:
-$\sqrt{M_{1}}$ $\leq\sqrt{m+1}/2-KDNL(N-1)/8\alpha+\sqrt{(\sqrt{m+1}/2-KDNL(N-1)/8\alpha)^{2}+LKN-1},$
-$0\leq\sqrt{M_{1}}$ $\leq\frac{L}{2N}\left(\frac{N\sqrt{m+1}}{L}-\frac{KDN^{2}(N-1)}{4\alpha}+\sqrt{\left(\frac{N\sqrt{m+1}}{L}-\frac{KDN^{2}(N-1)}{4\alpha}\right)^{2}+\frac{4N^{2}(LKN-1)}{L^{2}}}\right).$
+$0\leq\sqrt{M_{1}}\leq\xi(m)$
 So $\sqrt{M_{1}}$ lies between the roots of the polynomial
@@
 Therefore:
-$0\geq\frac{N}{L}M_{1}+\left(\frac{KDN^{2}(N-1)}{4\alpha}-\frac{N\sqrt{m+1}}{L}\right)\sqrt{M_{1}}+\frac{N-LKN^{2}}{L}$
-$\frac{N\sqrt{M_{1}}\sqrt{M_{1}+1}}{L}\leq\left(\frac{N\sqrt{m+1}}{L}-\frac{KDN^{2}(N-1)}{4\alpha}\right)\sqrt{M_{1}}+KN^{2}$
-$KDN(N-1)/2+2\alpha L^{-1}(\sqrt{M_{1}+1}-\sqrt{m+1})\leq\frac{2\alpha KN}{\sqrt{M_{1}}}$
+$\frac{N}{L}(M_{1}+1)+\left(\frac{KDN^{2}(N-1)}{4\alpha}-\frac{N\sqrt{m+1}}{L}\right)\sqrt{M_{1}}-KN^{2}\leq 0,$
+$\frac{N\sqrt{M_{1}}\sqrt{M_{1}+1}}{L}+\frac{KDN^{2}(N-1)\sqrt{M_{1}}}{4\alpha}-\frac{N\sqrt{M_{1}}\sqrt{m+1}}{L}\leq KN^{2},$
+$\frac{KDN(N-1)}{2}+2\alpha L^{-1}(\sqrt{M_{1}+1}-\sqrt{m+1})\leq\frac{2\alpha KN}{\sqrt{M_{1}}}.$
 Claim 38.
-For all $n\geq M_{1}$, we have $a_{n}\leq\epsilon_{m}/N$.
-Proof.
-We have assumed that
-$\sqrt{m+1}/2-KDNL(N-1)/8\alpha+\sqrt{(\sqrt{m+1}/2-KDNL(N-1)/8\alpha)^{2}+LKN-1}>\sqrt{2}$
-Then:
-$M_{1}$ $\geq\left(\sqrt{m+1}/2-KDNL(N-1)/8\alpha+\sqrt{(\sqrt{m+1}/2-KDNL(N-1)/8\alpha)^{2}+LKN-1}\right)^{2}-1$
-$\geq\left(\sqrt{m+1}/2-KDNL(N-1)/8\alpha+\sqrt{(\sqrt{m+1}/2-KDNL(N-1)/8\alpha)^{2}+LKN-1}-1\right)^{2}$
-$\sqrt{M_{1}}$ $\geq\sqrt{m+1}/2-KDNL(N-1)/8\alpha+\sqrt{(\sqrt{m+1}/2-KDNL(N-1)/8\alpha)^{2}+LKN-1}-1$
-$\geq(\sqrt{m+1}-KDNL(N-1)/4\alpha)-1$
-Hence:
-$a_{n}$ $\leq\frac{2\alpha}{M_{1}^{1/2}}$
-$\leq\frac{2\alpha}{\sqrt{m+1}-KDNL(N-1)/4\alpha-1}=\epsilon_{m}/N$
+For all $n\geq M_{1}$ and $m$ such that $1<\xi(m)$, we have $a_{n}\leq\epsilon_{m}/N$.
+Proof We have assumed that $1<\xi(m)$, so:
+$M_{1}$ $\geq\xi(m)^{2}-1,$
+$\geq\left(\xi(m)-1\right)^{2},$
+$\sqrt{M_{1}}$ $\geq\xi(m)-1,\geq(\sqrt{m+1}-KDNL(N-1)/4\alpha)-1.$
+Hence, by the step size definition:
+$a_{n}$ $\leq\frac{2\alpha}{\sqrt{M_{1}}},$
+$\leq\frac{2\alpha}{\sqrt{m+1}-KDNL(N-1)/4\alpha-1}=\epsilon_{m}/N.$
+The claim is proved. ∎
 Claim 39.
-For all $n\geq m$, we have $\max_{i}\Delta_{i}(\mathbf{t}_{M_{1}})+s_{M_{1}}-s_{m}\leq K\epsilon_{m}$.
-Proof.
-We note that
-$\max_{i}\Delta_{i}(\mathbf{t})$ $\leq\sum_{i}\Delta_{i}(\mathbf{t})$
-$=\sum_{k\leq K}\sum_{i\leq j\leq N}|t_{j}-x_{kj}-t_{i}+x_{ki}|$
-$=\sum_{k\leq K}\sum_{i\leq j\leq N}d_{{\rm tr}}(\mathbf{t},\mathbf{x}_{k}),$
-and hence
-$\max_{i}\Delta_{i}(\mathbf{t}_{M_{1}})$ $\leq\frac{KDN(N-1)}{2}.$
+For all $n\geq m$, we have $\max_{j}\Delta_{j}(\mathbf{t}_{M_{1}})+s_{M_{1}}-s_{m}\leq K\epsilon_{m}$.
+Proof We note that
+$\max_{j}\Delta_{j}(\mathbf{t}_{M_{1}})$ $\leq\sum_{j}\Delta_{j}(\mathbf{t}_{M_{1}}),$
+$=\sum_{k\leq K}\sum_{i\leq j\leq N}|t_{M_{1}j}-x_{kj}-t_{M_{1}i}+x_{ki}|,$
+$\leq\sum_{k\leq K}\sum_{i\leq j\leq N}d_{{\rm tr}}(\mathbf{t}_{M_{1}},\mathbf{x}_{k}),$
+$\leq\frac{KDN(N-1)}{2}.$
 The partial sums are bounded by:
@@
 $\leq\sum_{n=M_{1}+1}^{m}\frac{-\alpha}{Ln^{1/2}}$
+$\leq\sum_{n=M_{1}+1}^{m}-2\alpha L^{-1}(\sqrt{n+1}-\sqrt{n})$
 $\leq-2\alpha L^{-1}(\sqrt{m+1}-\sqrt{M_{1}+1})$
 Therefore:
-$\max_{i}\Delta_{i}(\mathbf{t}_{M_{1}})+s_{M_{1}}-s_{m}$ $\leq KDN(N-1)/2-2\alpha L^{-1}(\sqrt{m+1}-\sqrt{M_{1}+1})$
-$\leq\frac{2\alpha KN}{\sqrt{M_{1}}}$
-$\leq K\epsilon_{m}$
-Hence by Theorem 26, we have almost surely that
-$d_{{\rm tr}}(\mathbf{t}_{m},V)\leq\epsilon_{m}$
-The bound for the tropically quasiconvex case follows from Corollary 32.
+$\max_{i}\Delta_{i}(\mathbf{t}_{M_{1}})+s_{M_{1}}-s_{m}$ $\leq KDN(N-1)/2+2\alpha L^{-1}(\sqrt{M_{1}+1}-\sqrt{m+1}),$
+$\leq\frac{2\alpha KN}{\sqrt{M_{1}}},$
+$\leq K\epsilon_{m}.$
+Hence by Theorem 29, we have that
+$d_{{\rm tr}}(\mathbf{t}_{m},V)\leq\epsilon_{m}.$
+The bound for the tropically quasi-convex case follows from Corollary 34. ∎
 4 Numerical Experiments
-In this section we compare practical performance of classical descent, tropical descent, SGD, tropical SGD, Adam, Adamax, and Trop Adamax, with a focus on their stability behaviour.
-The implementation for all these experiments is available in an anonymous GitHub repository at https://github.com/Roroast/TropicalGradDescent.
+In this section we compare the practical performance of classical descent, tropical descent, SGD, tropical SGD, Adam, Adamax, and TrAdamax, with a focus on their stability behavior.
+The implementation for all these experiments is available in the GitHub repository at https://github.com/Roroast/TropicalGradDescent.
 4.1 Methods
-Descent
-We will perform classical descent (CD) and tropical descent (TD) using equivalent same step size sequences $a_{n}$ for both methods, which we fix as
+We begin by outlining our implementation of steepest descent, stochastic gradient descent (SGD), and Adam as well as their tropical variants.
+We take a heuristic approach to the tropicalization of SGD and Adam, replacing derivatives with tropical steepest descent directions in each case.
+Steepest Descent. We will perform classical descent (CD) and tropical descent (TD) using equivalent step size sequences $a_{n}$ for both methods, which we fix as
 $a_{n}=\frac{\gamma\|\nabla f(\mathbf{t}_{n})\|}{\sqrt{n}},$
 […]
-Stochastic Gradient Descent
-Stochastic gradient descent is a variant of gradient descent which looks to improve computational speed by approximating $\nabla f(t)$; rather than computing the loss and its gradient at $\mathbf{t}$ with respect to the entire dataset $f(\mathbf{t};X)$, we evaluate the loss and its gradient with respect to a random sample $\mathbf{x}_{n}\in X$, $f(\mathbf{t};\mathbf{x}_{n})$.
+Stochastic Gradient Descent.
+Stochastic gradient descent is a variant of gradient descent which looks to improve computational speed by approximating $\nabla f(t)$; rather than computing the loss and its derivative at $\mathbf{t}$ with respect to the entire dataset $f(\mathbf{t};X)$, we evaluate the loss and its derivative with respect to a random sample $\mathbf{x}_{n}\in X$, $f(\mathbf{t};\mathbf{x}_{n})$.
 This is simplified with the notation $f_{n}(\mathbf{t})$.
@@
 […]
-Adam Variants
-The Adam algorithm (Kingma & Ba, (2017)) uses gradient momentum to compound consistent but small effects.
+Adam Variants.
+The Adam algorithm (Kingma & Ba,, 2015) uses gradient momentum to compound consistent but small effects.
 This is done by taking first and second moments of the gradient over all steps using exponentially decaying weights.
 The Adamax algorithm is a variant of the Adam algorithm motivated by the use of the $\infty$-moments rather than second moments.
-To adapt the Adam algorithm for the tropical setting, we take moments of an un-normalized tropical descent direction.
-When looking to take higher moments of $\mathbf{m}_{n}$ over $n$, the $L_{\infty}$ is the natural option for the tropical setting.
-We therefore look to use the Adamax recursion relation on the $\infty$-moment estimate $\mathbf{u}_{n}$:
-$\mathbf{u}_{n+1}=\max(\beta_{2}\mathbf{u}_{n},\mathbf{d}_{n}).$
-We therefore define TropAdamax as in Algorithm 2.
-Algorithm 2 TropAdamax
+Algorithm 2 TrAdamax
 $\mathbf{m},\mathbf{u}\leftarrow\mathbf{0}$;
 for $n<$ max_steps do
-     $\mathbf{g}_{n}\leftarrow\nabla f(\mathbf{t}_{n-1})$;
-     $\mathbf{d}_{n}\leftarrow[\mathbf{g}_{n}>0]*(\max(\mathbf{g}_{n})-\min(\mathbf{g}_{n}))$;
+     $\mathbf{g}_{ni}\leftarrow\nabla f_{i}(\mathbf{t}_{n-1})$;
+     $\mathbf{d}_{ni}\leftarrow(\max_{j}(\mathbf{g}_{nj})-\min_{j}(\mathbf{g}_{nj}))*\mathbf{1}_{\mathbf{g}_{ni}>0}$;
      $\mathbf{m}_{n}\leftarrow\beta_{1}*\mathbf{m}_{n-1}+(1-\beta_{1})*\mathbf{d}_{n}$;
@@
      $\mathbf{t}_{n}\leftarrow\mathbf{t}_{n}-\alpha/(1-\beta_{1}^{n})*\mathbf{m}_{n}/\mathbf{u}_{n}$;
+To adapt the Adam algorithm for the tropical setting, we take moments of an unnormalized tropical descent direction.
+When looking to take higher moments of the gradient over $n$, the $L_{\infty}$ is the natural option for the tropical setting.
+We therefore look to use the Adamax recursion relation on the $\infty$-moment estimate $\mathbf{u}_{n}$:
+$\mathbf{u}_{n+1}=\max(\beta_{2}\mathbf{u}_{n},\mathbf{d}_{n}).$
+The tropicalized Adamax algorithm is then given by Algorithm 2.
 4.2 Computation
-We test the gradient methods discussed on 12 datasets of different sizes, shapes and dimensions; we consider datasets generated by a branching process, a coalescent process, or a Gaussian distribution on $\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$, for $N=6,28$ and samples of size 10 or 100.
-See Appendix A for further details on the simulated datasets.
-The learning rates used have been tuned for each gradient method and dataset size and dimension (Appendix C).
-In all experiments, random initialisations are taken from a Gaussian distribution on $\mathbb{R}^{N}$.
-We provide an overview of the general behaviour of each method for the various optimisation problems introduced in section 2.3.
+We now look to solve our centrality statistics, Wasserstein projection problems, and tropical linear regression using the 7 gradient methods discussed: classical descent (CD), tropical descent (TD), stochastic gradient descent (SGD), tropical stochastic gradient descent(TSGD), Adam, Adamax, and TrAdamax.
+We run computations for 12 datasets of different sizes, shapes and dimensions; motivated by phylogenetic data science applications, we consider datasets generated by branching processes, a coalescent processes, or Gaussian distributions on $\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$, for $N=6,28$ and samples of size 10 or 100.
+We take same random initializations for each gradient method, taken from a Gaussian distribution on $\mathbb{R}^{N}$.
+Appendix B provides further details on the simulated datasets and computational methodology used.
+For each gradient method, we use learning rates which have been optimized for each dataset size and dimension, across different distribution types (Appendix C).
 Rather than investigating the relative convergence rates of each method, we are interested in the relative stability of local minima.
-We therefore plot the CDF of the relative log error after 1000 steps over 50 random initialisations, which can highlight certain error values at which several initialisations have accumulated.
-When the CDF of one method dominates another at log error $=\epsilon$, we expect it has a higher probability of achieving a log error of at most $\epsilon$.
+We therefore plot the CDF of the relative log error after 1000 steps over 50 random initializations, which can highlight certain error values at which several initializations have accumulated.
+When the CDF of one method dominates another, it has a higher probability of achieving the same error and is therefore a more suitable method.
+As we will observe, tropical gradient methods are stable at fewer points than classical gradient methods; as a result, tropical methods typically have further to travel before reaching a stable point.
+Over few steps, this would be compensated for by selecting larger learning rates for tropical gradient methods relative to classical gradient methods, leading to less precise results in the case of tropical gradient methods.
+Our computations are run over 1000 steps, as this gives enough time for tropical gradient methods to stabilise, given our initialisations, without sacrificing learning rate; we see in Appendix C that neither classical nor tropical learning rates are consistently greater than the other.
 4.2.1 Centrality Statistics
-We first consider the behavior of tropical descent methods for the centrality statistics, Fermat–Weber points and Fréchet means.
-These optimisation problems are classically convex problems, as well as being both $\max$-tropical and $\min$-tropical location problems, and therefore lend themselves well to gradient methods.
-The experimental behaviour of these problems is similar, so in this section we include the results for the Fermat–Weber problem while the results for Fréchet means can be found in Appendix E.
-Figure 5 shows the CDF of the final log relative error of the non-stochastic methods for the 12 datasets.
-We note that the stochastic methods achieves the worst error for all 12 datasets, and TropSGD is often performing worse than SGD.
-The disparity between SGD and TropSGD is greatest when the dimensionality of the data is greater than size of the dataset, while the disparity is smallest when the dimensionality is much smaller than the dataset size.
-We see similar behaviour in the log error of Tropical Descent and TropAdamax; they perform on par with their classical counterparts for each dataset other than the 28 by 10 datasets, but when the data dimensionality exceeds the number of data points, the tropical gradient methods are not achieving the same accuracy.
-We also note that convergence of all methods is less consistent for coalescent data than branching or Gaussian data; while it is unclear why this is, we note that coalescent data is supported on the space of ultrametric trees, which is a tropical hyperplane (Ardila & Klivans, (2006); Page et al., (2020)) and may influence the uniqueness of minima.
-Figure 5: The CDF of the log relative error after 1000 steps for each gradient method across 50 initialisations when minimising the Fermat–Weber objective.
-Each subfigure corresponds to a dataset of different dimensionality (columns) and distribution (rows).
-Data CD TD SGD TropSGD Adam Adamax TropAdamax
-6$\times$10 Branching Data -4.60 -4.60 -4.48 -2.95 -4.60 -4.60 -4.60
-6$\times$10 Coalescent Data -4.58 -4.55 -4.38 -1.34 -4.58 -4.58 -4.55
-6$\times$10 Gaussian Data -4.60 -4.60 -4.44 -2.63 -4.58 -4.58 -4.60
-6$\times$100 Branching Data -4.60 -4.60 -4.45 -3.95 -4.59 -4.59 -4.59
-6$\times$100 Coalescent Data -4.55 -4.52 -4.37 -2.76 -4.54 -4.57 -4.54
-6$\times$100 Gaussian Data -4.60 -4.60 -4.45 -4.00 -4.59 -4.60 -4.59
-28$\times$10 Branching Data -4.60 -4.60 -4.15 -1.90 -4.60 -4.60 -4.56
-28$\times$10 Coalescent Data -4.52 -3.76 -3.99 -2.24 -4.52 -4.56 -3.84
-28$\times$10 Gaussian Data -4.59 -4.59 -4.09 -2.47 -4.60 -4.60 -4.47
-28$\times$100 Branching Data -4.59 -4.59 -4.05 -3.13 -4.59 -4.59 -4.59
-28$\times$100 Coalescent Data -4.49 -4.47 -3.86 -2.54 -4.38 -4.56 -4.48
-28$\times$100 Gaussian Data -4.59 -4.59 -4.09 -3.02 -4.59 -4.59 -4.56
-Table 1: The mean log relative error of each gradient method after 1000 steps for each dataset when minimising the Fermat–Weber objective function.
+We first consider the behavior of tropical descent methods for centrality statistics, Fermat–Weber points and Fréchet means.
+These optimization problems are classically convex problems, as well as being both $\max$-tropical and $\min$-tropical location problems, and therefore lend themselves well to gradient methods.
+The experimental behavior of these problems is similar, so in this section we include the results for the Fermat–Weber problem while the results for Fréchet means can be found in Appendix E.
+Figure 5 shows the CDF of the log relative error over 1000 steps of each gradient method for the 12 datasets.
+We note that the deterministic methods achieve similar results for most datasets, while the stochastic methods achieve the worst error for all 12 datasets and TSGD is performing worse than SGD.
+The disparity between SGD and TSGD is greatest when the dimensionality of the data is greater than the size of the dataset, and for coalescent data.
+We see similar behavior in the log error of Tropical Descent and TrAdamax; they perform on par with their classical counterparts for each dataset other than the 28 by 10 coalescent data, for which the tropical gradient methods are not achieving the same accuracy.
+While it is unclear why this is, we note that coalescent data is supported on the space of ultrametric trees, which is a tropical hyperplane (Ardila & Klivans,, 2006; Page et al.,, 2020) and may influence the uniqueness of minima.
+Figure 5: The CDF of the log relative error after 1000 steps for each gradient method across 50 initializations when minimizing the Fermat–Weber objective.
+Each sub-figure corresponds to a dataset of different dimensionality (columns) and distribution (rows).
+Table 1: The mean log relative error of each gradient method after 1000 steps for each dataset when minimizing the Fermat–Weber objective function.
 The minimal mean errors for each dataset are in bold.
+Data CD TD SGD TSGD Adam Adamax TrAdamax
+6$\times$10 Branching Data -4.60 -4.60 -4.50 -3.32 -4.60 -4.60 -4.60
+6$\times$10 Coalescent Data -4.58 -4.55 -4.40 -1.28 -4.57 -4.57 -4.57
+6$\times$10 Gaussian Data -4.60 -4.60 -4.44 -2.75 -4.60 -4.60 -4.56
+6$\times$100 Branching Data -4.60 -4.60 -4.39 -4.02 -4.54 -4.58 -4.59
+6$\times$100 Coalescent Data -4.55 -4.54 -4.25 -2.88 -4.56 -4.58 -4.55
+6$\times$100 Gaussian Data -4.60 -4.60 -4.39 -4.03 -4.52 -4.57 -4.59
+28$\times$10 Branching Data -4.60 -4.60 -4.15 -1.89 -4.60 -4.60 -4.56
+28$\times$10 Coalescent Data -4.53 -3.76 -3.96 -2.11 -4.52 -4.57 -3.87
+28$\times$10 Gaussian Data -4.58 -4.58 -4.07 -2.40 -4.60 -4.60 -4.47
+28$\times$100 Branching Data -4.59 -4.57 -4.04 -2.89 -4.59 -4.59 -4.59
+28$\times$100 Coalescent Data -4.51 -4.54 -3.91 -2.33 -4.39 -4.58 -4.39
+28$\times$100 Gaussian Data -4.59 -4.53 -4.10 -2.78 -4.59 -4.59 -4.59
 4.2.2 Wasserstein Projections
-The Wasserstein projection problems highlight the differences in convergent behaviour with and without the tropical gradient, as they are $\max$-tropical location problems, but not classically convex.
+The Wasserstein projection problems highlight the differences in convergent behavior between classical and tropical gradient methods, as they are $\min$-tropical location problems but not classically convex.
 The CDFs of the final log relative errors can be found in Figures 6 and 7.
-Figure 6: The CDF of the log relative error after 1000 steps for each gradient method across 50 initialisations when minimising the 2-Wasserstein projection objective.
-Each subfigure corresponds to a dataset of different dimensionality (columns) and distribution (rows).
-Figure 7: The CDF of the log relative error after 1000 steps for each gradient method across 50 initialisations when minimising the $\infty$-Wasserstein projection objective.
-Each subfigure corresponds to a dataset of different dimensionality (columns) and distribution (rows).
+Figure 6: The CDF of the log relative error after 1000 steps for each gradient method across 50 initializations when minimizing the 2-Wasserstein projection objective.
+Each sub-figure corresponds to a dataset of different dimensionality (columns) and distribution (rows).
+Figure 7: The CDF of the log relative error after 1000 steps for each gradient method across 50 initializations when minimizing the $\infty$-Wasserstein projection objective.
+Each sub-figure corresponds to a dataset of different dimensionality (columns) and distribution (rows).
 In Figures 6 and 7 we see large steps at various intervals in the CDF plots, which is indicative of stable local minima.
 This is particularly prevalent in low dimensional space.
-Classical descent, Adam, and Adamax are particularly susceptible to this, while tropical descent and TropAdamax are often more likely to pass these local minima.
+Classical descent, Adam, and Adamax are particularly susceptible to this, while tropical descent and TrAdamax are more likely to pass these local minima.
 This difference is most pronounced when solving the $\infty$-Wasserstein problem in low dimensions—we are almost certain to reach the minimum with small error, while classical methods have a 50-80% chance of achieving the same error.
-Data CD TD SGD TropSGD Adam Adamax TropAdamax
-6$\times$10 Branching Data -3.52 -4.13 -1.98 -2.17 -3.49 -3.51 -3.99
-6$\times$10 Coalescent Data -3.53 -4.31 -2.38 -3.17 -3.77 -3.85 -4.16
-6$\times$10 Gaussian Data -3.66 -3.72 -3.21 -2.76 -3.84 -3.85 -3.90
-6$\times$100 Branching Data -3.57 -4.38 -3.28 -1.64 -3.94 -3.94 -4.39
-6$\times$100 Coalescent Data -4.35 -4.54 -3.31 -2.54 -4.43 -4.44 -4.26
-6$\times$100 Gaussian Data -3.78 -4.15 -3.39 -2.69 -3.77 -3.79 -4.42
-28$\times$10 Branching Data -3.59 -4.29 -3.18 -2.03 -3.63 -3.64 -4.34
-28$\times$10 Coalescent Data -4.07 -3.80 -3.45 -2.50 -4.12 -4.09 -3.82
-28$\times$10 Gaussian Data -3.53 -3.89 -3.22 -2.18 -3.61 -3.58 -3.85
-28$\times$100 Branching Data -3.65 -4.54 -3.42 -2.64 -3.86 -3.87 -4.53
-28$\times$100 Coalescent Data -4.12 -4.57 -3.89 -3.44 -4.20 -4.21 -4.55
-28$\times$100 Gaussian Data -3.65 -4.50 -3.39 -2.67 -3.85 -3.86 -4.48
-Table 2: The mean log relative error of each gradient method after 1000 steps for each dataset when minimising the $2$-Wasserstein projection objective function.
+4.2.3 Linear Regression
+As the coalescent data lies on a tropical hyperplane (Ardila & Klivans,, 2006; Page et al.,, 2020), the minimum of the linear regression function for such data is exactly 0.
+Figure 8 therefore shows the CDF of the log error rather than the CDF of the log relative error.
+Table 2 shows the mean log errors of each gradient method and dataset, from which we see that the tropical methods outperform their classical counterparts for all but one dataset, often by a factor of two.
+In Figure 8 we see that the CDF of log errors for tropical descent and TrAdamax consistently dominates those of classical descent, Adam and Adamax, to the extent that the support of their CDFs does not intersect that of their classical counterparts for some datasets.
+We also note that unlike the previous problems, TSGD is outperforming SGD when solving the linear regression problem, often to the extent that TSGD outperforms all classical gradient methods.
+Figure 8: The CDF of the log relative error after 1000 steps for each gradient method across 50 initializations when minimizing the linear regression objective.
+Each sub-figure corresponds to a dataset of different dimensionality (columns) and distribution (rows).
+Table 2: The mean log relative error of each gradient method after 1000 steps for each dataset when minimizing the linear regression objective function.
 The minimal mean errors for each dataset are in bold.
-Data CD TD SGD TropSGD Adam Adamax TropAdamax
-6$\times$10 Branching Data -2.56 -4.29 -0.58 -0.94 -2.58 -2.59 -4.28
-6$\times$10 Coalescent Data -4.07 -4.48 -1.13 -1.38 -4.13 -4.09 -4.48
-6$\times$10 Gaussian Data -4.00 -4.44 -2.08 -3.10 -4.04 -4.02 -4.38
-6$\times$100 Branching Data -3.63 -4.46 -2.14 -1.17 -3.60 -3.65 -4.46
-6$\times$100 Coalescent Data -4.11 -4.51 -2.19 -2.30 -4.18 -4.14 -4.51
-6$\times$100 Gaussian Data -3.02 -4.36 -1.48 -1.23 -3.06 -3.09 -4.31
-28$\times$10 Branching Data -3.72 -3.66 -2.19 -1.14 -3.81 -3.81 -3.55
-28$\times$10 Coalescent Data -4.31 -4.13 -2.83 -1.99 -4.39 -4.39 -4.01
-28$\times$10 Gaussian Data -4.16 -3.47 -2.08 -0.96 -4.43 -4.36 -3.12
-28$\times$100 Branching Data -3.55 -3.86 -1.38 -1.16 -3.59 -3.49 -3.71
-28$\times$100 Coalescent Data -4.45 -4.12 -2.11 -2.05 -4.54 -4.54 -3.95
-28$\times$100 Gaussian Data -3.23 -3.61 -1.74 -1.61 -3.27 -3.22 -2.94
-Table 3: The mean log relative error of each gradient method after 1000 steps for each dataset when minimising the $\infty$-Wasserstein projection objective function.
-The minimal mean errors for each dataset are in bold.
-From tables 2 and 3, we see that when solving the 2-Wasserstein projection problem, tropical methods are optimal in all cases other than $28\times 10$ coalescent data.
-For the $\infty$-Wasserstein projection problem in small dimensions, tropical descent outperforms other methods by some margin, while the comparative performances are less consistent for the $\infty$-Wasserstein projection problem in high dimensions.
-4.2.3 Linear Regression
-The linear regression problem is a $\min$ tropical location problem, so we run the same code with input $-\mathbf{t}$.
-As the coalescent data lies on a tropical hyperplane (Ardila & Klivans, (2006); Page et al., (2020)), the minimum of the linear regression function for such data is exactly 0.
-Figure 8 therefore shows the CDF of the log error rather than the CDF of the log relative error.
-As we can see in table 4, the tropical methods outperform their classical counterparts for every dataset, often by a factor of two.
-The tropical descent method in particular is the most effective for all but one dataset.
-In Figure 8 we see that the CDF of log errors for tropical descent and TropAdamax consistently dominates those of classical descent, Adam and Adamax, to the extent that the support of their CDFs does not intersect that of their classical counterparts for some datasets.
-We also note that unlike the previous problems, TropSGD is outperforming SGD when solving the linear regression problem, often to the extent that TropSGD outperforms all classical gradient methods.
-Figure 8: The CDF of the log relative error after 1000 steps for each gradient method across 50 initialisations when minimising the linear regression objective.
-Each subfigure corresponds to a dataset of different dimensionality (columns) and distribution (rows).
-Data CD TD SGD TropSGD Adam Adamax TropAdamax
-6$\times$10 Branching Data -0.96 -5.96 -0.92 -2.10 -1.09 -1.07 -5.57
-6$\times$10 Coalescent Data -4.27 -6.25 -3.80 -5.78 -4.44 -4.54 -6.22
-6$\times$10 Gaussian Data -1.13 -5.42 -0.86 -1.33 -1.15 -1.14 -4.79
-6$\times$100 Branching Data -1.08 -5.34 -0.86 -1.54 -1.10 -1.11 -3.73
-6$\times$100 Coalescent Data -6.42 -6.35 -3.29 -5.69 -6.81 -6.88 -6.22
-6$\times$100 Gaussian Data -0.93 -5.43 -1.09 -1.93 -0.92 -0.92 -3.97
-28$\times$10 Branching Data -1.60 -5.83 -1.13 -4.47 -1.41 -1.48 -5.53
-28$\times$10 Coalescent Data -2.56 -6.52 -2.22 -5.50 -2.68 -2.61 -6.44
-28$\times$10 Gaussian Data -1.35 -5.68 -0.91 -3.99 -1.20 -1.23 -5.56
-28$\times$100 Branching Data -0.79 -4.33 -0.54 -1.46 -0.69 -0.70 -3.44
-28$\times$100 Coalescent Data -3.35 -5.34 -2.10 -6.04 -3.79 -3.94 -4.95
-28$\times$100 Gaussian Data -0.93 -3.88 -0.64 -1.22 -0.84 -0.85 -3.44
-Table 4: The mean log relative error of each gradient method after 1000 steps for each dataset when minimising the linear regression objective function.
-The minimal mean errors for each dataset are in bold.
-5 Discussion
+Data CD TD SGD TSGD Adam Adamax TrAdamax
+6$\times$10 Branching Data -0.97 -5.79 -0.56 -2.16 -0.80 -0.84 -4.39
+6$\times$10 Coalescent Data -2.70 -5.87 -2.11 -5.47 -2.97 -2.66 -6.58
+6$\times$10 Gaussian Data -1.19 -5.48 -0.80 -1.35 -1.05 -1.10 -4.10
+6$\times$100 Branching Data -0.79 -5.49 -0.71 -1.43 -0.72 -0.72 -3.53
+6$\times$100 Coalescent Data -5.54 -5.76 -3.11 -5.57 -6.90 -7.01 -6.52
+6$\times$100 Gaussian Data -1.04 -5.50 -0.88 -1.70 -1.02 -1.02 -3.81
+28$\times$10 Branching Data -1.49 -5.51 -0.93 -4.49 -1.14 -1.14 -5.57
+28$\times$10 Coalescent Data -2.43 -6.63 -1.77 -5.55 -2.16 -2.29 -6.18
+28$\times$10 Gaussian Data -1.26 -5.16 -0.79 -4.29 -1.08 -1.08 -5.72
+28$\times$100 Branching Data -0.90 -4.37 -0.61 -1.42 -0.85 -0.77 -3.32
+28$\times$100 Coalescent Data -3.33 -5.84 -2.17 -5.35 -3.02 -3.20 -4.93
+28$\times$100 Gaussian Data -0.91 -4.35 -0.62 -1.07 -0.96 -0.90 -3.51
+5 Conclusion and Discussion
 We have introduced the concept of tropical descent as the steepest descent method with respect to the tropical norm.
-We have demonstrated, both theoretically and experimentally, that it possesses well-behaved properties with respect to a wide class of optimisation problems on the tropical projective torus, including several key statistical problems for the analysis of phylogenetic data.
-However, our experimental results are stronger than our theoretical results suggest; the optimisation problems studied are location problems, rather than $\triangle$-star-quasi-convex problems, and yet tropical descent appears to consistently converge to a global minimum, not just the tropical convex hull of our dataset.
+We have demonstrated, both theoretically and experimentally, that it possesses well-behaved properties with respect to a wide class of optimization problems on the tropical projective torus, including several key statistical problems for the analysis of phylogenetic data.
+As well as proposing our own framework for tropical gradient methods, this work also provides a detailed review of the varying behavior of popular gradient methods when applied to tropical statistical optimization problems.
+Our theoretical results outline the relative strengths of tropical descent over classical descent for tropical location problems, but only provide global solvability for 1-sample problems.
+In practice, our experimental results go further; the optimization problems studied are location problems, rather than $\triangle$-star-quasi-convex problems, and yet tropical descent appears to consistently converge to a global minimum, not just the tropical convex hull of our dataset.
 It would therefore be a natural extension of this work to identify how much we can relax the $\triangle$-star-quasi-convexity condition while guaranteeing convergence.
-In this paper, we consider tree data with 4 to 8 leaves, and samples of size 10 to 100—at this scale, we have observed (Trop/)SGD to be a significant sacrifice in accuracy with little benefit in computational time.
-Furthermore, at this scale, the relative merits of SGD versus its tropical counterpart are inconsistent; TropSGD out-performs SGD when solving the linear regression problem, but performs worse when finding centrality statistics.
-However, for applications to larger datasets, we would expect the stochastic approaches to be more suitable.
-It would then be important to understand the theoretical differences between SGD and TropSGD in terms of their convergence guarantees in the tropical setting and their comparative performance in practice.
-Finally, as shown by Akian et al., (2023), the tropical linear regression problem is polynomial-time equivalent to mean payoff games - two-player perfect information games on a directed graph (Ehrenfeucht & Mycielski, (1979); Gurvich et al., (1988)).
-This family of problems of particular interest in game theory as they are known to be NP $\cap$ co-NP.
-While our methods are designed to approximate the minimum rather than compute it exactly, tropical descent may prove to be an efficient tool when solving large-scale mean payoff games.
-Acknowledgments
+In this paper we have considered tree data with 4 to 8 leaves, and samples of size 10 to 100—at this scale, we have observed (T/)SGD to be a significant sacrifice in accuracy with little benefit in computational time.
+Furthermore, at this scale, the relative merits of SGD versus its tropical counterpart are inconsistent; TSGD out-performs SGD when solving the linear regression problem, but performs worse when finding centrality statistics.
+In a similar vein, initial investigations suggest that the classical stochastic algorithm for the computation of Fréchet means—Sturm’s algorithm (Sturm,, 2003)—performs poorly for the computation of tropical Fréchet means due to the inconsistent curvature of the tropical projective torus (Améndola & Monod,, 2023).
+However, for applications to larger datasets and more complex objective functions, stochastic gradient methods may become necessary.
+It would then be important to understand the theoretical differences between SGD and TSGD in terms of their convergence guarantees in the tropical setting and their comparative performance at different scales of data dimension.
+The recent identification of tropical rational functions and ReLU neural networks (Zhang et al.,, 2018) has inspired new avenues of research in tropical applications, into the tropical geometry of neural networks (Brandenburg et al.,, 2024; Yoshida et al.,, 2024).
+It is therefore natural to consider the possible extensions of this work to neural network training; can we identify local tropical quasi-convexity in the loss functions of ReLU neural networks, and how can we extend tropical steepest descent to the space of tropical rational functions?
+Finally, as shown by (Akian et al.,, 2023), the tropical linear regression problem is polynomial-time equivalent to mean payoff games—two-player perfect information games on a directed graph (Ehrenfeucht & Mycielski,, 1979; Gurvich et al.,, 1988).
+This family of problems is of particular interest in game theory as they are known to be NP $\cap$ co-NP.
+In our numerical experiments, tropical gradient methods demonstrate a particularly strong performance in solving tropical linear regression, and while our methods are designed to approximate the minimum rather than compute it exactly, tropical descent may prove to be an efficient tool when solving large-scale mean payoff games.
+Acknowledgements
 R.T.
-receives partial funding from a Technical University of Munich–-Imperial College London Joint Academy of Doctoral Studies (JADS) award (2021 cohort, PIs Drton/Monod).
+receives partial funding from a Technical University of Munich–Imperial College London Joint Academy of Doctoral Studies (JADS) award (2021 cohort, PIs Drton/Monod).
 A.M.
-is supported by the Engineering and Physical Sciences Research Council under grant reference [EP/Y028872/1] and a London Mathematical Society Emmy Noether Fellowship.
-Appendix A Data Generation
-The same 12 datasets are used for each experiment in Section 4.
-These datasets live in $\mathbb{R}^{6}/\mathbb{R}\mathbf{1}$ or $\mathbb{R}^{28}/\mathbb{R}\mathbf{1}$; these dimensions correspond to the spaces of trees with 4 and 8 leaves respectively.
-Branching process datasets are generated using the R package ape (Paradis & Schliep, (2019)) with exponentially distributed branch lengths—this produces data supported on the space of metric trees, i.e., the tropical Grassmannian.
-Coalescent data has also been generated using ape and is supported on the space of ultrametric trees, which is a tropical linear space.
-Gaussian data is generated using NumPy (Harris et al., (2020)), and is supported on the general tropical projective torus.
-In the computation of the Wasserstein projection objectives, we require a secondary dataset of different dimensionality; for each of the 12 datasets, we generate datasets of the same size and distribution, but in $\mathbb{R}^{3}/\mathbb{R}\mathbf{1}$ or $\mathbb{R}^{21}/\mathbb{R}\mathbf{1}$ which contain the spaces of trees with 3 and 7 leaves respectively.
-Appendix B PCA and Logistic Regression
-Figure 9: The CDF of the log relative error after 1000 steps for each gradient method across 50 initialisations when minimising the PCA objective.
-Each subfigure corresponds to a dataset of different dimensionality (columns) and distribution (rows).
-Figure 10: The CDF of the log relative error after 1000 steps for each gradient method across 50 initialisations when minimising the logistic regression objective.
-Each subfigure corresponds to a dataset of different dimensionality (columns) and distribution (rows).
-As well as the location problems outlined in this paper, we performed the same numerical experiments with the tropical polytope PCA loss function (Yoshida et al., (2019)) and tropical logistic regression loss function (Aliatimis et al., (2023)).
+is supported by the Engineering and Physical Sciences Research Council under grant reference [EP/Y028872/1] and a London Mathematical Society Emmy Noether Fellowship [EN-2223-01].
+Declarations
+The authors have no relevant financial or non-financial interests to disclose.
+The authors have no competing interests to declare that are relevant to the content of this article.
+All authors certify that they have no affiliations with or involvement in any organization or entity with any financial interest or non-financial interest in the subject matter or materials discussed in this manuscript.
+The authors have no financial or proprietary interests in any material discussed in this article.
+Appendix A PCA and Logistic Regression
+Figure 9: The CDF of the log relative error after 1000 steps for each gradient method across 50 initializations when minimizing the PCA objective.
+Each sub-figure corresponds to a dataset of different dimensionality (columns) and distribution (rows).
+Figure 10: The CDF of the log relative error after 1000 steps for each gradient method across 50 initializations when minimizing the logistic regression objective.
+Each sub-figure corresponds to a dataset of different dimensionality (columns) and distribution (rows).
+As well as the location problems outlined in this paper, we performed the same numerical experiments with the tropical polytope PCA loss function (Yoshida et al.,, 2019) and tropical logistic regression loss function (Aliatimis et al.,, 2024).
 These loss functions take multiple points in $\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ as argument, and we execute our gradient methods on each argument simultaneously.
@@
 These problems do not lend themselves well to gradient methods, as the relative log error is rather large for all of our proposed methods.
-However, for both problems the tropical descent and TropAdamax methods are outperforming the classical methods quite consistently.
+However, for both problems the tropical descent and TrAdamax methods are outperforming the classical methods quite consistently.
+Appendix B Methodology
+The same 12 datasets are used for each experiment in Section 4.
+These datasets live in $\mathbb{R}^{6}/\mathbb{R}\mathbf{1}$ or $\mathbb{R}^{28}/\mathbb{R}\mathbf{1}$; these dimensions correspond to the spaces of trees with 4 and 8 leaves respectively.
+Branching process datasets are generated using the R package ape (Paradis & Schliep,, 2019) with exponentially distributed branch lengths—this produces data supported on the space of metric trees, i.e., the tropical Grassmannian.
+Coalescent data has also been generated using ape and is supported on the space of ultrametric trees, which is a tropical linear space.
+Gaussian data is generated using NumPy (Harris et al.,, 2020), and is supported on the general tropical projective torus.
+In the computation of the Wasserstein projection objectives, we require a secondary dataset of different dimensionality; for each of the 12 datasets, we generate datasets of the same size and distribution, but in $\mathbb{R}^{3}/\mathbb{R}\mathbf{1}$ or $\mathbb{R}^{21}/\mathbb{R}\mathbf{1}$ which contain the spaces of trees with 3 and 7 leaves respectively.
+We use PyTorch (Paszke et al.,, 2019) for the computation and differentiation of each objective function.
+Max-tropical descent directions are used for tropical gradient methods; we therefore optimize the min-tropical location problems—Wasserstein projections and tropical linear regression—with respect to $-\mathbf{t}$.
 Appendix C Learning Rate Tuning
-The Adam algorithms—Adam, Adamax and TropAdamax—are run using the default $\beta=(0.9,0.999)$ parameters.
-We then tune the learning rates used for each method; Gradient Descent, Tropical Descent, SGD, Tropical SGD, Adam, Adamax, and TropAdamax.
-We perform an initial test of learning rates $\gamma=0.0001,0.0003,0.001,0.003,0.01,0.03,0.1,0.3,1,3,10,30$ over 1000 steps for each algorithm, with 5 random initialisations for each dataset.
+The Adam algorithms — Adam, Adamax and TrAdamax — are run using the default $\beta=(0.9,0.999)$ parameters.
+We then tune the learning rates used for each method; Gradient Descent, Tropical Descent, SGD, Tropical SGD, Adam, Adamax, and TrAdamax.
+We perform an initial test of learning rates $\gamma$ on a logarithmic scale from 0.0001 to 30 over 1000 steps for each algorithm, with 5 random initializations for each dataset.
 From this, we narrowed our search to take 18 initializations at 0.5 intervals on the log scale in near optimal regions for each problem, gradient method, and data dimensionality.
-We then selected the learning rate with minimal mean log relative error in each case; tables 5 - 11 show the selected learning rates.
-Data Dim Data Count CD TD SGD TropSGD Adam Adamax TropAdamax
-6 10 0.135 0.135 0.135 0.0821 0.00409 0.00409 0.0111
-100 0.135 0.135 0.0821 0.135 0.00409 0.00409 0.00674
-28 10 0.368 0.368 0.368 0.607 0.00674 0.00674 0.0183
-100 0.368 0.223 0.368 0.607 0.00674 0.00674 0.0111
+We then selected the learning rate with minimal mean log relative error in each case; Tables 3, 4, 5, 6, 7, 8 and 9 show the selected learning rates.
 Table 3: Learning rates for each gradient method when solving the Fermat–Weber problem.
-Data Dim Data Count CD TD SGD TropSGD Adam Adamax TropAdamax
+Data Dim Data Count CD TD SGD TSGD Adam Adamax TrAdamax
 6 10 0.135 0.135 0.135 0.135 0.00674 0.00674 0.00674
-100 0.135 0.0821 0.0821 0.223 0.00674 0.00674 0.0111
-28 10 0.368 0.223 0.223 0.607 0.00674 0.00674 0.0183
-100 0.368 0.135 0.223 0.607 0.00674 0.00674 0.0111
+100 0.135 0.135 0.135 0.135 0.00248 0.00248 0.00674
+28 10 0.368 0.368 0.368 1 0.00674 0.00674 0.0183
+100 0.368 0.135 0.368 1 0.00674 0.00674 0.0183
 Table 4: Learning rates for each gradient method when solving the Fréchet mean problem.
-Data Dim Data Count CD TD SGD TropSGD Adam Adamax TropAdamax
-6 10 1.65 7.39 0.0821 0.223 0.135 0.223 0.0821
-100 1.65 4.48 0.368 0.135 0.0821 0.135 1.00
-28 10 1.65 0.368 0.607 0.607 0.00674 0.0111 0.0183
-100 2.72 2.72 0.368 0.607 0.0302 0.0821 0.0498
+Data Dim Data Count CD TD SGD TSGD Adam Adamax TrAdamax
+6 10 0.135 0.135 0.135 0.135 0.00674 0.00674 0.00674
+100 0.135 0.135 0.135 0.135 0.00674 0.00248 0.00674
+28 10 0.368 0.368 0.368 1 0.00674 0.00674 0.0183
+100 0.368 0.135 0.368 1 0.00674 0.00674 0.0183
 Table 5: Learning rates for each gradient method when solving the 2-Wasserstein projection problem.
-Data Dim Data Count CD TD SGD TropSGD Adam Adamax TropAdamax
-6 10 0.0821 0.135 0.0302 0.223 0.00674 0.00674 0.0183
-100 0.368 0.135 0.607 0.0498 0.00248 0.00409 0.0183
-28 10 0.223 0.607 0.223 0.607 0.00409 0.0111 0.0821
-100 0.223 0.607 0.223 0.607 0.00409 0.0111 0.0821
+Data Dim Data Count CD TD SGD TSGD Adam Adamax TrAdamax
+6 10 2.72 1 1 0.135 0.135 0.368 0.368
+100 2.72 2.72 0.368 0.135 0.135 0.135 0.135
+28 10 0.368 0.368 0.368 1 0.00674 0.0183 0.0183
+100 2.72 2.72 0.368 1 0.0498 0.135 0.0498
 Table 6: Learning rates for each gradient method when solving the $\infty$-Wasserstein projection problem.
-Data Dim Data Count CD TD SGD TropSGD Adam Adamax TropAdamax
-6 10 2.72 4.48 0.607 1.65 0.135 0.135 0.368
-100 2.72 4.48 1.00 0.607 0.0821 0.135 0.607
-28 10 7.39 4.48 0.368 1.65 0.0183 0.0183 0.223
-100 4.48 7.39 0.368 0.607 0.0302 0.0821 0.0821
+Data Dim Data Count CD TD SGD TSGD Adam Adamax TrAdamax
+6 10 0.0498 0.135 0.0498 0.368 0.00674 0.00674 0.0183
+100 0.0498 0.135 0.135 2.72 0.00674 0.00674 0.0183
+28 10 0.368 1 0.368 1 0.00674 0.0183 0.0498
+100 0.368 1 0.368 1 0.00674 0.0183 0.0498
 Table 7: Learning rates for each gradient method when solving the polytope PCA problem.
-Data Dim Data Count CD TD SGD TropSGD Adam Adamax TropAdamax
-6 10 7.39 12.2 1.00 2.72 1.00 1.00 0.607
-100 7.39 12.2 1.00 0.368 0.607 1.00 1.00
-28 10 20.1 245 7.39 90.0 4.48 4.48 2.72
-100 20.1 90.0 4.48 90.0 1.65 7.39 2.72
+Data Dim Data Count CD TD SGD TSGD Adam Adamax TrAdamax
+6 10 2.72 2.72 1 1 0.135 0.135 0.368
+100 2.72 7.39 1 1 0.135 0.135 1
+28 10 7.39 7.39 0.368 1 0.0183 0.0183 0.135
+100 2.72 7.39 0.368 1 0.0498 0.0498 0.0498
 Table 8: Learning rates for each gradient method when solving the logistic regression problem.
-Data Dim Data Count CD TD SGD TropSGD Adam Adamax TropAdamax
-6 10 0.0302 0.0821 0.223 0.0821 0.00248 0.00248 0.0111
-100 0.0498 0.0821 0.368 0.0821 0.00248 0.00248 0.0111
-28 10 0.368 0.0821 0.0821 0.135 0.0111 0.0302 0.0111
-100 0.223 0.607 1.00 12.2 0.0015 0.00409 0.0498
+Data Dim Data Count CD TD SGD TSGD Adam Adamax TrAdamax
+6 10 7.39 2.72 1 2.72 0.368 1 0.0498
+100 1 0.368 1 2.72 0.00248 0.00248 0.00248
+28 10 1 0.368 1 2.72 0.00248 0.00248 0.00674
+100 1 0.368 1 7.39 0.00248 0.00248 0.00674
 Table 9: Learning rates for each gradient method when solving the linear regression problem.
+Data Dim Data Count CD TD SGD TSGD Adam Adamax TrAdamax
+6 10 0.368 0.135 0.368 0.135 0.00248 0.0183 0.00674
+100 0.135 0.135 0.368 0.0498 0.00248 0.00248 0.00674
+28 10 0.368 0.0498 0.0498 0.135 0.00248 0.00248 0.0183
+100 1 0.368 1 7.39 0.0498 0.0498 0.0498
 Appendix D Runtimes
-Experiments were implemented on a AMD EPYC 7742 node using a single core, 8 GB. Table 12 shows the average time taken per initalisation for each gradient method and optimisation problem.
-Stats Prob CD TD SGD TropSGD Adam Adamax TropAdamax
-Fermat–Weber 0.18s 0.19s 0.26s 0.26s 0.22s 0.23s 0.26s
-Fréchet Mean 0.23s 0.23s 0.30s 0.30s 0.28s 0.28s 0.31s
-2-Wasserstein Projection 1.18s 1.18s 1.75s 1.77s 1.23s 1.24s 1.27s
-$\infty$-Wasserstein Projection 1.17s 1.17s 1.71s 1.71s 1.21s 1.21s 1.24s
-Polytope PCA 0.60s 0.61s 0.78s 0.78s 0.64s 0.64s 0.68s
-Logistic Regression 0.45s 0.45s 0.62s 0.62s 0.68s 0.50s 0.53s
-Linear Regression 0.22s 0.22s 0.27s 0.28s 0.26s 0.27s 0.30s
-Table 12: Average runtimes for 1000 steps of each gradient method when minimising each objective function.
+Experiments were implemented on a AMD EPYC 7742 node using a single core, 8 GB. Table 10 shows the average time taken per initialization for each gradient method and optimization problem.
+Table 10: Average runtimes for 1000 steps of each gradient method when minimizing each objective function.
+Stats Prob CD TD SGD TSGD Adam Adamax TrAdamax
+Fermat–Weber 0.480 0.520 0.435 0.499 0.391 0.460 0.464
+Fréchet Mean 0.347 0.363 0.301 0.342 0.294 0.422 0.370
+2-Wasserstein Projection 2.739 2.134 2.392 2.209 2.341 2.246 2.272
+$\infty$-Wasserstein Projection 2.947 3.049 2.772 2.565 2.501 3.162 2.807
+Polytope PCA 1.117 1.081 1.014 1.052 1.037 1.174 1.029
+Logistic Regression 0.987 1.015 0.962 0.934 0.943 1.084 0.943
+Linear Regression 0.494 0.499 0.386 0.420 0.386 0.537 0.433
 Appendix E Further Figures
+E.1 Fréchet Means
 Here we include the results for experiments run for the computation of Fréchet means.
-Data CD TD SGD TropSGD Adam Adamax TropAdamax
-6$\times$10 Branching Data -4.59 -4.56 -4.01 -2.93 -4.58 -4.59 -4.54
-6$\times$10 Coalescent Data -4.58 -4.49 -4.06 -1.86 -4.57 -4.58 -4.50
-6$\times$10 Gaussian Data -4.59 -4.35 -4.10 -2.63 -4.58 -4.59 -4.32
-6$\times$100 Branching Data -4.60 -4.59 -4.36 -3.90 -4.59 -4.59 -4.59
-6$\times$100 Coalescent Data -4.57 -4.57 -4.44 -3.14 -4.54 -4.56 -4.55
-6$\times$100 Gaussian Data -4.59 -4.59 -4.44 -4.03 -4.59 -4.60 -4.59
-28$\times$10 Branching Data -4.57 -3.69 -3.66 -1.99 -4.54 -4.57 -3.79
-28$\times$10 Coalescent Data -4.56 -3.72 -4.01 -2.57 -4.52 -4.57 -3.71
-28$\times$10 Gaussian Data -4.56 -4.05 -3.71 -2.49 -4.54 -4.57 -4.02
-28$\times$100 Branching Data -4.59 -4.58 -3.99 -3.23 -4.58 -4.59 -4.58
-28$\times$100 Coalescent Data -4.54 -4.56 -4.14 -2.87 -4.46 -4.57 -4.52
-28$\times$100 Gaussian Data -4.59 -4.55 -3.90 -3.01 -4.59 -4.59 -4.58
-Table 13: The mean log relative error of each gradient method after 1000 steps for each dataset when minimising the Fréchet mean objective function.
+Table 11: The mean log relative error of each gradient method after 1000 steps for each dataset when minimizing the Fréchet mean objective function.
 The minimal mean errors for each dataset are in bold.
-Figure 11: The CDF of the log relative error after 1000 steps for each gradient method across 50 initialisations when minimising the Fréchet mean objective.
-Each subfigure corresponds to a dataset of different dimensionality (columns) and distribution (rows).
+Data CD TD SGD TSGD Adam Adamax TrAdamax
+6$\times$10 Branching Data -4.59 -4.54 -3.98 -2.99 -4.58 -4.59 -4.53
+6$\times$10 Coalescent Data -4.58 -4.47 -4.07 -1.82 -4.57 -4.58 -4.45
+6$\times$10 Gaussian Data -4.59 -4.33 -4.09 -2.63 -4.58 -4.59 -4.35
+6$\times$100 Branching Data -4.59 -4.59 -4.35 -4.03 -4.59 -4.58 -4.59
+6$\times$100 Coalescent Data -4.57 -4.56 -4.36 -3.29 -4.55 -4.56 -4.57
+6$\times$100 Gaussian Data -4.59 -4.60 -4.39 -4.12 -4.59 -4.58 -4.59
+28$\times$10 Branching Data -4.58 -3.73 -3.78 -1.98 -4.55 -4.58 -3.83
+28$\times$10 Coalescent Data -4.56 -3.67 -3.98 -2.46 -4.53 -4.57 -3.70
+28$\times$10 Gaussian Data -4.55 -4.07 -3.91 -2.41 -4.55 -4.57 -4.04
+28$\times$100 Branching Data -4.59 -4.58 -4.03 -2.98 -4.58 -4.59 -4.57
+28$\times$100 Coalescent Data -4.54 -4.56 -4.08 -2.66 -4.47 -4.57 -4.46
+28$\times$100 Gaussian Data -4.59 -4.56 -4.08 -2.78 -4.59 -4.59 -4.58
+Figure 11: The CDF of the log relative error after 1000 steps for each gradient method across 50 initializations when minimizing the Fréchet mean objective.
+Each sub-figure corresponds to a dataset of different dimensionality (columns) and distribution (rows).
+E.2 Wasserstein Projections
+From Tables 12 and 13, we see that when solving the 2-Wasserstein projection problem, tropical methods are optimal in all cases other than $28\times 10$ coalescent data.
+For the $\infty$-Wasserstein projection problem in small dimensions, tropical descent outperforms other methods by some margin, while the comparative performances are less consistent for the $\infty$-Wasserstein projection problem in high dimensions.
+Table 12: The mean log relative error of each gradient method after 1000 steps for each dataset when minimizing the $2$-Wasserstein projection objective function.
+The minimal mean errors for each dataset are in bold.
+Data CD TD SGD TSGD Adam Adamax TrAdamax
+6$\times$10 Branching Data -3.02 -4.09 -2.16 -2.11 -2.95 -3.00 -3.89
+6$\times$10 Coalescent Data -3.71 -3.92 -2.44 -3.13 -3.78 -3.68 -4.12
+6$\times$10 Gaussian Data -3.59 -3.83 -2.87 -2.75 -3.61 -3.57 -3.89
+6$\times$100 Branching Data -3.44 -4.24 -2.91 -1.60 -3.79 -3.79 -4.02
+6$\times$100 Coalescent Data -4.26 -4.55 -3.27 -2.58 -4.39 -4.45 -4.50
+6$\times$100 Gaussian Data -3.69 -4.17 -3.33 -2.73 -3.78 -3.78 -4.18
+28$\times$10 Branching Data -3.39 -4.31 -3.16 -2.03 -3.53 -3.55 -4.32
+28$\times$10 Coalescent Data -4.11 -3.79 -3.39 -2.42 -4.17 -4.12 -3.80
+28$\times$10 Gaussian Data -3.43 -3.91 -3.21 -2.15 -3.50 -3.44 -3.83
+28$\times$100 Branching Data -3.61 -4.55 -3.37 -2.65 -3.85 -3.85 -4.51
+28$\times$100 Coalescent Data -4.11 -4.57 -3.89 -3.32 -4.21 -4.19 -4.56
+28$\times$100 Gaussian Data -3.66 -4.52 -3.41 -2.61 -3.82 -3.80 -4.49
+Table 13: The mean log relative error of each gradient method after 1000 steps for each dataset when minimizing the $\infty$-Wasserstein projection objective function.
+The minimal mean errors for each dataset are in bold.
+Data CD TD SGD TSGD Adam Adamax TrAdamax
+6$\times$10 Branching Data -1.63 -4.32 -0.57 -0.91 -1.68 -1.69 -4.32
+6$\times$10 Coalescent Data -4.00 -4.47 -0.99 -1.43 -4.03 -4.04 -4.46
+6$\times$10 Gaussian Data -3.77 -4.42 -1.76 -3.33 -3.76 -3.75 -4.40
+6$\times$100 Branching Data -3.35 -4.48 -1.94 -0.95 -3.34 -3.36 -4.47
+6$\times$100 Coalescent Data -3.85 -4.51 -2.06 -2.21 -3.84 -3.85 -4.51
+6$\times$100 Gaussian Data -3.63 -4.40 -1.61 -1.22 -3.70 -3.63 -4.36
+28$\times$10 Branching Data -3.68 -3.53 -2.22 -1.17 -4.11 -3.93 -3.17
+28$\times$10 Coalescent Data -4.22 -3.95 -2.76 -1.92 -4.42 -4.36 -3.94
+28$\times$10 Gaussian Data -4.02 -3.14 -2.13 -0.95 -4.43 -4.30 -2.85
+28$\times$100 Branching Data -3.46 -3.58 -1.37 -1.19 -3.68 -3.59 -3.39
+28$\times$100 Coalescent Data -4.41 -3.92 -2.09 -2.00 -4.54 -4.51 -3.81
+28$\times$100 Gaussian Data -2.98 -3.30 -1.64 -1.61 -3.07 -3.04 -2.80

### v3

@@
 Abstract
-We propose a gradient descent method for solving optimization problems arising in settings of tropical geometry—a variant of algebraic geometry that has become increasingly studied in applications such as computational biology, economics, and computer science.
-Our approach takes advantage of the polyhedral and combinatorial structures arising in tropical geometry to propose a versatile approach for approximating local minima in tropical statistical optimization problems—a rapidly growing body of work in recent years.
+We propose a gradient descent method for solving optimization problems arising in settings of tropical geometry—a variant of algebraic geometry that has attracted growing interest in applications such as computational biology, economics, and computer science.
+Our approach takes advantage of the polyhedral and combinatorial structures arising in tropical geometry to propose a versatile method for approximating local minima in tropical statistical optimization problems—a rapidly growing body of work in recent years.
 Theoretical results establish global solvability for 1-sample problems and a convergence rate matching classical gradient descent.
-Numerical experiments demonstrate the method’s superior performance over classical descent for tropical optimization problems which exhibit tropical convexity but not classical convexity.
-We also demonstrate the seamless integration of tropical descent into advanced optimization methods, such as Adam, offering improved overall performance.
+Numerical experiments demonstrate the method’s superior performance compared to classical gradient descent for tropical optimization problems which exhibit tropical convexity but not classical convexity.
+We also demonstrate the seamless integration of tropical descent into advanced optimization methods, such as Adam, offering improved overall accuracy.
 Keywords:
@@
 Evaluating polynomials with these operations results in piecewise linear functions.
-Tropical geometry is a relatively young field of pure mathematics established in the 1990s (with roots dating back further) and has recently become an area of active interest in computational and applied mathematics for its relevance to the statistical analyses of phylogenetic trees (Monod et al.,, 2018), the problem of dynamic programming in computer science (Maclagan & Sturmfels,, 2015), and in mechanism design for two-player games in game theory (Lin & Tran,, 2019).
-Following the identification of the tropical Grassmannian and the space of phylogenetic trees (Speyer & Sturmfels,, 2004) in particular, there has been surge of research (Barnhill et al.,, 2024b; Lin & Yoshida,, 2018; Comăneci & Joswig,, 2024; Monod et al.,, 2018) which necessitates the development of new computational techniques for optimisation in tropical geometric settings.
-The training of machine learning models and the vast majority of computational tasks in statistics entail solving optimization problems; this paper is motivated by statistical optimization problems in tropical geometric settings.
-Initial research into tropical statistics considered optimization problems such as the identification of Fermat–Weber points which, due to the piecewise linear nature of tropical geometry, can be re-framed as a linear program (Comăneci & Joswig,, 2024; Lin & Yoshida,, 2018).
-Unfortunately, the complexity of these linear programs often scales with the size of our dataset, so gradient descent is more commonly implemented in practice (Barnhill et al.,, 2024a, b).
-More recently—following the identification of ReLU neural networks and tropical rational functions (Zhang et al.,, 2018)—more sophisticated neural network optimization problems are being studied in the tropical setting (Pasque et al.,, 2024; Yoshida et al.,, 2024), which necessitates the use of gradient methods for optimization in the tropical setting.
+Tropical geometry is a relatively young field of pure mathematics established in the 1990s (with roots dating back further) and has recently become an area of active interest in computational and applied mathematics for its relevance to the statistical analyses of phylogenetic trees (monod2018StatPerspective), the problem of dynamic programming in computer science (maclagan2015introduction), and in mechanism design for two-player games in game theory (LIN2019133).
+Following the identification of the tropical Grassmannian and the space of phylogenetic trees (speyer2004grassmannian) in particular, there has been a surge of research (barnhill2023tml; lin2018FermatWeber; comuaneci2023transportation; monod2018StatPerspective) which necessitates the development of new computational techniques for optimization in tropical geometric settings.
+Training machine learning models and many computational tasks in statistics entail solving optimization problems; this paper is motivated by statistical optimization problems in tropical geometric settings.
+Early research in tropical statistics focused on optimization problems such as the identification of Fermat–Weber points which, due to the piecewise linear nature of tropical geometry, can be re-framed as a linear program (comuaneci2023transportation; lin2018FermatWeber).
+Unfortunately, the complexity of these linear programs often scales poorly with the size of our dataset, so gradient methods are more commonly implemented in practice (barnhill2024polytropes; barnhill2023tml).
+More recently—following the identification of ReLU neural networks and tropical rational functions (zhang2018tropical)—more sophisticated neural network optimization problems are being studied in the tropical setting (pasque2024tropDBs; yoshida2023tropNNs), which necessitates the use of gradient methods for optimization in the tropical setting.
 In this paper, we address the natural question—given the need for gradient methods in tropical data science, how can we tailor classical gradient methods to the tropical setting?
@@
 The remainder of this paper is organized as follows.
-The following section contains the necessary foundations of tropical geometry, an introduction to tropical location problems (Comăneci,, 2024), and the array of statistical optimization tasks studied in this paper.
-Section 3 contains the theoretical contributions of this paper; we formulate steepest descent with respect to the tropical norm and present the theoretical guarantees for tropical descent, most notably its convergence for a wide class of tropically quasi-convex functions and a convergence rate of $O(1/\sqrt{k})$.
+The following section contains the necessary foundations of tropical geometry, an introduction to tropical location problems (comuaneci2024location), and an overview of the statistical optimization tasks studied in this paper.
+Section 3 contains the theoretical contributions of this paper; we formulate steepest descent with respect to the tropical norm and present the theoretical guarantees for tropical descent, most notably its convergence for a wide class of tropically quasi-convex functions and a convergence rate of $O(1/\sqrt{m})$.
 In Section 4, we perform a comprehensive numerical study of our proposed tropical descent method and its performance in the statistical optimization problems of interest.
@@
 2 Background and Preliminaries
-In this section, we overview the background on tropical geometry and present the statistical optimization problems which we study in the tropical setting.
+In this section, we present a summary of tropical geometry and the statistical optimization problems which we study in the tropical setting.
 2.1 Tropical Geometry
-Here, we review the algebraic and geometric structure of the tropical projective torus, the state space for phylogenetic data and the domain of our tropical statistical optimization problems.
+We begin by reviewing the algebraic and geometric structure of the tropical projective torus, the state space for phylogenetic data and the domain of our tropical statistical optimization problems.
 Definition 1 (Tropical Algebra).
@@
 $a\boxplus b=\max\{a,b\},\quad a\odot b=a+b.$
-The additive identity is $\infty$- ∞ and the multiplicative identity is 00. Tropical subtraction is not defined; tropical division is given by classical subtraction.
+The additive identity is -$\infty$ and the multiplicative identity is 0. Tropical subtraction is not defined; tropical division is given by classical subtraction.
 The tropical algebra defined above is sometimes referred to as the max-plus algebra; we can define the min-plus algebra analogously, taking $\min\{a,b\}=:a\oplus b$ as the additive operation on $\mathbb{R}\cup\{\infty\}$.
-These are algebraically equivalent through negation.
-Unless specified, we use the max-plus convention as this is better suited for phylogenetic statistics on the tropical projective torus (Speyer & Sturmfels,, 2004).
+These are algebraically equivalent under negation.
+Unless specified, we use the max-plus convention as this is better suited for phylogenetic statistics on the tropical projective torus (speyer2004grassmannian).
 Definition 2 (Tropical Projective Torus).
-The $n$-dimensional tropical projective torus is a quotient space constructed by endowing $\mathbb{R}^{N}$ with the equivalence relation
-$\mathbf{x}\sim\mathbf{y}\Leftrightarrow\exists\,a:\>\mathbf{x}=a\odot\mathbf{y};$ (1)
+The N1N-1-dimensional tropical projective torus is a quotient space constructed by endowing $\mathbb{R}^{N}$ with the equivalence relation
+$\mathbf{x}\sim\mathbf{y}\Leftrightarrow\exists\,a{\,\in\mathbb{R}}:\>\mathbf{x}=a\odot\mathbf{y};$ (1)
 it is denoted by $\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$. The generalized Hilbert projective metric, also referred to as the tropical metric, is given by
-$d_{{\rm tr}}(\mathbf{x},\mathbf{y})=\max_{i}\{x_{i}-y_{i}\}-\min_{i}\{x_{i}-y_{i}\}=\max_{i,j}\{x_{i}-y_{i}-x_{j}+y_{j}\}.$
+$d_{{\rm tr}}(\mathbf{x},\mathbf{y})=\max_{i}{(}x_{i}-y_{i}{)}-\min_{i}{(}x_{i}-y_{i}{)}=\max_{i,j}{(}x_{i}-y_{i}-x_{j}+y_{j}{)}.$
 This metric is induced by the tropical norm, which is given by
@@
 While a single point in the tropical projective torus is given by some ray $\mathbf{x}+\mathbb{R}\mathbf{1}$, we note that there is a unique representative of this equivalence class whose coordinates sum to zero.
-In taking such representatives, we can identify the tropical projective torus $\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ and the hyperplane =$\mathcal{H}=\{\mathbf{x}\in\mathbb{R}^{N}:\sum x_{i}=0\}$.
+In taking such representatives, we can identify the tropical projective torus $\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ and the hyperplane $\mathcal{H}=\{\mathbf{x}\in\mathbb{R}^{N}:\sum x_{i}=0\}$.
 Throughout this paper, we use this homeomorphism to visualize the tropical projective torus (e.g., Figures 1 and 2).
-The tropical projective torus is the ambient space containing the tropical Grassmannian, which is equivalent to the space of phylogenetic trees (Speyer & Sturmfels,, 2004).
-This has sparked the study of tropical data science, formalizing statistical techniques which respect the tropical geometry of the state space (Monod et al.,, 2018; Yoshida,, 2022).
-The generalized Hilbert projective metric defined above is widely accepted and preferred for the development of geometric statistical tools on the tropical projective torus (Yoshida,, 2022).
-However, recent work has shown theoretical advantages to the use of an asymmetric metric (Comăneci & Joswig,, 2024).
+The tropical projective torus is the ambient space containing the tropical Grassmannian, which is equivalent to the space of phylogenetic trees (speyer2004grassmannian).
+This has sparked the study of tropical data science, formalizing statistical techniques which respect the tropical geometry of the state space (monod2018StatPerspective; yoshida2020science).
+The generalized Hilbert projective metric defined above is widely accepted and preferred for the development of geometric statistical tools on the tropical projective torus (yoshida2020science).
+However, recent work has shown theoretical advantages to the use of an asymmetric metric (comuaneci2023transportation).
 Definition 4 (Tropical Asymmetric Distance).
@@
 Figure 1: The $d_{\mathrm{tr}}$, $d_{\triangle_{\min}}$ and $d_{\triangle_{\max}}$ tropical unit balls in $\mathbb{R}^{3}/\mathbb{R}\mathbf{1}\cong\mathcal{H}=\{\sum x_{i}=0\}$.
+The solid lines show coordinate directions.
 We now define the primary geometric objects of interest in the tropical projective torus; hyperplanes, lines, and convex hulls.
@@
 The tropical hyperplane $\mathcal{H}_{\mathbf{a}}$ defined by the linear form $a_{1}\odot x_{1}\boxplus\cdots\boxplus{a_{N}\odot x_{N}}$ is given by
-$\mathcal{H}_{\mathbf{a}}=\{\mathbf{x}:\exists\,i\neq j\text{ s.t. }a_{i}+x_{i}=a_{j}+x_{j}=\max\{a_{k}+x_{k}\}\}.$
+$\mathcal{H}_{\mathbf{a}}=\{\mathbf{x}{\,\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}}:\exists\,i\neq j\text{ s.t. }a_{i}+x_{i}=a_{j}+x_{j}=\max_{k}{(}a_{k}+x_{k}{)}\}.$
 Definition 7 (Tropical Line Segment).
@@
 The dashed lines show coordinate directions.
-Definition 8 (Tropical Convex Hull (Develin & Sturmfels,, 2004)).
-A set $S\subset\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ is $\max$-tropically convex if it contains $\alpha\odot\mathbf{x}\boxplus\beta\odot\mathbf{y}$ for all $\mathbf{x},\mathbf{y}\in S$ and $\alpha,\beta\in\mathbb{R}$.
+Definition 8 (Tropical Convex Hull (develin2004tropical)).
+A set $S\subset\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ is max-tropically convex if it contains $\alpha\odot\mathbf{x}\boxplus\beta\odot\mathbf{y}$ for all $\mathbf{x},\mathbf{y}\in S$ and $\alpha,\beta\in\mathbb{R}$.
 For a finite subset $X=\{\mathbf{x}_{1},\ldots,\mathbf{x}_{{K}}\}\subset\mathbb{R}^{{N}}/\mathbb{R}\mathbf{1}$, the max-plus tropical convex hull of $X$ is the set of all max-tropical linear combinations of points in $X$,
@@
 Tropical location problems are a new avenue of research, motivated by the optimization of statistical loss functions; they involve the minimization over some dataset of some loss function which heuristically increases with distance from some kernel.
-Here we outline the relevant definitions of tropical location problems and their theoretical behavior as presented by Comăneci, (2024).
-We introduce the problem of tropical linear regression (Akian et al.,, 2023) as motivation.
+Here we outline the relevant definitions of tropical location problems and their theoretical behavior as presented by comuaneci2024location.
+We introduce the problem of tropical linear regression (akian2023tropical) as motivation.
 2.2.1 Motivating Example
 The problem of tropical linear regression looks to find the best-fit tropical hyperplane which minimizes the maximal distance to a set of data points.
-While this is an inherently statistical optimization problem, it has recently been proven to be polynomial-time equivalent to solving mean payoff games (Akian et al.,, 2023).
-A tropical hyperplane is uniquely defined by a cone point $\mathbf{t}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$, allowing us to formulate the best-fit tropical hyperplane as an optimization problem over $\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$.
-Definition 9 (Tropical Linear Regression (Akian et al.,, 2023)).
+While this is an inherently statistical optimization problem, it has recently been proven to be polynomial-time equivalent to solving mean payoff games (akian2023tropical).
+A tropical hyperplane is uniquely defined by a cone point $\mathbf{t}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$, allowing us to formulate the best-fit tropical hyperplane as an optimization problem over ${\mathbf{t}\in\,}\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$.
+Definition 9 (Tropical Linear Regression (akian2023tropical)).
 Let $X=\{\mathbf{x}_{1},\dots,\mathbf{x}_{K}\}\subset\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ be our dataset.
 The tropical linear regression problem finds the vertex $\mathbf{t}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ of the best-fit hyperplane:
-$\min_{\mathbf{t}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}}f(\mathbf{t})=\max_{k\leq K}d_{\mathrm{tr}}(\mathbf{x},\mathcal{H}_{\mathbf{t}})=\max_{k\leq K}(\mathbf{x}_{k}-\mathbf{t})_{i}-(\mathbf{x}_{k}-\mathbf{t})_{j}$
-where $i,j$ are such that $\forall\,\ell\neq i,j:(\mathbf{x}_{k}-\mathbf{t})_{i}\geq(\mathbf{x}_{k}-\mathbf{t})_{j}\geq(\mathbf{x}_{k}-\mathbf{t})_{\ell}$.
+$\min_{\mathbf{t}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}}f(\mathbf{t})=\max_{k\leq K}d_{\mathrm{tr}}(\mathbf{x}_{k},\mathcal{H}_{\mathbf{t}})=\max_{k\leq K}{\left(\textstyle\max_{i}(\mathbf{x}_{k}-\mathbf{t})_{i}-\text{2ndmax}_{i}(\mathbf{x}_{k}-\mathbf{t})_{i}\right)}$
+where $\text{2ndmax}_{i}y_{i}$ denotes some $y_{i}$ such that $\exists\,j\neq i$ such that $\forall\,k\neq i,j$: $y_{j}\geq y_{i}\geq y_{k}$.
+This expression for the tropical distance between $\mathbf{x}$ and $\mathcal{H}_{\mathbf{t}}$ was first stated by gartner2008tropical, and has been frequently used in the ongoing study of tropical support vector machines (yoshida2023tropical; tang2024tropical).
 In Figure 3, we see a heatmap of this objective function over the tropical projective plane for a sample of size 3.
@@
 The linear regression objective has sparse gradient almost everywhere, which produces valleys.
-These properties are considered unfavourable extreme cases for gradient methods, but are particularly common in tropical optimization problems.
+These properties are considered unfavorable extreme cases for gradient methods, but are particularly common in tropical optimization problems.
 While the tropical linear regression problem is not convex, it is quasi-convex along tropical line segments; that is, its sub-level sets are tropically convex.
@@
 2.2.2 Quasi-Convexity of Tropical Functions
-In formalizing tropical location problems, the work of Comăneci, (2024) first considers functions which increase along tropical geodesics from a kernel $\mathbf{v}$; that is, for any point $\mathbf{t}$ on a geodesic from $\mathbf{v}$ to $\mathbf{w}$, we have that $f(\mathbf{t})\leq f(\mathbf{w})$.
-This property is equivalent to star-convexity of sublevel sets.
+In formalizing tropical location problems, the work of comuaneci2024location first considers functions which increase along tropical geodesics from a kernel $\mathbf{v}$; that is, for any point $\mathbf{t}$ on a geodesic from $\mathbf{v}$ to $\mathbf{w}$, we have that $f(\mathbf{t})\leq f(\mathbf{w})$.
+This property is equivalent to star-convexity of sub-level sets.
 To formally state this, we define (oriented) geodesic segments and $\triangle_{\min}$-star convexity of sets.
 Definition 10 (Oriented Geodesic Segment).
-The (oriented) geodesic segment between $\mathbf{a},\mathbf{b}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$, is given by $[\mathbf{a},\mathbf{b}]_{\triangle_{\min}}\coloneqq\{\mathbf{x}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}:d_{\triangle_{\min}}(\mathbf{a},\mathbf{x})+d_{\triangle_{\min}}(\mathbf{x},\mathbf{b})=d_{\triangle_{\min}}(\mathbf{a},\mathbf{b})\}$.
+The (oriented) geodesic segment between $\mathbf{a},\mathbf{b}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$, is given by $\mathbf{a},\mathbf{b}]_{\triangle_{\min}}\coloneqq\{\mathbf{x}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}:d_{\triangle_{\min}}(\mathbf{a},\mathbf{x})+d_{\triangle_{\min}}(\mathbf{x},\mathbf{b})=d_{\triangle_{\min}}(\mathbf{a},\mathbf{b})\}$.
 The oriented geodesic segment is the union of all geodesics from $\mathbf{a}$ to $\mathbf{b}$ with respect to $d_{\triangle_{\min}}$.
 Definition 11 ($\triangle_{\min}$-Star Convex Set).
-A set $K\subseteq\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ is a $\triangle_{\min}$-star-convex set with kernel $\mathbf{v}$ if, for every $\mathbf{w}\in K$, we have $[\mathbf{v},\mathbf{w}]_{\triangle_{\min}}\subseteq K$.
-Using the above definitions, we can characterize a loss function which is increasing along tropical geodesics as a function with $\triangle_{\min}$-star-convex sublevel sets.
+A set $S\subseteq\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ is a $\triangle_{\min}$-star-convex set with kernel $\mathbf{v}$ if, for every $\mathbf{w}\in S$, we have $\mathbf{v},\mathbf{w}]_{\triangle_{\min}}\subseteq S$.
+Using the above definitions, we can characterize a loss function which is increasing along tropical geodesics as a function with $\triangle_{\min}$-star-convex sub-level sets.
 The following definition gives a closed form expression for such functions, while Theorem 13 proves their equivalence.
-Definition 12 ($\triangle_{\min}$-Star-Quasi-Convex Functions (Comăneci,, 2024)).
+Definition 12 ($\triangle_{\min}$-Star-Quasi-Convex Functions (comuaneci2024location)).
 A function $f:\mathbb{R}^{N}/\mathbb{R}\mathbf{1}\rightarrow\mathbb{R}$ is $\triangle_{\min}$-star-quasi-convex with kernel $\mathbf{v}$ if $f(\mathbf{x})=\hat{\gamma}(\mathbf{x}-\mathbf{v})$ for some $\gamma:\mathbb{R}^{N}_{\geq 0}\rightarrow\mathbb{R}$ which is increasing in every coordinate, where we define $\hat{\gamma}(\mathbf{x})\coloneqq\gamma(\mathbf{x}-(\min_{i}x_{i})\mathbf{1})$.
-Theorem 13 (Theorem 17 of (Comăneci,, 2024)).
+Theorem 13 (Theorem 17 of comuaneci2024location).
 A continuous function $f:\mathbb{R}^{N}/\mathbb{R}\mathbf{1}\rightarrow\mathbb{R}$ is $\triangle_{\min}$-star-quasi-convex with kernel $\mathbf{v}$ if and only if all of its non-empty sub-level sets are $\triangle_{\min}$-star convex with kernel $\mathbf{v}$.
@@
 We can define $\triangle_{\max}$-star-quasi-convex functions similarly, but we also note that a function $f(\mathbf{t})$ is a $\triangle_{\max}$-star-quasi-convex function if and only if $f(-\mathbf{t})$ is a $\triangle_{\min}$-star-quasi-convex function.
-These $\triangle_{\min}$-star-convex functions act as a loss function with respect to a single data point.
+These $\triangle_{\min}$-star-convex functions serve as a loss function with respect to a single data point.
 For example, the tropical metric $d_{\mathrm{tr}}$ is both $\triangle_{\min}$ and $\triangle_{\max}$-star-quasi-convex.
@@
 We limit our theoretical considerations to $\triangle_{\min}$-star-quasi-convex functions to utilize their closed form expression given in Definition 12, however numerical experiments demonstrate strong performance of our methodology for more general tropical quasi-convex functions.
-Proposition 15 (Proposition 8 of (Comăneci,, 2024)).
-Any $\triangle_{\min}$-star-convex set is $\min$-tropically convex. Hence, any $\triangle_{\min}$-star-quasi-convex function is tropically quasi-convex in that it has tropically convex sublevel sets.
+Proposition 15 (Proposition 8 of comuaneci2024location).
+Any $\triangle_{\min}$-star-convex set is $\min$-tropically convex. Hence, any $\triangle_{\min}$-star-quasi-convex function is tropically quasi-convex in that it has tropically convex sub-level sets.
 2.2.3 Tropical Location Problems
-The $\triangle_{\min}$-star-convex functions defined above are generally a measure of closeness to the kernel $\mathbf{v}$, but in minimizing a statistical loss function we look at closeness to a sample of points.
-A location problem (Laporte et al.,, 2019) measures the closeness to a sample rather than a single kernel.
-Definition 16 (Min(/max)-Tropical Location Problem (Comăneci,, 2024)).
-Let $X=\{\mathbf{x}_{1},\dots,\mathbf{x}_{K}\}\subset\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ be our dataset.
+The $\triangle_{\min}$-star-convex functions defined above are generally a measure of closeness to the kernel $\mathbf{v}$, but in minimizing a statistical loss function we consider proximity to a sample of points.
+A location problem (laporte2019introduction) measures the closeness to a sample rather than a single kernel.
+Definition 16 (Min(/max)-Tropical Location Problem (comuaneci2024location)).
+Consider a dataset $X=\{\mathbf{x}_{1},\dots,\mathbf{x}_{K}\}\subset\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$.
 For $k\leq K$, let $h_{k}$ be a $\triangle_{\min}$-star-quasi-convex function with kernel $\mathbf{x}_{k}$, and let $g:\mathbb{R}^{K}\rightarrow\mathbb{R}$ be increasing in every coordinate.
@@
 A max-tropical location problem is defined similarly, where the $h_{k}$ are $\triangle_{\max}$-star-quasi-convex.
-The main result of (Comăneci,, 2024) is the following: tropical location problems contain a minimum in the tropical convex hull of its dataset $X$.
-Theorem 17 (Theorem 20 of (Comăneci,, 2024)).
+The main result of comuaneci2024location is the following: tropical location problems contain a minimum in the tropical convex hull of its dataset $X$.
+Theorem 17 (Theorem 20 of comuaneci2024location).
 Let $f$ be the objective function of a min-tropical location problem.
@@
 Conversely, max-tropical location problems have minima belonging to $\operatorname{tconv}_{\min}(X)$.
-We can now revisit our motivating example (Definition 9), showing that tropical linear regression is not just a min-tropical location problem, but also demonstrates tropical quasi-convexity in that it has tropically convex sublevel sets.
+We can now revisit our motivating example (Definition 9), showing that tropical linear regression is not just a min-tropical location problem, but also demonstrates tropical quasi-convexity in that it has tropically convex sub-level sets.
 Proposition 18.
-The tropical linear regression problem is a $\min$-tropical location problem with tropically convex sublevel sets.
+The tropical linear regression problem is a $\min$-tropical location problem with tropically convex sub-level sets.
 Proof
-$f(\mathbf{t})$ $=\max_{k\leq K}\min_{j}[\max_{i}[x_{ki}-t_{i}]-x_{kj}+t_{j}]$
-$=\max_{k\leq K}\min_{j}[t_{j}-x_{kj}-\min_{i}[t_{i}-x_{ki}]]$
+$f(\mathbf{t})$ $={\max_{k\leq K}\left(\textstyle\max_{i}(\mathbf{x}_{k}-\mathbf{t})_{i}-\text{2ndmax}_{i}(\mathbf{x}_{k}-\mathbf{t})_{i}\right)}$
+$=\max_{k\leq K}\text{2ndmin}_{j}(\max_{i}(x_{ki}-t_{i})-x_{kj}+t_{j})$
+$=\max_{k\leq K}\text{2ndmin}_{j}(t_{j}-x_{kj}-\min_{i}(t_{i}-x_{ki}{))}$
 This is a min-tropical location problem with:
@@
 $h_{k}(\mathbf{t})$ $=\hat{\gamma}(\mathbf{t}-\mathbf{x}_{k}),$
-$\gamma(\mathbf{t})$ $=\min_{j}t_{i}.$
+$\gamma(\mathbf{t})$ $=\text{2ndmin}_{i}t_{i}.$
 This $g$ and $\gamma$ are increasing in each coordinate and each $h_{k}$ has kernel $\mathbf{x}_{k}$, satisfying the conditions for a min-tropical location problem.
-As the $h_{k}$ are $\triangle_{\min}$-star-convex, their sublevel sets are min-tropically convex by Proposition 15.
-This is preserved when taking a maximum over samples, so $f$ also has $\min$-tropically convex sublevel sets.
+As the $h_{k}$ are $\triangle_{\min}$-star-convex, their sub-level sets are min-tropically convex by Proposition 15.
+This is preserved when taking a maximum over samples, so $f$ also has $\min$-tropically convex sub-level sets.
 ∎
@@
 2.3 Tropical Data Science
-While young, the field of tropical data science is motivated by the crucial question of statistical learning from the increasing volume of phylogenetic tree data (Kahn,, 2011).
-In contrast to the BHV (Billera et al.,, 2001) or Robinson–Foulds metrics (Robinson & Foulds,, 1981), the tropical interpretation of tree space provides both interpretable geometry and computational efficiency (Steel & Penny,, 1993; Monod et al.,, 2018).
-Many of the statistical methods defined for the tropical setting are framed as optimisation problems (Barnhill et al.,, 2024b), and in this subsection we review some such tropical statistical optimization problems which we use to test our tropical descent methodology.
+While young, the field of tropical data science is motivated by the crucial question of statistical learning from the increasing volume of phylogenetic tree data (kahn2011future).
+In contrast to the BHV (billera2001geometry) or Robinson–Foulds metrics (ROBINSON1981131), the tropical interpretation of tree space provides both interpretable geometry and computational efficiency (penny1993distributions; monod2018StatPerspective).
+Many of the statistical methods defined for the tropical setting are framed as optimization problems (barnhill2023tml), and in this subsection we review some such tropical statistical optimization problems which we use to test our tropical descent methodology.
 We note the varying degrees of convexity demonstrated by each of them.
@@
 2.3.1 Centrality Statistics
-Our first problems of interest, Fermat–Weber points and Fréchet means, are centrality statistics for data on a general metric space; Fermat–Weber points (Drezner et al.,, 2004) act as a generalized median, while Fréchet means (Fréchet,, 1948) are a generalization of the mean.
-Definition 19 (Tropical Fermat–Weber Points (Lin & Yoshida,, 2018)).
+Our first problems of interest, Fermat–Weber points and Fréchet means, are centrality statistics for data on a general metric space; Fermat–Weber points (wesolowsky1993weber) act as a generalized median, while Fréchet means (frechet1948elements) are a generalization of the mean.
+Definition 19 (Tropical Fermat–Weber Points (lin2018FermatWeber)).
 Let $X=\{\mathbf{x}_{1},\dots,\mathbf{x}_{K}\}\subset\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ be our dataset. A solution to the following optimization problem is a tropical Fermat–Weber point:
 $\min_{\mathbf{t}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}}f(\mathbf{t})=\frac{1}{K}\sum_{k\leq K}d_{{\rm tr}}(\mathbf{x}_{k},\mathbf{t}).$
-Definition 20 (Tropical Fréchet means (Monod et al.,, 2018)).
+Definition 20 (Tropical Fréchet means (monod2018StatPerspective)).
 Let $X=\{\mathbf{x}_{1},\dots,\mathbf{x}_{K}\}\subset\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ be our dataset. A solution to the following optimization problem is a tropical Fréchet mean:
@@
 2.3.2 Wasserstein Projections
-The next statistical optimization problem was motivated by Cai & Lim, (2022), and looks to identify a tropical projection which minimizes the Wasserstein distance between samples in different tropical projective tori.
-Definition 21 (Tropical Wasserstein Projections (Talbut et al.,, 2023)).
+The next statistical optimization problem was motivated by cai2022distances, and looks to identify a tropical projection which minimizes the Wasserstein distance between samples in different tropical projective tori.
+Definition 21 (Tropical Wasserstein Projections (talbut2023probability)).
 Let $X=\{\mathbf{x}_{1},\dots,\mathbf{x}_{K}\}\subset\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ be a dataset in one space, and $Y=\{\mathbf{y}_{1},\dots,\mathbf{y}_{K}\}\subset\mathbb{R}^{M}/\mathbb{R}\mathbf{1}$ be a dataset in a different space.
@@
 Then the tropical $p$-Wasserstein projection problem is the minimization of the following objective function:
-$\min_{\mathbf{t}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}}f_{p}(\mathbf{t})=\left(\frac{1}{K}\sum_{k\leq K}\|[\max_{i\in J_{j}}(\mathbf{x}_{k}-\mathbf{t})_{i}]_{j}-\mathbf{y}_{k}\|_{\mathrm{tr}}^{p}\right)^{1/p}$
+$\min_{\mathbf{t}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}}f_{p}(\mathbf{t})=\left(\frac{1}{K}\sum_{k\leq K}\|(\max_{i\in J_{j}}(\mathbf{x}_{k}-\mathbf{t})_{i})_{{j\leq M}}-\mathbf{y}_{k}\|_{\mathrm{tr}}^{p}\right)^{1/p}$
 The tropical $\infty$-Wasserstein projection problem is the minimization of:
-$\min_{\mathbf{t}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}}f_{\infty}(\mathbf{t})=\max_{k\leq K}\|[\max_{i\in J_{j}}(\mathbf{x}_{k}-\mathbf{t})_{i}]_{j}-\mathbf{y}_{k}\|_{\mathrm{tr}}$
-In contrast to the work by Lee et al., (2022) which looks to solve the optimal transport problem for intrinsic measures on $\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$, this $p$-Wasserstein projection problem is computing an optimal projection for empirical samples.
+$\min_{\mathbf{t}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}}f_{\infty}(\mathbf{t})=\max_{k\leq K}\|(\max_{i\in J_{j}}(\mathbf{x}_{k}-\mathbf{t})_{i})_{{j\leq M}}-\mathbf{y}_{k}\|_{\mathrm{tr}}$
+In contrast to the work by lee2021tropical which looks to solve the optimal transport problem for intrinsic measures on $\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$, this $p$-Wasserstein projection problem is computing an optimal projection for empirical samples.
 Proposition 22.
-The tropical $p,\infty$-Wasserstein projections are min-tropical location problems, and the $\infty$-Wasserstein projection has $\min$-tropically convex sublevel sets.
+The tropical $p,\infty$-Wasserstein projections are min-tropical location problems, and the $\infty$-Wasserstein projection has $\min$-tropically convex sub-level sets.
 Proof Let $\mathbf{z}_{k}=(x_{ki}-y_{kj})_{i\leq N}$ where $j$ is such that $J_{j}$ is the unique partition set containing $i$. We then have that
-$\|[\max_{i\in J_{j}}(\mathbf{x}_{k}-\mathbf{t})_{i}]_{j}-\mathbf{y}_{j}\|_{\mathrm{tr}}$ $=\max_{b}\max_{a\in J_{b}}[z_{ka}-t_{a}]-\min_{j}\max_{i\in J_{j}}[z_{ki}-t_{i}]$
-$=\max_{a}[z_{ka}-t_{a}]+\max_{j}\min_{i\in J_{j}}[t_{i}-z_{ki}]$
-$=\max_{j}[\min_{i\in J_{j}}[t_{i}-z_{ki}-\min_{a}[t_{a}-z_{ka}]]]$
+$\|(\max_{i\in J_{j}}(\mathbf{x}_{k}-\mathbf{t})_{i})_{{j\leq M}}-\mathbf{y}_{k}\|_{\mathrm{tr}}$ $=\max_{b}\max_{a\in J_{b}}(z_{ka}-t_{a})-\min_{j}\max_{i\in J_{j}}(z_{ki}-t_{i})$
+$=\max_{a}(z_{ka}-t_{a})+\max_{j}\min_{i\in J_{j}}(t_{i}-z_{ki})$
+$=\max_{j}(\min_{i\in J_{j}}(t_{i}-z_{ki}-\min_{a}(t_{a}-z_{ka}{)))}$
 The $p,\infty$-Wasserstein projections can then be written as
-$f_{p}(\mathbf{t})$ $=\left(\frac{1}{K}\sum_{k\leq K}\left(\max_{j}[\min_{i\in J_{j}}[t_{i}-z_{ki}-\min_{a}[t_{a}-z_{ka}]]]\right)^{p}\right)^{1/p}$
-$f_{\infty}(\mathbf{t})$ $=\max_{k\leq K}\max_{j}[\min_{i\in J_{j}}[t_{i}-z_{ki}-\min_{a}[t_{a}-z_{ka}]]]$
+$f_{p}(\mathbf{t})$ $=\left(\frac{1}{K}\sum_{k\leq K}\left(\max_{j}(\min_{i\in J_{j}}(t_{i}-z_{ki}-\min_{a}(t_{a}-z_{ka}{)))}\right)^{p}\right)^{1/p}$
+$f_{\infty}(\mathbf{t})$ $=\max_{k\leq K}\max_{j}(\min_{i\in J_{j}}(t_{i}-z_{ki}-\min_{a}(t_{a}-z_{ka}{)))}$
 These are max-tropical location problems with:
@@
 The $g_{p},g_{\infty}$ and $\gamma$ are increasing in each coordinate and each $h_{k}$ has kernel $\mathbf{z}_{k}$, satisfying the conditions for a min-tropical location problem.
-As the $h_{k}$ are $\triangle_{\min}$-star-convex, their sublevel sets are min-tropically convex by Proposition 15.
-This is preserved when taking a maximum over samples, so $f_{\infty}$ also has $\min$-tropically convex sublevel sets.
+As the $h_{k}$ are $\triangle_{\min}$-star-convex, their sub-level sets are min-tropically convex by Proposition 15.
+This is preserved when taking a maximum over samples, so $f_{\infty}$ also has $\min$-tropically convex sub-level sets.
 ∎
@@
 We treat optimization over $\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ as optimization over $\mathbb{R}^{N}$ with the knowledge that $f$ is constant along $\mathbf{1}$-rays.
-We assume we can compute some gradient-like function $\nabla f:\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}$ which is satisfying Assumption 23.
+We assume we can compute some gradient-like function $\nabla f:\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}$ which satisfies Assumption 23.
 Assumption 23.
@@
 For example, if $\nabla f(\mathbf{t})$ is given by some Clarke gradient, that is, some element of the generalized Clarke derivative of $f$, then Lemma 25 shows that Assumption 23 holds.
-Definition 24 (Generalized Clarke Derivative (2.13, I.I of Kiwiel, (1985))).
+Definition 24 (Generalized Clarke Derivative (2.13, I.I of kiwiel2006methods)).
 We define the set $M_{f}(\mathbf{t})$ by
@@
 As $\gamma$ is increasing in every coordinate, $h_{k}=\hat{\gamma}(\mathbf{y}_{\ell}-\mathbf{x}_{k})$ is strictly decreasing in the $j^{th}$ coordinate if and only if $j\in\text{argmin}_{i}(y_{\ell i}-x_{ki})$.
-This is preserved through the limit as $\ell\rightarrow\infty$, so we conclude that $j\in\text{argmin}_{i}[t_{i}-x_{i}]$.
+This is preserved through the limit as $\ell\rightarrow\infty$, so we conclude that $j\in\text{argmin}_{i}(t_{i}-x_{i})$.
 ∎
-In practice, rather than computing elements of the generalized Clarke derivative, we use PyTorch automatic differentiation (Paszke et al.,, 2019).
+In practice, rather than computing elements of the generalized Clarke derivative, we use PyTorch automatic differentiation (paszke2019pytorch).
 The derivatives computed by PyTorch are dependent on the computational representation of a function, but are designed to respect the chain rule.
-We must therefore ensure that our computational representation of $g$ has non-negative derivatives in each $h_{k}$, while each $h_{k}$ is computed such that its $t_{j}$ derivative is negative only if $i\in\text{argmin}_{i}(t_{i}-x_{ki})$.
+We must therefore ensure that our computational representation of $g$ has non-negative derivatives in each $h_{k}$, while each $h_{k}$ is computed such that its $t_{j}$ derivative is negative only if $j\in\text{argmin}_{i}(t_{i}-x_{ki})$.
 For such a computational representation of $f$, Assumption 23 will hold by the chain rule.
@@
 In defining tropical descent, we refer back to the motivation behind classical gradient descent—steepest descent directions.
-Definition 26 (Steepest Descent Direction (Boyd & Vandenberghe,, 2004)).
-[…]
+Definition 26 (Steepest Descent Direction (boyd2004convex)).
+A normalized steepest descent direction at $\mathbf{t}$ with respect to the norm $\|\cdot\|$ is:
 $\mathbf{d}=\text{argmin}\{\nabla f(\mathbf{t})^{\top}\mathbf{v}:\|\mathbf{v}\|=1\}.$
-With respect to the Euclidean norm, the normalized derivative is a steepest descent direction.
-When we use a the tropical norm, the steepest descent directions are characterized by the following lemma.
+With respect to the Euclidean norm, the normalized derivative is a steepest descent direction. The steepest descent directions under the tropical norm are characterized by the following lemma.
 Lemma 27.
-Suppose $\nabla f\neq\mathbf{0}$. Then the direction $\mathbf{v}$ is a normalized tropical steepest descent direction iff
-$v_{i}=1$ $\Leftrightarrow\nabla f_{i}<0$
-$v_{i}=0$ $\Leftrightarrow\nabla f_{i}>0$
+Suppose $\nabla f\neq\mathbf{0}$. Then the direction $\mathbf{v}$ is a normalized tropical steepest descent direction (up to $\sim$ equivalence) iff
+$\nabla f_{i}<0$ $\Rightarrow v_{i}=1,$
+$\nabla f_{i}>0$ $\Rightarrow v_{i}=0.$
 Proof WLOG, we assume $0\leq v_{i}\leq 1$. Then
@@
 Throughout the rest of this paper, we use the min-tropical descent direction as convention, and will specify when the max-tropical descent direction is used instead.
-Tropical descent therefore refers to the method outlined by Algorithm 1, for some specified step size sequence $(a_{n})_{n\geq 1}$, while we refer to classical gradient descent as classical descent.
+Tropical descent therefore refers to the method outlined by Algorithm 1, for some specified step size sequence $(a_{m})_{m\geq 1}$, while we refer to classical gradient descent as classical descent.
 Algorithm 1 Tropical Descent
-for $n<$ max_steps do
-     $\mathbf{g}_{ni}\leftarrow\nabla f_{i}(\mathbf{t}_{n-1})$;
-     $\mathbf{d}_{ni}\leftarrow\mathbf{1}_{\mathbf{g}_{ni}<0}$;
-     $\mathbf{t}_{n}\leftarrow\mathbf{t}_{n-1}+a_{n}\mathbf{d}_{n}$;
-We note that over a large number of steps, Algorithm 1 can produce large $t_{ni}$ values as each coordinate is non-decreasing in $n$.
-If this results in overflow or precision errors, a normalisation operation can be included in each step, such as subtracting $\frac{1}{N}\sum_{i}t_{ni}$ from each $t_{ni}$.
+for $m<$ max_steps do
+  $\mathbf{g}_{mi}\leftarrow\nabla f(\mathbf{t}_{m-1})_{i}$;
+  $\mathbf{d}_{mi}\leftarrow\mathbf{1}_{\mathbf{g}_{mi}<0}$;
+  $\mathbf{t}_{m}\leftarrow\mathbf{t}_{m-1}+a_{m}\mathbf{d}_{m}$;
+We note that over a large number of steps, Algorithm 1 can produce large $t_{mi}$ values as each coordinate is non-decreasing in $m$.
+If this results in overflow or precision errors, a normalization operation can be included in each step, such as subtracting $\frac{1}{N}\sum_{{i\leq N}}t_{mi}$ from each $t_{mi}$.
 The function $f$ is invariant under such a translation due to the equivalence relation on the tropical projective torus.
-In (Barnhill et al.,, 2024a, b), classical descent is implemented for the computation of Fermat–Weber points; this is the most natural choice as the Fermat–Weber problem is classically convex.
+We note that barnhill2024polytropes and barnhill2023tml implement classical descent for the computation of Fermat–Weber points; this is the most natural choice as the Fermat–Weber problem is classically convex.
 However, the linear regression and Wasserstein projection problems lack classical convexity, and hence do not consistently produce accurate solutions via classical descent.
@@
 We show that tropical descent must converge to the tropical convex hull of the data, and hence will necessarily find a minimum of a $\triangle_{\min}$-star-convex problem.
-Throughout this section, we assume the step sizes $a_{n}$ satisfy $a_{n}\rightarrow 0,\sum a_{n}=\infty$.
-We first highlight the key difference in stability between classical descent and tropical descent; if some component of the derivative is locally zero, tropical descent cannot converge in that neighbourhood.
+Throughout this section, we assume the step sizes $a_{m}$ satisfy $a_{m}\rightarrow 0,\sum a_{m}=\infty$.
+We first highlight the key difference in stability between classical descent and tropical descent; if some component of the derivative is locally zero, tropical descent cannot converge in that neighborhood.
 Proposition 28.
-Let $\mathbf{t}_{n}$ be some sequence defined by tropical descent.
-Assume that for all $n$, $\nabla f(\mathbf{t}_{n})\neq\mathbf{0}$.
+Let $\mathbf{t}_{m}$ be some sequence defined by tropical descent.
+Assume that for all $m$, $\nabla f(\mathbf{t}_{m})\neq\mathbf{0}$.
 Suppose in some bounded open $U$, $\nabla f_{i}$ is identically 0.
-Then no point in $\operatorname{int}U$ is a convergence point for $\mathbf{t}_{n}$; in particular, if $\mathbf{t}_{m}\in U$ there is some $n>m$ such that $\mathbf{t}_{n}\notin U$.
+Then no point in $U$ is a convergence point for $\mathbf{t}_{m}$; in particular, if $\mathbf{t}_{m}\in U$ there is some $n>m$ such that $\mathbf{t}_{n}\notin U$.
 Proof Assume otherwise.
@@
 At each step, we increase at least one coordinate by $a_{n}$, so as $n\rightarrow\infty$:
-$\sum_{j}t_{nj}\geq\sum_{j}t_{mj}+\sum_{n\geq m}a_{n}\rightarrow\infty.$
+$\sum_{j}t_{nj}\geq\sum_{j}t_{mj}+\sum_{n\geq m^{\prime}\geq m}a_{m^{\prime}}\rightarrow\infty.$
 Therefore $\|\mathbf{t}_{n}\|_{{\rm tr}}$ will become arbitrarily large, as $\max_{j}t_{nj}\rightarrow\infty$ while $\min_{j}t_{nj}\leq t_{mi}$. Then we must have left $U$; contradiction. ∎
@@
 Suppose $f$ is a min-tropical location problem with respect to the dataset $X=\{\mathbf{x}_{1},\dots,\mathbf{x}_{K}\}$.
-Let $(\mathbf{t}_{n})_{n\geq 1}$ be a sequence of points defined by tropical descent, such that for all $n$: $\nabla f(\mathbf{t}_{n})\neq 0$.
+Let $(\mathbf{t}_{m})_{m\geq 1}$ be a sequence of points defined by tropical descent, such that for all $m$: $\nabla f(\mathbf{t}_{m})\neq 0$.
 Let $V=\operatorname{tconv}_{\max}(\mathbf{x}_{1},\dots,\mathbf{x}_{K})$, and define
-$\Delta_{j}(\mathbf{t})=\sum_{k\in[K]}\sum_{i\in[N]}[t_{j}-x_{kj}-t_{i}+x_{ki}]^{+}.$
+$\Delta_{j}(\mathbf{t})=\sum_{{k\leq K}}\sum_{{i\leq N}}[t_{j}-x_{kj}-t_{i}+x_{ki}]^{+}.$
 For any $\epsilon>0$, let $M_{1}$ be such that for all $m\geq M_{1}$, $a_{m}<\epsilon/N$.
-Let $M_{2}$ be such that for all $m\geq M_{2}$ $s_{m}\geq\max_{i}\Delta_{i}(\mathbf{t}_{M_{1}})+s_{M_{1}}-K\epsilon$, where $s_{m}$ is the partial sum of the step size sequence $a_{m}$.
+Let $M_{2}$ be such that for all $m\geq M_{2}$ $s_{m}\geq\max_{j}\Delta_{j}(\mathbf{t}_{M_{1}})+s_{M_{1}}-K\epsilon$, where $s_{m}$ is the partial sum of the step size sequence $a_{m}$.
 Then for all $m\geq M_{2}$ we have:
-$d_{{\rm tr}}(\mathbf{t}_{n},V)\leq\epsilon.$
+$d_{{\rm tr}}(\mathbf{t}_{m},V)\leq\epsilon.$
 Proof For each $i\leq N$, we define the functions
-$\delta_{j}(\mathbf{t})=\min_{k\in[K]}\left[\sum_{i\in[N]}[t_{j}-x_{kj}-t_{i}+x_{ki}]^{+}\right].$
+$\delta_{j}(\mathbf{t})=\min_{{k\leq K}}\left[\sum_{{i\leq N}}[t_{j}-x_{kj}-t_{i}+x_{ki}]^{+}\right].$
 Claim 30.
 For all $\mathbf{t}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$, we have $d_{{\rm tr}}(\mathbf{t},V)\leq\max_{j}\delta_{j}(\mathbf{t})$.
-Proof As discussed by Maclagan & Sturmfels, (2015), there is a well-defined metric projection map $\pi_{V}$ onto $V$ given by:
+Proof As discussed by maclagan2015introduction, there is a well-defined metric projection map $\pi_{V}$ onto $V$ given by:
 $\pi_{V}(\mathbf{t})=\lambda_{1}\odot\mathbf{x}_{1}\boxplus\dots\boxplus\lambda_{K}\odot\mathbf{x}_{K},$
-[…]
+ $\text{ where }\lambda_{k}=\max\{\lambda\in\mathbb{R}:\,\lambda\odot\mathbf{x}_{k}\boxplus\mathbf{t}=\mathbf{t}\}=\min_{i}(t_{i}-x_{ki}).$
 The distance from $\mathbf{t}$ to $V$ is then given by $d(\mathbf{t},\pi_{V}(\mathbf{t}))$, and we have:
-$d(\mathbf{t},V)$ $=d(\mathbf{t},\pi_{V}(t))$
-$=\|(\max_{k}(\min_{i}(t_{i}-x_{ki})-t_{j}+x_{kj})_{j\leq N}\|.$
+$d(\mathbf{t},V)$ $=d(\mathbf{t},\pi_{V}(\mathbf{t}))$
+$=\|(\max_{k}(\min_{i}(t_{i}-x_{ki})-t_{j}+x_{kj})_{j\leq N}\|_{{\rm tr}}.$
 We note that for all $j,k$, the term $\min_{i}(t_{i}-x_{ki})-t_{j}+x_{kj}$ is negative. Hence:
 $d(\mathbf{t},V)$ $\leq 0-\min_{j}(\max_{k}(\min_{i}(t_{i}-x_{ki})-t_{j}+x_{kj}))$
-$\leq\max_{j}(\min_{k}(t_{j}-x_{kj}-\min_{i}(t_{i}-x_{ki})))$
+$=\max_{j}(\min_{k}(t_{j}-x_{kj}-\min_{i}(t_{i}-x_{ki})))$
 $\leq\max_{j}\delta_{j}(\mathbf{t}).$
@@
 If $\delta_{i}(\mathbf{t}_{m})=0$ and $m\geq M_{1}$, then $\delta_{i}(\mathbf{t}_{m+1})<\epsilon$.
-Proof For all $k\in[K],j\in[N]$, we have that $t_{mj}-x_{kj}-t_{mi}+x_{ki}$ is $1$-Lipschitz in $\mathbf{t}$ (with respect to $d_{\mathrm{tr}}$).
+Proof For all $k\leq K,j\leq N$}, we have that $t_{mj}-x_{kj}-t_{mi}+x_{ki}$ is $1$-Lipschitz in $\mathbf{t}$ (with respect to $d_{\mathrm{tr}}$).
 Therefore $\delta_{j}$ is $N$-Lipschitz in $\mathbf{t}$.
 By the tropical descent rule:
-$\delta_{j}(t_{m+1})$ $\leq\delta_{j}(\mathbf{t}_{m})+Nd_{\mathrm{tr}}(\mathbf{t}_{m},\mathbf{t}_{m+1})$
+$\delta_{j}(\mathbf{t}_{m+1})$ $\leq\delta_{j}(\mathbf{t}_{m})+Nd_{\mathrm{tr}}(\mathbf{t}_{m},\mathbf{t}_{m+1})$
 $=0+Na_{m}$
@@
 Then as in the proof of the previous claim, at each step we have $\nabla f_{j}\geq 0$, $d_{mj}=0$ and $t_{mj}$ is fixed for $M_{1}\leq m\leq M_{2}$.
-We have assumed that for each $m$, $\nabla f\neq\mathbf{0}$ and so there is some $i$ such that $\nabla f_{i}<0$, and by Lemma 25, there is some $k$ such that $i\in\text{argmin}_{\ell}t_{m\ell}-v_{k\ell}$.
+We have assumed that for each $m$, $\nabla f\neq\mathbf{0}$ and so there is some $i$ such that $\nabla f_{i}<0$, and by Lemma 25, there is some $k$ such that $i\in\text{argmin}_{\ell}(t_{m\ell}-x_{k\ell})$.
 Hence:
-$\forall\,\ell:\quad t_{mj}-v_{kj}-t_{mi}+v_{ki}$ $\geq t_{mj}-v_{kj}-t_{m\ell}+v_{k\ell},$
-$\Rightarrow\quad t_{mj}-v_{kj}-t_{mi}+v_{ki}$ $\geq\delta_{j}(\mathbf{t}_{m})/N,$
+$\forall\,\ell:\quad t_{mj}-x_{kj}-t_{mi}+x_{ki}$ $\geq t_{mj}-x_{kj}-t_{m\ell}+x_{k\ell},$
+$\Rightarrow\quad t_{mj}-x_{kj}-t_{mi}+x_{ki}$ $\geq\delta_{j}(\mathbf{t}_{m})/N,$
 $>\epsilon/N,$
@@
 Contradiction. So the claim is proved. ∎
-Finally, from the claims above, we conclude that $\forall m\geq M_{2},i\in[N]$, we have $\delta_{i}(\mathbf{t}_{m})<\epsilon$, and so by our first claim, we have $d_{{\rm tr}}(\mathbf{t}_{m},V)\leq\epsilon$.
+Finally, from the claims above, we conclude that $\forall m\geq M_{2},{i\leq N}$, we have $\delta_{i}(\mathbf{t}_{m})\leq\epsilon$, and so by our first claim, we have $d_{{\rm tr}}(\mathbf{t}_{m},V)\leq\epsilon$.
 ∎
-As proven by Comăneci, (2024), tropical location problems have minima in the tropical convex hull of the data.
+As proven by comuaneci2024location, tropical location problems have minima in the tropical convex hull of the data.
 The theorem above says that according to tropical descent, only such minima can be stable.
-We cannot do much better in terms of convergence results for general location problems; they can have disconnected sublevel sets so we would not expect gradient methods to find a global minimum.
+We cannot do much better in terms of convergence results for general location problems; they can have disconnected sub-level sets so we would not expect gradient methods to find a global minimum.
 We would require some degree of convexity to guarantee global solvability, which is exactly what the following corollary gives us; tropical descent will necessarily converge to a minimum of any $\triangle_{\min}$-star-quasi-convex function.
 Corollary 34.
-Let $f(\mathbf{t})$ be a $\triangle_{\min}$-star-quasi-convex function, and let $(\mathbf{t}_{n})_{n\geq 1}$ be a sequence of points defined by tropical descent such that for all $n$: $\nabla f(\mathbf{t}_{n})\neq 0$.
-Then there is some global minimum $\mathbf{t}^{*}$ such that for all $m\leq M_{2}$, $d_{{\rm tr}}(\mathbf{t}_{m},\mathbf{t}^{*})\leq\epsilon$ where $M_{2}$ is as in Theorem 29.
+Let $f(\mathbf{t})$ be a $\triangle_{\min}$-star-quasi-convex function, and let $(\mathbf{t}_{m})_{m\geq 1}$ be a sequence of points defined by tropical descent such that for all $m$: $\nabla f(\mathbf{t}_{m})\neq 0$.
+Then there is some global minimum $\mathbf{t}^{*}$ such that for all $m\geq M_{2}$, $d_{{\rm tr}}(\mathbf{t}_{m},\mathbf{t}^{*})\leq\epsilon$ where $M_{2}$ is as in Theorem 29.
+This corollary follows from considering the 1-sample case of tropical location problems; Corollary 34 establishes convergence to the kernel of a $\triangle_{\min}$-star-quasi-convex function, which must necessarily be a global minimum.
+This convergence result holds even if $f$ does not have a unique minimum — it may be that $f$ has multiple global minima but a unique kernel, or $f$ may have a full dimensional set of global minima upon which $\nabla f(\mathbf{t})=\mathbf{0}$ and tropical descent terminates.
 As a further corollary, we note that we can achieve the same result for max-tropical location problems by using max-tropical descent directions.
@@
 Suppose $f$ is a max-tropical location problem with respect to the dataset $X=\{\mathbf{x}_{1},\dots,\mathbf{x}_{K}\}$.
-Let $(\mathbf{t}_{n})_{n\geq 1}$ be a sequence of points defined via max-tropical descent directions, such that for all $n$: $\nabla f(\mathbf{t}_{n})\neq 0$.
+Let $(\mathbf{t}_{m})_{m\geq 1}$ be a sequence of points defined via max-tropical descent directions, such that for all $m$: $\nabla f(\mathbf{t}_{m})\neq 0$.
 Let $V=\operatorname{tconv}_{\min}(\mathbf{x}_{1},\dots,\mathbf{x}_{K})$, and define
-$\Delta^{\prime}_{i}(\mathbf{t})=\sum_{k\in[K]}\sum_{j\in[N]}[t_{j}-x_{kj}-t_{i}+x_{ki}]^{+}.$
+$\Delta^{\prime}_{i}(\mathbf{t})=\sum_{{k\leq K}}\sum_{{j\leq N}}[t_{j}-x_{kj}-t_{i}+x_{ki}]^{+}.$
 For any initial $\mathbf{t}_{0}$ and $\epsilon>0$, let $M_{1}$ be such that for all $m\geq M_{1}$, $a_{m}\leq\epsilon/N$.
@@
 Then for all $m\geq M_{2}$ we have:
-$d_{{\rm tr}}(\mathbf{t}_{n},V)\leq\epsilon.$
+$d_{{\rm tr}}(\mathbf{t}_{m},V)\leq\epsilon.$
 Remark 36.
-The results above indicate convergence to $V$ or $\mathbf{t}^{*}$ in the tropical topology of $\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$, but $\mathbf{t}_{n}$ cannot not converge in the Euclidean topology of $\mathbb{R}^{N}$ due to the divergent partial sums of $\alpha_{n}$.
-We conclude by establishing the convergence rate of tropical descent in practice; we take step sizes given by $a_{m}=\alpha m^{-1/2}\|\nabla f(\mathbf{t}_{m})\|_{{\rm tr}}$, and assume derivatives have bounded tropical norm as is the case for piecewise linear 1-Lipschitz objective functions $h_{k}$.
-In this case, we prove convergence at a rate of $O(\sqrt{m})$.
+The results above indicate convergence to $V$ or $\mathbf{t}^{*}$ in the tropical topology of $\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$, but $\mathbf{t}_{m}$ cannot converge in the Euclidean topology of $\mathbb{R}^{N}$ due to the divergent partial sums of $a_{m}$.
+We conclude by establishing a foundational bound for the convergence rate of tropical descent; we take step sizes given by $a_{m}=\alpha m^{-1/2}\|\nabla f(\mathbf{t}_{m})\|_{{\rm tr}}$, and assume derivatives have bounded tropical norm as is the case for piecewise linear 1-Lipschitz objective functions (with finitely many linear pieces).
+In this case, we prove convergence at a rate of $O(1/\sqrt{m})$.
 Proposition 37.
 Suppose $f$ is a min-tropical location problem as in Theorem 29.
-Assume that there is some $L\in\mathbb{N}$ such that for all $\mathbf{t}_{m}$, $1/L\leq\|\nabla f(\mathbf{t}_{m})\|_{{\rm tr}}\leq 2$, and assume further there is some $D$ such that $\forall k\leq K,n\in\mathbb{N}:\,d_{{\rm tr}}(\mathbf{t}_{n},\mathbf{x}_{k})\leq D$.
+Assume that there is some $L\in\mathbb{N}$ such that for all $\mathbf{t}_{m}$, $1/L\leq\|\nabla f(\mathbf{t}_{m})\|_{{\rm tr}}\leq 2$, and assume further there is some $D$ such that $\forall m\in\mathbb{N}:\,\sum_{k\leq K}d_{\triangle_{\max}}(\mathbf{x}_{k},\mathbf{t}_{m})\leq KD$.
 We define $a_{m}=\alpha m^{-1/2}\|\nabla f(\mathbf{t}_{m})\|_{{\rm tr}}$.
 Then for all $m$ satisfying
-$1<\xi(m)\coloneqq\frac{\sqrt{m+1}}{2}-\frac{KDNL(N-1)}{8\alpha}+\sqrt{\left(\frac{\sqrt{m+1}}{2}-\frac{KDNL(N-1)}{8\alpha}\right)^{2}+LKN-1},$
+${m\geq\max\left\{\left(\sqrt{2}+\frac{KDL}{2\alpha}-\sqrt{LKN-1}\right)^{2}-1,\left(\frac{2\alpha N}{D}\right)^{2}\right\},}$ (2)
 we have that:
-$d_{{\rm tr}}(\mathbf{t}_{m},V)\leq\frac{2\alpha N}{\sqrt{m+1}-KDNL(N-1)/4\alpha-1}=\epsilon_{m}.$
+$d_{{\rm tr}}(\mathbf{t}_{m},V)\leq{\frac{2\sqrt{2}\alpha N}{\sqrt{m+1}-KDL/2\alpha+\sqrt{LKN-1}-\sqrt{2}}}=\epsilon_{m}.$
 In particular, if $f$ is a $\triangle_{\min}$-star-quasi-convex function, then there is some global minimum $\mathbf{t}^{*}$ such that
-$d_{{\rm tr}}(\mathbf{t}_{m},\mathbf{t}^{*})\leq\frac{2\alpha N}{\sqrt{m+1}-DNL(N-1)/4\alpha-1}.$
-Proof We fix $m$ and define:
-$M_{1}=\left\lfloor\xi(m)^{2}\right\rfloor.$
-We note that:
-$0\leq\sqrt{M_{1}}\leq\xi(m)$
-So $\sqrt{M_{1}}$ lies between the roots of the polynomial
-$p(t)=\frac{N}{L}t^{2}+\left(\frac{KDN^{2}(N-1)}{4\alpha}-\frac{N\sqrt{m+1}}{L}\right)t+\frac{N-LKN^{2}}{L}.$
+$d_{{\rm tr}}(\mathbf{t}_{m},\mathbf{t}^{*})\leq{\frac{2\sqrt{2}\alpha N}{\sqrt{m+1}-DL/2\alpha+\sqrt{LN-1}-\sqrt{2}}=\epsilon^{\prime}_{m}.}$
+In order to use Theorem 29, we will need to find some $M_{1}$ such that $a_{M_{1}}<\epsilon/N$ (our step size is small enough) and $s_{m}\geq\max_{j}\Delta_{j}(\mathbf{t}_{M_{1}})+s_{M_{1}}-K\epsilon$ (we have taken enough steps to get close to $V$).
+Our condition that $m$ is large enough such that $\sqrt{m+1}-KDL/2\alpha+\sqrt{LKN-1}\geq\sqrt{2}$ ensures that such an $M_{1}$ exists, while $\sqrt{m}\geq 2\alpha N/D$ ensures that $M_{1}\leq m$.
+We can then take the largest such $M_{1}$ to ensure that the corresponding $\epsilon$ is as tight as possible.
+Proof We fix $m$ satisfying 2 and define:
+${\xi(m)}$ ${\coloneqq\frac{\sqrt{m+1}}{2}-\frac{KDL}{4\alpha}+\sqrt{\left(\frac{\sqrt{m+1}}{2}-\frac{KDL}{4\alpha}\right)^{2}+LKN-1},}$
+$M_{1}$ $\coloneqq\left\lfloor\xi(m)^{2}\right\rfloor.$
+We first note that by the quadratic mean inequality and as $\sqrt{m+1}-KDL/2\alpha+\sqrt{LKN-1}\geq\sqrt{2}$:
+$\xi(m)$ $=\frac{\sqrt{m+1}}{2}-\frac{KDL}{4\alpha}+\sqrt{\left(\frac{\sqrt{m+1}}{2}-\frac{KDL}{4\alpha}\right)^{2}+LKN-1}$
+$\geq\frac{\sqrt{m+1}}{2}-\frac{KDL}{4\alpha}+\frac{1}{\sqrt{2}}\left(\frac{\sqrt{m+1}}{2}-\frac{KDL}{4\alpha}+\sqrt{LKN-1}\right)$
+$\geq\frac{1}{\sqrt{2}}\left(\sqrt{m+1}-\frac{KDL}{2\alpha}+\sqrt{LKN-1}\right)\geq 1.$
+Therefore we have that
+${1}\leq\sqrt{M_{1}}\leq\xi(m),$
+and so $\sqrt{M_{1}}$ lies between the roots of the polynomial
+$p(t)=L^{-1}t^{2}+\left(KD/2\alpha-L^{-1}\sqrt{m+1}\right)t+L^{-1}-KN$}.
 Therefore:
-$\frac{N}{L}(M_{1}+1)+\left(\frac{KDN^{2}(N-1)}{4\alpha}-\frac{N\sqrt{m+1}}{L}\right)\sqrt{M_{1}}-KN^{2}\leq 0,$
-$\frac{N\sqrt{M_{1}}\sqrt{M_{1}+1}}{L}+\frac{KDN^{2}(N-1)\sqrt{M_{1}}}{4\alpha}-\frac{N\sqrt{M_{1}}\sqrt{m+1}}{L}\leq KN^{2},$
-$\frac{KDN(N-1)}{2}+2\alpha L^{-1}(\sqrt{M_{1}+1}-\sqrt{m+1})\leq\frac{2\alpha KN}{\sqrt{M_{1}}}.$
+${L^{-1}}(M_{1}+1)+\left({KD/2\alpha}-{L^{-1}}\sqrt{m+1}\right)\sqrt{M_{1}}-KN\leq 0,$
+${L^{-1}}\sqrt{M_{1}}\sqrt{M_{1}+1}+{KD\sqrt{M_{1}}/2\alpha}-{L^{-1}}\sqrt{M_{1}}\sqrt{m+1}\leq KN,$
+${KD}+2\alpha L^{-1}(\sqrt{M_{1}+1}-\sqrt{m+1})\leq\frac{2\alpha KN}{\sqrt{M_{1}}}.$
+We now show that $M_{1}\leq m$. Suppose otherwise; then by the inequality above:
+$KD$ $\leq\frac{2\alpha KN}{\sqrt{M_{1}}}+2\alpha L^{-1}(\sqrt{m+1}-\sqrt{M_{1}+1})$
+$<\frac{2\alpha KN}{\sqrt{m}}$
+This contradicts our assumption that $m\geq(2\alpha N/D)^{2}$, so we conclude that $M_{1}\leq m$.
 Claim 38.
-For all $n\geq M_{1}$ and $m$ such that $1<\xi(m)$, we have $a_{n}\leq\epsilon_{m}/N$.
-Proof We have assumed that $1<\xi(m)$, so:
+For all $n\geq{m}$ we have $a_{n}\leq\epsilon_{m}/N$.
+Proof We have shown that $\xi(m)\geq 1$, so:
 $M_{1}$ $\geq\xi(m)^{2}-1,$
 $\geq\left(\xi(m)-1\right)^{2},$
-$\sqrt{M_{1}}$ $\geq\xi(m)-1,\geq(\sqrt{m+1}-KDNL(N-1)/4\alpha)-1.$
-Hence, by the step size definition:
+$\sqrt{M_{1}}$ $\geq\xi(m)-1,$
+$\geq{\frac{1}{\sqrt{2}}\left(\sqrt{m+1}-\frac{KDL}{2\alpha}+\sqrt{LKN-1}\right)-1}.$
+Hence, by the step size definition and the fact that $n\geq m\geq M_{1}$:
 $a_{n}$ $\leq\frac{2\alpha}{\sqrt{M_{1}}},$
-$\leq\frac{2\alpha}{\sqrt{m+1}-KDNL(N-1)/4\alpha-1}=\epsilon_{m}/N.$
+$\leq{\frac{2\sqrt{2}\alpha}{\sqrt{m+1}-KDL/2\alpha+\sqrt{LKN-1}-\sqrt{2}}}=\epsilon_{m}/N.$
 The claim is proved. ∎
 Claim 39.
-For all $n\geq m$, we have $\max_{j}\Delta_{j}(\mathbf{t}_{M_{1}})+s_{M_{1}}-s_{m}\leq K\epsilon_{m}$.
+We have $\max_{j}\Delta_{j}(\mathbf{t}_{M_{1}})+s_{M_{1}}-s_{m}\leq K\epsilon_{m}$.
 Proof We note that
-$\max_{j}\Delta_{j}(\mathbf{t}_{M_{1}})$ $\leq\sum_{j}\Delta_{j}(\mathbf{t}_{M_{1}}),$
-$=\sum_{k\leq K}\sum_{i\leq j\leq N}|t_{M_{1}j}-x_{kj}-t_{M_{1}i}+x_{ki}|,$
-$\leq\sum_{k\leq K}\sum_{i\leq j\leq N}d_{{\rm tr}}(\mathbf{t}_{M_{1}},\mathbf{x}_{k}),$
-$\leq\frac{KDN(N-1)}{2}.$
-The partial sums are bounded by:
+$\Delta_{j}(\mathbf{t}_{M_{1}})$ $=\sum_{k\leq K}\sum_{i\leq N}[t_{M_{1}j}-x_{kj}-t_{M_{1}i}+x_{ki}]^{+},$
+$\leq\sum_{k\leq K}d_{\triangle_{\max}}(\mathbf{x}_{k},\mathbf{t}_{M_{1}}),$
+$\leq KD.$
+And therefore we have $\max_{j}\Delta_{j}(\mathbf{t}_{M_{1}})\leq KD$. The partial sums are bounded by:
 $s_{M_{1}}-s_{m}$ $\leq\sum_{n=M_{1}+1}^{m}-a_{n}$
@@
 $\leq\sum_{n=M_{1}+1}^{m}-2\alpha L^{-1}(\sqrt{n+1}-\sqrt{n})$
-$\leq-2\alpha L^{-1}(\sqrt{m+1}-\sqrt{M_{1}+1})$
+$\leq 2\alpha L^{-1}{(\sqrt{M_{1}+1}-\sqrt{m+1})}.$
 Therefore:
-$\max_{i}\Delta_{i}(\mathbf{t}_{M_{1}})+s_{M_{1}}-s_{m}$ $\leq KDN(N-1)/2+2\alpha L^{-1}(\sqrt{M_{1}+1}-\sqrt{m+1}),$
-$\leq\frac{2\alpha KN}{\sqrt{M_{1}}},$
-$\leq K\epsilon_{m}.$
+$\max_{i}\Delta_{i}(\mathbf{t}_{M_{1}})+s_{M_{1}}-s_{m}$ $\leq{KD}+2\alpha L^{-1}(\sqrt{M_{1}+1}-\sqrt{m+1}),$
+$\leq\frac{2\alpha KN}{\sqrt{M_{1}}}\leq K\epsilon_{m}.$
 Hence by Theorem 29, we have that
@@
 The bound for the tropically quasi-convex case follows from Corollary 34. ∎
+Our error bound in Proposition 12 is a useful foundational result on the convergence behavior of tropical descent; for sufficiently nice problems, tropical descent will converge at a rate of $O(1/\sqrt{m})$.
+However, a more detailed study of the numerical analysis of tropical descent would be valuable.
+We outline possible questions below, as well as presenting a comparison of our error bounds with tropical descent for tropical linear regression.
+Firstly, the constant term $LD/2\alpha$ in the denominator of $\epsilon_{m}$ will typically be larger than necessary, resulting in a weaker error bound.
+In particular, this constant term is quick to dominate for larger problems.
+Further work improving these constant factors would produce much tighter and more robust error bounds.
+Secondly, the error $\epsilon_{m}$ bounds the distance to a tropical convex hull $V$, which is rarely a useful quantity to bound.
+We would rather use $\epsilon^{\prime}_{m}$, which bounds the distance to a global minimum, but our tropical location problems are not known to be $\triangle_{\min}$-star-quasi-convex.
+It would be particularly useful to generalize our error bound $\epsilon_{m}^{\prime}$ to such problems which exhibit some local (but not global) $\triangle_{\min}$-star-quasi-convexity.
+For a qualitative view on Proposition 37, we compare $\epsilon_{m}^{\prime}$ and the convergence rate of tropical descent for tropical linear regression.
+As noted, tropical linear regression is tropically quasi-convex (Proposition 18) but not known to be $\triangle_{\min}$-star-quasi-convex, so does not meet the assumptions of Proposition 37; despite this, we see $\epsilon_{m}^{\prime}$ is often an effective upper bound on the error in $\mathbf{t}$.
+Figure 5: The mean log error bound ${\rm log}\,\epsilon_{m}^{\prime}$ and the mean log tropical distance to the estimated minimum $\mathbf{t}^{*}$ when applying tropical descent (TD) with $\alpha=1$ to solve the tropical linear regression problem.
+For each dataset, we take 50 random initialization points and compute ${\rm log}\,\epsilon_{m}^{\prime}$ and ${\rm log}\,d_{\mathrm{tr}}(\mathbf{t}_{m},\mathbf{t}^{*})$ independently, then take an average.
+We also record the 10th and 90th percentile values to estimate an 80% confidence interval.
+We perform tropical descent for tropical linear regression using 12 different datasets; these are generated via branching processes, coalescent processes, or a Gaussian distribution on $\mathbb{R}^{6}/\mathbb{R}\mathbf{1}$ and $\mathbb{R}^{28}/\mathbb{R}\mathbf{1}$.
+We consider sample sizes of 10 or 100.
+These are the same datasets used in Section 4, and further details on their generation can be found in Appendix B.
+We run computations using a learning rate of $\alpha=1$, for steps up to $m\leq 3000$.
+The tropical linear regression problem has $\|\nabla f(\mathbf{t}_{m})\|=2$ almost everywhere, so we take $L=1$ when computing $\epsilon_{m}^{\prime}$.
+We estimate $\mathbf{t}^{*}$ to be the mean of the final 10 steps of tropical descent (TD), and compute a tight estimate for $D$ by taking
+$D\approx\max_{m\leq 3000}d_{\triangle_{\max}}(\mathbf{t}^{*},\mathbf{t}_{m}).$
+Figure 5 shows the mean log error bound $\epsilon_{m}^{\prime}$ with the mean log tropical distance to $\mathbf{t}^{*}$.
+Examining Figure 5, we first note that the error $\epsilon_{m}^{\prime}$ is orders of magnitude smaller for the lower dimensional problems, which is primarily a result of the constant terms in the denominator of $\epsilon_{m}^{\prime}$.
+As we have noted, it would be of significant value to reduce these constant terms in future numerical analytical studies of tropical descent.
+These constant terms are also the cause of noise in the mean error bound for high dimensional problems, as we require more steps for the error to be well-defined.
+We also note that although $\epsilon_{m}^{\prime}$ is a reasonable upper bound for the distance to $\mathbf{t}^{*}$ in many cases, this upper bound does not hold for the low dimensional coalescent datasets.
+It is unclear whither this is due to poor estimates for $D$ and $\mathbf{t}^{*}$, or due to tropical linear regression not being $\triangle_{\min}$-star-quasi-convex.
+We conclude by noting that we are observing a convergence rate of $O(1/\sqrt{m})$ for most experiments, with this rate appearing particularly distinctly for branching process and Gaussian data.
 4 Numerical Experiments
-In this section we compare the practical performance of classical descent, tropical descent, SGD, tropical SGD, Adam, Adamax, and TrAdamax, with a focus on their stability behavior.
+In this section we compare the practical performance of classical descent, tropical descent, SGD, tropical SGD, Adam, Adamax, and tropical Adamax, with a focus on their stability behavior.
 The implementation for all these experiments is available in the GitHub repository at https://github.com/Roroast/TropicalGradDescent.
@@
 We take a heuristic approach to the tropicalization of SGD and Adam, replacing derivatives with tropical steepest descent directions in each case.
-Steepest Descent. We will perform classical descent (CD) and tropical descent (TD) using equivalent step size sequences $a_{n}$ for both methods, which we fix as
-$a_{n}=\frac{\gamma\|\nabla f(\mathbf{t}_{n})\|}{\sqrt{n}},$
-[…]
+Steepest Descent. We will perform classical descent (CD) and tropical descent (TD) using equivalent step size sequences $a_{m}$ for both methods, which we fix as
+$a_{m}=\frac{\gamma\|\nabla f(\mathbf{t}_{m})\|}{\sqrt{m}},$
+where $\|\cdot\|$ is the Euclidean norm for classical descent and the tropical norm for tropical descent.
+When $\nabla f(\mathbf{t})$ is bounded this is a well-behaved step rule; $a_{m}$ will converge to 0.
+Furthermore, for the piecewise linear problems, $\|\nabla f(\mathbf{t}_{m})\|$ is bounded away from 0 almost everywhere which is not a global minimum, so $\sum a_{m}=\infty$.
 Stochastic Gradient Descent.
-Stochastic gradient descent is a variant of gradient descent which looks to improve computational speed by approximating $\nabla f(t)$; rather than computing the loss and its derivative at $\mathbf{t}$ with respect to the entire dataset $f(\mathbf{t};X)$, we evaluate the loss and its derivative with respect to a random sample $\mathbf{x}_{n}\in X$, $f(\mathbf{t};\mathbf{x}_{n})$.
-This is simplified with the notation $f_{n}(\mathbf{t})$.
-As with gradient descent, we use the step size sequence give by
-$a_{n}=\frac{\gamma\|\nabla f_{n}(\mathbf{t}_{n})\|}{\sqrt{n}},$
-[…]
+Stochastic gradient descent is a variant of gradient descent which looks to improve computational speed by approximating $\nabla f(\mathbf{t})$; rather than computing the loss and its derivative at $\mathbf{t}$ with respect to the entire dataset $f(\mathbf{t};X)$, we evaluate the loss and its derivative with respect to a random sample $\mathbf{x}_{m}\in X$, $f(\mathbf{t};\mathbf{x}_{m})$.
+This is simplified with the notation $f_{m}(\mathbf{t})$.
+As with gradient descent, we use the step size sequence given by
+$a_{m}=\frac{\gamma\|\nabla f_{m}(\mathbf{t}_{m})\|}{\sqrt{m}},$
+where $\|\cdot\|$ is the Euclidean norm for stochastic gradient descent (SGD) and the tropical norm for tropical stochastic gradient descent (TSGD).
 Adam Variants.
-The Adam algorithm (Kingma & Ba,, 2015) uses gradient momentum to compound consistent but small effects.
+The Adam algorithm (kingma2017adam) uses gradient momentum to compound consistent but small effects.
 This is done by taking first and second moments of the gradient over all steps using exponentially decaying weights.
 The Adamax algorithm is a variant of the Adam algorithm motivated by the use of the $\infty$-moments rather than second moments.
+To adapt the Adam algorithm for the tropical setting, we take moments of an unnormalized tropical descent direction.
+When looking to take higher moments of the gradient over $m$, the $L_{\infty}$ is the natural option for the tropical setting.
+We therefore look to use the Adamax recursion relation on the $\infty$-moment estimate $\mathbf{u}_{m}$:
+$\mathbf{u}_{m+1}=\max(\beta_{2}\mathbf{u}_{m},\mathbf{d}_{m}).$
+The tropicalized Adamax algorithm is then given by Algorithm 2.
 Algorithm 2 TrAdamax
-$\mathbf{m},\mathbf{u}\leftarrow\mathbf{0}$;
-for $n<$ max_steps do
-     $\mathbf{g}_{ni}\leftarrow\nabla f_{i}(\mathbf{t}_{n-1})$;
-     $\mathbf{d}_{ni}\leftarrow(\max_{j}(\mathbf{g}_{nj})-\min_{j}(\mathbf{g}_{nj}))*\mathbf{1}_{\mathbf{g}_{ni}>0}$;
-     $\mathbf{m}_{n}\leftarrow\beta_{1}*\mathbf{m}_{n-1}+(1-\beta_{1})*\mathbf{d}_{n}$;
-     $\mathbf{u}_{n}\leftarrow\max(\beta_{2}*\mathbf{u}_{n-1},|\mathbf{d}_{n}|)$;
-     $\mathbf{t}_{n}\leftarrow\mathbf{t}_{n}-\alpha/(1-\beta_{1}^{n})*\mathbf{m}_{n}/\mathbf{u}_{n}$;
-To adapt the Adam algorithm for the tropical setting, we take moments of an unnormalized tropical descent direction.
-When looking to take higher moments of the gradient over $n$, the $L_{\infty}$ is the natural option for the tropical setting.
-We therefore look to use the Adamax recursion relation on the $\infty$-moment estimate $\mathbf{u}_{n}$:
-$\mathbf{u}_{n+1}=\max(\beta_{2}\mathbf{u}_{n},\mathbf{d}_{n}).$
-The tropicalized Adamax algorithm is then given by Algorithm 2.
+${\mathbf{v}},\mathbf{u}\leftarrow\mathbf{0}$;
+for $m<$ max_steps do
+  $\mathbf{g}_{mi}\leftarrow\nabla f(\mathbf{t}_{m-1})_{i}$;
+  $\mathbf{d}_{mi}\leftarrow(\max_{j}(\mathbf{g}_{mj})-\min_{j}(\mathbf{g}_{mj}))*\mathbf{1}_{\mathbf{g}_{mi}>0}$;
+  ${\mathbf{v}_{m}}\leftarrow\beta_{1}*{\mathbf{v}_{m-1}}+(1-\beta_{1})*\mathbf{d}_{m}$;
+  $\mathbf{u}_{m}\leftarrow\max(\beta_{2}*\mathbf{u}_{m-1},|\mathbf{d}_{m}|)$;
+  $\mathbf{t}_{m}\leftarrow\mathbf{t}_{m}-\alpha/(1-\beta_{1}^{m})*{\mathbf{v}_{m}}/\mathbf{u}_{m}$;
 4.2 Computation
 We now look to solve our centrality statistics, Wasserstein projection problems, and tropical linear regression using the 7 gradient methods discussed: classical descent (CD), tropical descent (TD), stochastic gradient descent (SGD), tropical stochastic gradient descent(TSGD), Adam, Adamax, and TrAdamax.
-We run computations for 12 datasets of different sizes, shapes and dimensions; motivated by phylogenetic data science applications, we consider datasets generated by branching processes, a coalescent processes, or Gaussian distributions on $\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$, for $N=6,28$ and samples of size 10 or 100.
-We take same random initializations for each gradient method, taken from a Gaussian distribution on $\mathbb{R}^{N}$.
+We run computations for 12 datasets of different sizes, shapes and dimensions; motivated by phylogenetic data science applications, we consider datasets generated by branching processes, coalescent processes, or Gaussian distributions on $\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$, for $N=6,28$ and samples of size 10 or 100.
+We take the same random initializations for each gradient method, taken from a Gaussian distribution on $\mathbb{R}^{N}$.
 Appendix B provides further details on the simulated datasets and computational methodology used.
@@
 Rather than investigating the relative convergence rates of each method, we are interested in the relative stability of local minima.
-We therefore plot the CDF of the relative log error after 1000 steps over 50 random initializations, which can highlight certain error values at which several initializations have accumulated.
-When the CDF of one method dominates another, it has a higher probability of achieving the same error and is therefore a more suitable method.
+We therefore plot the CDF of the relative log error after 1000 steps over 50 random initializations, to reveal concentrations of error values where multiple initializations stabilize.
+When the CDF of one method dominates another, it has a higher probability of achieving the same error or less and is therefore a more suitable method.
 As we will observe, tropical gradient methods are stable at fewer points than classical gradient methods; as a result, tropical methods typically have further to travel before reaching a stable point.
 Over few steps, this would be compensated for by selecting larger learning rates for tropical gradient methods relative to classical gradient methods, leading to less precise results in the case of tropical gradient methods.
-Our computations are run over 1000 steps, as this gives enough time for tropical gradient methods to stabilise, given our initialisations, without sacrificing learning rate; we see in Appendix C that neither classical nor tropical learning rates are consistently greater than the other.
+Our computations are run over 1000 steps, as this gives enough time for tropical gradient methods to stabilize, given our initializations, without sacrificing learning rate; we see in Appendix C that neither classical nor tropical learning rates are consistently greater than the other.
 4.2.1 Centrality Statistics
-We first consider the behavior of tropical descent methods for centrality statistics, Fermat–Weber points and Fréchet means.
+We first consider the behavior of tropical descent methods for Fermat–Weber points and Fréchet means.
 These optimization problems are classically convex problems, as well as being both $\max$-tropical and $\min$-tropical location problems, and therefore lend themselves well to gradient methods.
 The experimental behavior of these problems is similar, so in this section we include the results for the Fermat–Weber problem while the results for Fréchet means can be found in Appendix E.
-Figure 5 shows the CDF of the log relative error over 1000 steps of each gradient method for the 12 datasets.
+Figure 6 shows the CDF of the log relative error after 1000 steps of each gradient method for the 12 datasets.
 We note that the deterministic methods achieve similar results for most datasets, while the stochastic methods achieve the worst error for all 12 datasets and TSGD is performing worse than SGD.
 The disparity between SGD and TSGD is greatest when the dimensionality of the data is greater than the size of the dataset, and for coalescent data.
-We see similar behavior in the log error of Tropical Descent and TrAdamax; they perform on par with their classical counterparts for each dataset other than the 28 by 10 coalescent data, for which the tropical gradient methods are not achieving the same accuracy.
-While it is unclear why this is, we note that coalescent data is supported on the space of ultrametric trees, which is a tropical hyperplane (Ardila & Klivans,, 2006; Page et al.,, 2020) and may influence the uniqueness of minima.
-Figure 5: The CDF of the log relative error after 1000 steps for each gradient method across 50 initializations when minimizing the Fermat–Weber objective.
-Each sub-figure corresponds to a dataset of different dimensionality (columns) and distribution (rows).
-Table 1: The mean log relative error of each gradient method after 1000 steps for each dataset when minimizing the Fermat–Weber objective function.
+We see similar behavior in the log error of tropical descent and TrAdamax; they perform on par with their classical counterparts for each dataset other than the 28 by 10 coalescent data, for which the tropical gradient methods are not achieving the same accuracy (Table 1).
+While it is unclear why this is, we note that coalescent data is supported on the space of ultrametric trees, which is a tropical hyperplane (ardila2006bergman; page2020tropical) and may influence the stability of minima.
+Figure 6: The CDF of the log relative error after 1000 steps for each gradient method with tuned learning rates across 50 random (Gaussian) initializations when minimizing the Fermat–Weber objective.
+Each subfigure corresponds to a dataset with sample size $K=10,100$ of dimensionality $N=6,28$, sampled from a branching process, coalescent process, or a Gaussian distribution.
+Table 1: The mean log relative error of each gradient method after 1000 steps with tuned learning rates for branching process, coalescent process, and Gaussian datasets of size $K=10,100$ in $\mathbb{R}^{6}/\mathbb{R}\mathbf{1}$, $\mathbb{R}^{28}/\mathbb{R}\mathbf{1}$ when minimizing the Fermat–Weber objective function.
 The minimal mean errors for each dataset are in bold.
 Data CD TD SGD TSGD Adam Adamax TrAdamax
-6$\times$10 Branching Data -4.60 -4.60 -4.50 -3.32 -4.60 -4.60 -4.60
-6$\times$10 Coalescent Data -4.58 -4.55 -4.40 -1.28 -4.57 -4.57 -4.57
-6$\times$10 Gaussian Data -4.60 -4.60 -4.44 -2.75 -4.60 -4.60 -4.56
-6$\times$100 Branching Data -4.60 -4.60 -4.39 -4.02 -4.54 -4.58 -4.59
-6$\times$100 Coalescent Data -4.55 -4.54 -4.25 -2.88 -4.56 -4.58 -4.55
-6$\times$100 Gaussian Data -4.60 -4.60 -4.39 -4.03 -4.52 -4.57 -4.59
-28$\times$10 Branching Data -4.60 -4.60 -4.15 -1.89 -4.60 -4.60 -4.56
-28$\times$10 Coalescent Data -4.53 -3.76 -3.96 -2.11 -4.52 -4.57 -3.87
-28$\times$10 Gaussian Data -4.58 -4.58 -4.07 -2.40 -4.60 -4.60 -4.47
-28$\times$100 Branching Data -4.59 -4.57 -4.04 -2.89 -4.59 -4.59 -4.59
-28$\times$100 Coalescent Data -4.51 -4.54 -3.91 -2.33 -4.39 -4.58 -4.39
-28$\times$100 Gaussian Data -4.59 -4.53 -4.10 -2.78 -4.59 -4.59 -4.59
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 4.2.2 Wasserstein Projections
-The Wasserstein projection problems highlight the differences in convergent behavior between classical and tropical gradient methods, as they are $\min$-tropical location problems but not classically convex.
-The CDFs of the final log relative errors can be found in Figures 6 and 7.
-Figure 6: The CDF of the log relative error after 1000 steps for each gradient method across 50 initializations when minimizing the 2-Wasserstein projection objective.
-Each sub-figure corresponds to a dataset of different dimensionality (columns) and distribution (rows).
-Figure 7: The CDF of the log relative error after 1000 steps for each gradient method across 50 initializations when minimizing the $\infty$-Wasserstein projection objective.
-Each sub-figure corresponds to a dataset of different dimensionality (columns) and distribution (rows).
-In Figures 6 and 7 we see large steps at various intervals in the CDF plots, which is indicative of stable local minima.
+The Wasserstein projection problems highlight the differences in stability behavior between classical and tropical gradient methods, as they are $\min$-tropical location problems but not classically convex.
+The CDFs of the final log relative errors can be found in Figures 7 and 8.
+Figure 7: The CDF of the log relative error after 1000 steps for each gradient method with tuned learning rates across 50 random (Gaussian) initializations when minimizing the 2-Wasserstein projection objective.
+Each subfigure corresponds to a dataset with sample size $K=10,100$ of dimensionality $N=6,28$, sampled from a branching process, coalescent process, or a Gaussian distribution.
+Figure 8: The CDF of the log relative error after 1000 steps for each gradient method with tuned learning rates across 50 random (Gaussian) initializations when minimizing the $\infty$-Wasserstein projection objective.
+Each subfigure corresponds to a dataset with sample size $K=10,100$ of dimensionality $N=6,28$, sampled from a branching process, coalescent process, or a Gaussian distribution.
+In Figures 7 and 8 we see large steps at various intervals in the CDF plots, which is indicative of stable local minima.
 This is particularly prevalent in low dimensional space.
 Classical descent, Adam, and Adamax are particularly susceptible to this, while tropical descent and TrAdamax are more likely to pass these local minima.
-This difference is most pronounced when solving the $\infty$-Wasserstein problem in low dimensions—we are almost certain to reach the minimum with small error, while classical methods have a 50-80% chance of achieving the same error.
+This difference is most pronounced when solving the $\infty$-Wasserstein problem in low dimensions—we are almost certain to reach the minimum with relative error $\leq e^{-3}$, while classical methods have a 50-80% chance of achieving the same error.
 4.2.3 Linear Regression
-As the coalescent data lies on a tropical hyperplane (Ardila & Klivans,, 2006; Page et al.,, 2020), the minimum of the linear regression function for such data is exactly 0.
+As the coalescent data lies on a tropical hyperplane (ardila2006bergman; page2020tropical), the minimum of the linear regression function for such data is exactly 0.
 Figure 9 therefore shows the CDF of the log error rather than the CDF of the log relative error.
+Table 2: The mean log relative error of each gradient method after 1000 steps with tuned learning rates for branching process, coalescent process, and Gaussian datasets of size $K=10,100$ in $\mathbb{R}^{6}/\mathbb{R}\mathbf{1}$, $\mathbb{R}^{28}/\mathbb{R}\mathbf{1}$ when minimizing the linear regression objective function.
+The minimal mean errors for each dataset are in bold.
+Data CD TD SGD TSGD Adam Adamax TrAdamax
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 Table 2 shows the mean log errors of each gradient method and dataset, from which we see that the tropical methods outperform their classical counterparts for all but one dataset, often by a factor of two.
@@
 We also note that unlike the previous problems, TSGD is outperforming SGD when solving the linear regression problem, often to the extent that TSGD outperforms all classical gradient methods.
-Figure 8: The CDF of the log relative error after 1000 steps for each gradient method across 50 initializations when minimizing the linear regression objective.
-Each sub-figure corresponds to a dataset of different dimensionality (columns) and distribution (rows).
-Table 2: The mean log relative error of each gradient method after 1000 steps for each dataset when minimizing the linear regression objective function.
-The minimal mean errors for each dataset are in bold.
-Data CD TD SGD TSGD Adam Adamax TrAdamax
-6$\times$10 Branching Data -0.97 -5.79 -0.56 -2.16 -0.80 -0.84 -4.39
-6$\times$10 Coalescent Data -2.70 -5.87 -2.11 -5.47 -2.97 -2.66 -6.58
-6$\times$10 Gaussian Data -1.19 -5.48 -0.80 -1.35 -1.05 -1.10 -4.10
-6$\times$100 Branching Data -0.79 -5.49 -0.71 -1.43 -0.72 -0.72 -3.53
-6$\times$100 Coalescent Data -5.54 -5.76 -3.11 -5.57 -6.90 -7.01 -6.52
-6$\times$100 Gaussian Data -1.04 -5.50 -0.88 -1.70 -1.02 -1.02 -3.81
-28$\times$10 Branching Data -1.49 -5.51 -0.93 -4.49 -1.14 -1.14 -5.57
-28$\times$10 Coalescent Data -2.43 -6.63 -1.77 -5.55 -2.16 -2.29 -6.18
-28$\times$10 Gaussian Data -1.26 -5.16 -0.79 -4.29 -1.08 -1.08 -5.72
-28$\times$100 Branching Data -0.90 -4.37 -0.61 -1.42 -0.85 -0.77 -3.32
-28$\times$100 Coalescent Data -3.33 -5.84 -2.17 -5.35 -3.02 -3.20 -4.93
-28$\times$100 Gaussian Data -0.91 -4.35 -0.62 -1.07 -0.96 -0.90 -3.51
+Figure 9: The CDF of the log error after 1000 steps for each gradient method with tuned learning rates across 50 random (Gaussian) initializations when minimizing the linear regression objective.
+Each subfigure corresponds to a dataset with sample size $K=10,100$ of dimensionality $N=6,28$, sampled from a branching process, coalescent process, or a Gaussian distribution.
+4.3 Tropical Statistics for Applications
+We conclude our computational experiments by using tropical gradient methods to solve tropical statistical problems which appear in phylogenetic data analysis and game theory.
+In particular we look to use tropical centrality statistics to estimate the species tree of a multi-species coalescent model (liu2019modern) , and we use tropical linear regression to estimate the hidden preference factors in the Auction model introduced by akian2023tropical, which is a variant of the first-price sealed-bid auction model (krishna2009auction).
+4.3.1 Multi-species Coalescent Model
+A species tree refers to a rooted metric tree which represents the evolutionary relationships between a set of species.
+However, for a specific gene, we may not observe the same branching pattern, instead observing a gene-specific evolution which is referred to as a gene tree.
+This raises one of the key questions in phylogenetics; how do we estimate the species tree from a set of gene trees?
+The multi-species coalescent model is a probabilistic model for gene trees given a certain species tree (see liu2019modern for a detailed introduction), which is particularly valuable for formulating Bayesian estimators for species trees and producing simulated data with which to test species tree estimation methodologies.
+In this section, we consider the tropical Fermat-Weber and Fréchet means of gene trees as estimators for the underlying species tree.
+We fix a species tree with 8 leaves (see Figure 10), then use Dendropy (dendropy5) to sample 100 gene trees from a multi-species coalescent model with this underlying species tree and normalize this dataset to have average tropical norm 1.
+Figure 10: The underlying species tree used for the multi-species coalescent model generating our dataset of gene trees.
+The topology of this tree can be represented by the Newick string “(((((F,E),(H,G)),(D,C)),B),A)”.
+We then compute a Fermat-Weber point and Fréchet mean for this dataset using each of our gradient methods for 1000 steps with their trained learning rates (Appendix C).
+We perform these computations from 100 random initializations for each gradient method, then project the result to ultrametric space using the single-linkage algorithm, which is known to coincide with tropical projection (single-linkage; ardila2005subdominant), and record the topology of the resulting tree.
+The counts for each observed tree topology are presented in Tables 3 and 4.
+Newick Tree Adam Adamax CD SGD TD TrAdamax TSGD
+((B,A),(((F,E),(H,G)),(D,C))); 0 0 0 0 0 0 2
+(((((F,E),(H,G)),(D,C)),A),B); 0 0 0 3 0 0 0
+(((((F,E),(H,G)),(D,C)),B),A); 100 100 100 97 100 100 98
+Table 3: The counts of each Newick string (up to tree topology) returned by the tropical projection of a Fermat–Weber point computed via each gradient method.
+Each method is run for 1000 steps with trained learning rate, from 100 random initializations.
+Newick Tree Adam Adamax CD SGD TD TrAdamax TSGD
+((B,A),(((F,E),(H,G)),(D,C))); 0 0 0 0 0 0 5
+(((((F,E),(H,G)),(D,C)),A),B); 0 0 0 1 0 0 1
+(((((F,E),(H,G)),(D,C)),B),A); 100 100 100 99 100 100 94
+Table 4: The counts of each Newick string (up to tree topology) returned by the tropical projection of a Fréchet means computed via each gradient method.
+Each method is run for 1000 steps with trained learning rate, from 100 random initializations.
+We first note that in the vast majority of cases, our calculations recover the tree topology of the original species tree, suggesting that for this multi-species coalescent dataset, the tropical Fermat–Weber and Frécht mean points are sufficiently good estimators for the species tree.
+When comparing the performance of each gradient method, we notice that the stochastic methods are the only methods to not recover the species tree topology, with Fréchet means computed via TSGD producing the highest error rate.
+The relative performance of tropical Fermat–Weber and Fréchet mean points as estimators for multi-species coalescent species trees is currently an open topic of study (cox2025tropical).
+In exploring the consistency of these estimators it world be natural for future work to also consider the relative performance of tropical gradient methods for their computation.
+4.3.2 Auction Model with Hidden Preference Factors
+As a final computational demonstration, we use tropical gradient methods to solve the problem of tropical linear regression using the toy auction data example given by akian2023tropical.
+For $N$ firms and $K$ products, we let $(p_{ij})_{i\leq N,j\leq K}$ denote the price firm $i$ offers for product $j$.
+Assuming the offered prices are in a state of equilibrium, that is, no firm can raise their price and still win, differences in proposed price from one firm relative to another is indicative of a hidden preference factor $f_{i}$ in the market; that is, two firms’ offers for product $j$ are considered equal when $p_{i_{1}j}f_{i_{1}}=p_{i_{2}j}f_{i_{2}}$.
+By performing tropical linear regression on the data $V_{ij}=\exp(-p_{ij})$, we can detect these hidden preference factors.
+For further details of this model and its assumptions, see akian2023tropical.
+For our computations, we take $N=3$ and $K=6,100$, using the same toy datasets discussed in Section 7.2 of akian2023tropical, which have true hidden preference factors $f=(1,0.8,0.6)$.
+We then perform each gradient method for 1000 steps over 100 random initializations, using learning rates which have been tuned to these datasets (Appendix C).
+Our estimates for $f$ are then given by $\hat{f}\coloneqq\exp(\mathbf{t}_{1000}-\max(\mathbf{t}_{1000}))$, such that $\max_{i}\hat{f}=1$.
+K=6 K=100
+Adam (0.76 $\pm$ 0.33, 0.70 $\pm$ 0.32, 0.56 $\pm$ 0.21) (0.77 $\pm$ 0.33, 0.71 $\pm$ 0.32, 0.58 $\pm$ 0.21)
+Adamax (0.76 $\pm$ 0.33, 0.70 $\pm$ 0.32, 0.56 $\pm$ 0.21) (0.77 $\pm$ 0.33, 0.71 $\pm$ 0.32, 0.59 $\pm$ 0.21)
+CD (0.71 $\pm$ 0.43, 0.60 $\pm$ 0.42, 0.45 $\pm$ 0.30) (0.78 $\pm$ 0.33, 0.70 $\pm$ 0.32, 0.58 $\pm$ 0.20)
+SGD (0.73 $\pm$ 0.44, 0.57 $\pm$ 0.43, 0.41 $\pm$ 0.32) (0.77 $\pm$ 0.33, 0.70 $\pm$ 0.32, 0.58 $\pm$ 0.21)
+TD (1.00 $\pm$ 0.00, 0.81 $\pm$ 0.02, 0.60 $\pm$ 0.01) (1.00 $\pm$ 0.02, 0.75 $\pm$ 0.08, 0.62 $\pm$ 0.05)
+TrAdamax (1.00 $\pm$ 0.00, 0.81 $\pm$ 0.00, 0.60 $\pm$ 0.00) (0.93 $\pm$ 0.19, 0.79 $\pm$ 0.16, 0.65 $\pm$ 0.08)
+TSGD (1.00 $\pm$ 0.00, 0.82 $\pm$ 0.02, 0.60 $\pm$ 0.01) (1.00 $\pm$ 0.04, 0.79 $\pm$ 0.04, 0.61 $\pm$ 0.01)
+Table 5: The mean($\pm$ 1 std.) estimators $\hat{f}\coloneqq\exp(\mathbf{t}_{1000}-\max(\mathbf{t}_{1000}))$ for the hidden preference factors in the Auction model (akian2023tropical) when computed using gradient methods to solve the tropical linear regression problem.
+The true hidden preference factors are $f=(1,0.8,0.6)$.
+akian2023tropical uses the projective Krasnoselskii-Mann iteration to solve this tropical linear regression problem, which produced an estimate of $f^{\text{reg}}=(1,0.81,0.605)$ for the $K=6$ toy dataset.
+Table 5 shows the resulting mean estimates for these hidden preference factors when using different gradient methods to solve the tropical linear regression problem.
+We see that the accuracy of the tropical gradient methods vastly exceeds each of the classical gradient methods, and that these gradient methods produce particularly consistent results for the small sample problem.
+These computations are limited to a single Auction model, but show promising results.
+For future work, it would be of interest to explore how the precision of these methods changes with sample size and dimension, as well as a running time comparison with existing methods such as the projective Krasnoselskii-Mann iteration used by akian2023tropical or combinatorial simplex algorithms (allamigeon2014combinatorial).
 5 Conclusion and Discussion
 We have introduced the concept of tropical descent as the steepest descent method with respect to the tropical norm.
-We have demonstrated, both theoretically and experimentally, that it possesses well-behaved properties with respect to a wide class of optimization problems on the tropical projective torus, including several key statistical problems for the analysis of phylogenetic data.
+We have demonstrated, both theoretically and experimentally, that it exhibits desirable properties with respect to a wide class of optimization problems on the tropical projective torus, including several key statistical problems for the analysis of phylogenetic data.
 As well as proposing our own framework for tropical gradient methods, this work also provides a detailed review of the varying behavior of popular gradient methods when applied to tropical statistical optimization problems.
 Our theoretical results outline the relative strengths of tropical descent over classical descent for tropical location problems, but only provide global solvability for 1-sample problems.
-In practice, our experimental results go further; the optimization problems studied are location problems, rather than $\triangle$-star-quasi-convex problems, and yet tropical descent appears to consistently converge to a global minimum, not just the tropical convex hull of our dataset.
-It would therefore be a natural extension of this work to identify how much we can relax the $\triangle$-star-quasi-convexity condition while guaranteeing convergence.
+In practice, our experimental results go much further; the optimization problems studied are location problems, rather than $\triangle_{\min}$-star-quasi-convex problems, and yet tropical descent appears to consistently converge to a global minimum, not just the tropical convex hull of our dataset.
+It would therefore be a natural extension of this work to identify how much we can relax the $\triangle_{\min}$-star-quasi-convexity condition while guaranteeing global solvability.
 In this paper we have considered tree data with 4 to 8 leaves, and samples of size 10 to 100—at this scale, we have observed (T/)SGD to be a significant sacrifice in accuracy with little benefit in computational time.
 Furthermore, at this scale, the relative merits of SGD versus its tropical counterpart are inconsistent; TSGD out-performs SGD when solving the linear regression problem, but performs worse when finding centrality statistics.
-In a similar vein, initial investigations suggest that the classical stochastic algorithm for the computation of Fréchet means—Sturm’s algorithm (Sturm,, 2003)—performs poorly for the computation of tropical Fréchet means due to the inconsistent curvature of the tropical projective torus (Améndola & Monod,, 2023).
+In a similar vein, initial investigations suggest that the classical stochastic algorithm for the computation of Fréchet means—Sturm’s algorithm (sturm2003probability)—performs poorly for the computation of tropical Fréchet means due to the inconsistent curvature of the tropical projective torus (amendola2021invitation).
 However, for applications to larger datasets and more complex objective functions, stochastic gradient methods may become necessary.
 It would then be important to understand the theoretical differences between SGD and TSGD in terms of their convergence guarantees in the tropical setting and their comparative performance at different scales of data dimension.
-The recent identification of tropical rational functions and ReLU neural networks (Zhang et al.,, 2018) has inspired new avenues of research in tropical applications, into the tropical geometry of neural networks (Brandenburg et al.,, 2024; Yoshida et al.,, 2024).
-It is therefore natural to consider the possible extensions of this work to neural network training; can we identify local tropical quasi-convexity in the loss functions of ReLU neural networks, and how can we extend tropical steepest descent to the space of tropical rational functions?
-Finally, as shown by (Akian et al.,, 2023), the tropical linear regression problem is polynomial-time equivalent to mean payoff games—two-player perfect information games on a directed graph (Ehrenfeucht & Mycielski,, 1979; Gurvich et al.,, 1988).
+The recent identification of tropical rational functions and ReLU neural networks (zhang2018tropical) has inspired new avenues of research in tropical applications, such as the tropical geometry of neural networks (brandenburg2024real; yoshida2023tropNNs).
+It is therefore natural to consider the possible extensions of this work to neural network training; can we identify local tropical quasi-convexity in the loss functions of ReLU neural networks, and how can we extend tropical descent to the space of tropical rational functions?
+Finally, as shown by akian2023tropical, the tropical linear regression problem is polynomial-time equivalent to mean payoff games—two-player perfect information games on a directed graph (ehrenfeucht1979positional; gurvich1988cyclic).
 This family of problems is of particular interest in game theory as they are known to be NP $\cap$ co-NP.
@@
 A.M.
-is supported by the Engineering and Physical Sciences Research Council under grant reference [EP/Y028872/1] and a London Mathematical Society Emmy Noether Fellowship [EN-2223-01].
+is supported by EPSRC AI Hub on Mathematical Foundations of Intelligence: An “Erlangen Programme” for AI No.
+EP/Y028872/1.
 Declarations
@@
 Appendix A PCA and Logistic Regression
-Figure 9: The CDF of the log relative error after 1000 steps for each gradient method across 50 initializations when minimizing the PCA objective.
-Each sub-figure corresponds to a dataset of different dimensionality (columns) and distribution (rows).
-Figure 10: The CDF of the log relative error after 1000 steps for each gradient method across 50 initializations when minimizing the logistic regression objective.
-Each sub-figure corresponds to a dataset of different dimensionality (columns) and distribution (rows).
-As well as the location problems outlined in this paper, we performed the same numerical experiments with the tropical polytope PCA loss function (Yoshida et al.,, 2019) and tropical logistic regression loss function (Aliatimis et al.,, 2024).
+As well as the location problems outlined in this paper, we performed the same numerical experiments with the tropical polytope PCA loss function (yoshida2019tropicalPCA) and tropical logistic regression loss function (aliatimis2023logistic).
 These loss functions take multiple points in $\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ as argument, and we execute our gradient methods on each argument simultaneously.
-Figures 9 and 10 show the resulting CDF of our gradient methods for these problems.
+Figures 11 and 12 show the resulting CDF of our gradient methods for these problems.
 These problems do not lend themselves well to gradient methods, as the relative log error is rather large for all of our proposed methods.
 However, for both problems the tropical descent and TrAdamax methods are outperforming the classical methods quite consistently.
+Figure 11: The CDF of the log relative error after 1000 steps for each gradient method with tuned learning rates across 50 random (Gaussian) initializations when minimizing the PCA objective.
+Each subfigure corresponds to a dataset with sample size $K=10,100$ of dimensionality $N=6,28$, sampled from a branching process, coalescent process, or a Gaussian distribution.
+Figure 12: The CDF of the log relative error after 1000 steps for each gradient method with tuned learning rates across 50 random (Gaussian) initializations when minimizing the logistic regression objective.
+Each subfigure corresponds to a dataset with sample size $K=10,100$ of dimensionality $N=6,28$, sampled from a branching process, coalescent process, or a Gaussian distribution.
 Appendix B Methodology
-The same 12 datasets are used for each experiment in Section 4.
+The same 12 datasets are used for each experiment in Subsections 3.2 and 4.
 These datasets live in $\mathbb{R}^{6}/\mathbb{R}\mathbf{1}$ or $\mathbb{R}^{28}/\mathbb{R}\mathbf{1}$; these dimensions correspond to the spaces of trees with 4 and 8 leaves respectively.
-Branching process datasets are generated using the R package ape (Paradis & Schliep,, 2019) with exponentially distributed branch lengths—this produces data supported on the space of metric trees, i.e., the tropical Grassmannian.
+Branching process datasets are generated using the R package ape (paradis2019ape) with exponentially distributed branch lengths—this produces data supported on the space of metric trees, i.e., the tropical Grassmannian.
 Coalescent data has also been generated using ape and is supported on the space of ultrametric trees, which is a tropical linear space.
-Gaussian data is generated using NumPy (Harris et al.,, 2020), and is supported on the general tropical projective torus.
+Gaussian data is generated using NumPy (harris2020array), and is supported on the general tropical projective torus.
 In the computation of the Wasserstein projection objectives, we require a secondary dataset of different dimensionality; for each of the 12 datasets, we generate datasets of the same size and distribution, but in $\mathbb{R}^{3}/\mathbb{R}\mathbf{1}$ or $\mathbb{R}^{21}/\mathbb{R}\mathbf{1}$ which contain the spaces of trees with 3 and 7 leaves respectively.
-We use PyTorch (Paszke et al.,, 2019) for the computation and differentiation of each objective function.
-Max-tropical descent directions are used for tropical gradient methods; we therefore optimize the min-tropical location problems—Wasserstein projections and tropical linear regression—with respect to $-\mathbf{t}$.
+We use PyTorch (paszke2019pytorch) for the computation and differentiation of each objective function.
+Max-tropical descent directions are used for tropical gradient methods; we therefore optimize the min-tropical location problems—Wasserstein projections and tropical linear regression—with respect to $\mathbf{t}$.
 Appendix C Learning Rate Tuning
 The Adam algorithms—Adam, Adamax and TrAdamax—are run using the default $\beta=(0.9,0.999)$ parameters.
-We then tune the learning rates used for each method; Gradient Descent, Tropical Descent, SGD, Tropical SGD, Adam, Adamax, and TrAdamax.
-We perform an initial test of learning rates $\gamma$ on a logarithmic scale from 0.0001 to 30 over 1000 steps for each algorithm, with 5 random initializations for each dataset.
-From this, we narrowed our search to take 18 initializations at 0.5 intervals on the log scale in near optimal regions for each problem, gradient method, and data dimensionality.
-We then selected the learning rate with minimal mean log relative error in each case; Tables 3, 4, 5, 6, 7, 8 and 9 show the selected learning rates.
-Table 3: Learning rates for each gradient method when solving the Fermat–Weber problem.
-Data Dim Data Count CD TD SGD TSGD Adam Adamax TrAdamax
+We then tune the learning rates used for each method; CD, TD, SGD, TSGD, Adam, Adamax, and TrAdamax.
+We take 10 initializations at unit intervals on a log scale in the region $e^{-6},e^{4}]$ for each problem, gradient method, and data dimensionality.
+We then selected the learning rate with minimal mean log relative error in each case; Tables 6, 7, 8, 9, 10, 11 and 12 show the selected learning rates.
+For the Auction data in Subsection 4.3.2, we take 10 initializations at unit intervals in the region of $e^{-10},e^{5}]$ for each dataset; Table 13 shows the selected learning rates which minimize the mean log relative error.
+Table 6: Tuned learning rates for each gradient method when solving the Fermat–Weber problem.
+Data
+Dim
+Data
+Count
+CD TD SGD TSGD Adam Adamax TrAdamax
 6 10 0.135 0.135 0.135 0.135 0.00674 0.00674 0.00674
@@
 100 0.368 0.135 0.368 1 0.00674 0.00674 0.0183
-Table 4: Learning rates for each gradient method when solving the Fréchet mean problem.
-Data Dim Data Count CD TD SGD TSGD Adam Adamax TrAdamax
+Table 7: Tuned learning rates for each gradient method when solving the Fréchet mean problem.
+Data
+Dim
+Data
+Count
+CD TD SGD TSGD Adam Adamax TrAdamax
 6 10 0.135 0.135 0.135 0.135 0.00674 0.00674 0.00674
@@
 100 0.368 0.135 0.368 1 0.00674 0.00674 0.0183
-Table 5: Learning rates for each gradient method when solving the 2-Wasserstein projection problem.
-Data Dim Data Count CD TD SGD TSGD Adam Adamax TrAdamax
+Table 8: Tuned learning rates for each gradient method when solving the 2-Wasserstein projection problem.
+Data
+Dim
+Data
+Count
+CD TD SGD TSGD Adam Adamax TrAdamax
 6 10 2.72 1 1 0.135 0.135 0.368 0.368
@@
 100 2.72 2.72 0.368 1 0.0498 0.135 0.0498
-Table 6: Learning rates for each gradient method when solving the $\infty$-Wasserstein projection problem.
-Data Dim Data Count CD TD SGD TSGD Adam Adamax TrAdamax
+Table 9: Tuned learning rates for each gradient method when solving the $\infty$-Wasserstein projection problem.
+Data
+Dim
+Data
+Count
+CD TD SGD TSGD Adam Adamax TrAdamax
 6 10 0.0498 0.135 0.0498 0.368 0.00674 0.00674 0.0183
@@
 100 0.368 1 0.368 1 0.00674 0.0183 0.0498
-Table 7: Learning rates for each gradient method when solving the polytope PCA problem.
-Data Dim Data Count CD TD SGD TSGD Adam Adamax TrAdamax
+Table 10: Tuned learning rates for each gradient method when solving the polytope PCA problem.
+Data
+Dim
+Data
+Count
+CD TD SGD TSGD Adam Adamax TrAdamax
 6 10 2.72 2.72 1 1 0.135 0.135 0.368
@@
 100 2.72 7.39 0.368 1 0.0498 0.0498 0.0498
-Table 8: Learning rates for each gradient method when solving the logistic regression problem.
-Data Dim Data Count CD TD SGD TSGD Adam Adamax TrAdamax
+Table 11: Tuned learning rates for each gradient method when solving the logistic regression problem.
+Data
+Dim
+Data
+Count
+CD TD SGD TSGD Adam Adamax TrAdamax
 6 10 7.39 2.72 1 2.72 0.368 1 0.0498
@@
 100 1 0.368 1 7.39 0.00248 0.00248 0.00674
-Table 9: Learning rates for each gradient method when solving the linear regression problem.
-Data Dim Data Count CD TD SGD TSGD Adam Adamax TrAdamax
+Table 12: Tuned learning rates for each gradient method when solving the linear regression problem.
+Data
+Dim
+Data
+Count
+CD TD SGD TSGD Adam Adamax TrAdamax
 6 10 0.368 0.135 0.368 0.135 0.00248 0.0183 0.00674
@@
 100 1 0.368 1 7.39 0.0498 0.0498 0.0498
+Table 13: Tuned learning rates for each gradient method when solving the linear regression problem with Auction data.
+Data
+Dim
+Data
+Count
+CD TD SGD TSGD Adam Adamax TrAdamax
+3 6 2.72 0.368 7.39 0.0498 0.000912 0.000912 0.0183
+100 0.0183 2.72 0.0183 0.0498 0.000912 0.000912 0.00248
 Appendix D Run Times
-Experiments were implemented on a AMD EPYC 7742 node using a single core, 8 GB. Table 10 shows the average time taken per initialization for each gradient method and optimization problem.
+Experiments were implemented on a AMD EPYC 7742 node using a single core, 96 GB. Table 14 shows the average time taken per initialization for each gradient method and optimization problem.
 Table 14: Average run times for 1000 steps of each gradient method when minimizing each objective function.
@@
 Fréchet Mean 0.347 0.363 0.301 0.342 0.294 0.422 0.370
-2-Wasserstein Projection 2.739 2.134 2.392 2.209 2.341 2.246 2.272
-$\infty$-Wasserstein Projection 2.947 3.049 2.772 2.565 2.501 3.162 2.807
+2-Wasserstein
+Projection
+2.739 2.134 2.392 2.209 2.341 2.246 2.272
+$\infty$-Wasserstein
+Projection
+2.947 3.049 2.772 2.565 2.501 3.162 2.807
 Polytope PCA 1.117 1.081 1.014 1.052 1.037 1.174 1.029
@@
 E.1 Fréchet Means
-Here we include the results for experiments run for the computation of Fréchet means.
-Table 11: The mean log relative error of each gradient method after 1000 steps for each dataset when minimizing the Fréchet mean objective function.
+Here we include the results for experiments run for the computation of Fréchet means. (Table 15, Figure 13)
+Table 15: The mean log relative error of each gradient method after 1000 steps with tuned learning rates for branching process, coalescent process, and Gaussian datasets of size $K=10,100$ in $\mathbb{R}^{6}/\mathbb{R}\mathbf{1}$, $\mathbb{R}^{28}/\mathbb{R}\mathbf{1}$ when minimizing the Fréchet mean objective function.
 The minimal mean errors for each dataset are in bold.
 Data CD TD SGD TSGD Adam Adamax TrAdamax
-6$\times$10 Branching Data -4.59 -4.54 -3.98 -2.99 -4.58 -4.59 -4.53
-6$\times$10 Coalescent Data -4.58 -4.47 -4.07 -1.82 -4.57 -4.58 -4.45
-6$\times$10 Gaussian Data -4.59 -4.33 -4.09 -2.63 -4.58 -4.59 -4.35
-6$\times$100 Branching Data -4.59 -4.59 -4.35 -4.03 -4.59 -4.58 -4.59
-6$\times$100 Coalescent Data -4.57 -4.56 -4.36 -3.29 -4.55 -4.56 -4.57
-6$\times$100 Gaussian Data -4.59 -4.60 -4.39 -4.12 -4.59 -4.58 -4.59
-28$\times$10 Branching Data -4.58 -3.73 -3.78 -1.98 -4.55 -4.58 -3.83
-28$\times$10 Coalescent Data -4.56 -3.67 -3.98 -2.46 -4.53 -4.57 -3.70
-28$\times$10 Gaussian Data -4.55 -4.07 -3.91 -2.41 -4.55 -4.57 -4.04
-28$\times$100 Branching Data -4.59 -4.58 -4.03 -2.98 -4.58 -4.59 -4.57
-28$\times$100 Coalescent Data -4.54 -4.56 -4.08 -2.66 -4.47 -4.57 -4.46
-28$\times$100 Gaussian Data -4.59 -4.56 -4.08 -2.78 -4.59 -4.59 -4.58
-Figure 11: The CDF of the log relative error after 1000 steps for each gradient method across 50 initializations when minimizing the Fréchet mean objective.
-Each sub-figure corresponds to a dataset of different dimensionality (columns) and distribution (rows).
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Figure 13: The CDF of the log relative error after 1000 steps for each gradient method with tuned learning rates across 50 random (Gaussian) initializations when minimizing the Fréchet mean objective.
+Each subfigure corresponds to a dataset with sample size $K=10,100$ of dimensionality $N=6,28$, sampled from a branching process, coalescent process, or a Gaussian distribution.
 E.2 Wasserstein Projections
-From Tables 12 and 13, we see that when solving the 2-Wasserstein projection problem, tropical methods are optimal in all cases other than $28\times 10$ coalescent data.
+From Tables 16 and 17, we see that when solving the 2-Wasserstein projection problem, tropical methods are optimal in all cases other than $28\times 10$ coalescent data.
 For the $\infty$-Wasserstein projection problem in small dimensions, tropical descent outperforms other methods by some margin, while the comparative performances are less consistent for the $\infty$-Wasserstein projection problem in high dimensions.
-Table 12: The mean log relative error of each gradient method after 1000 steps for each dataset when minimizing the $2$-Wasserstein projection objective function.
+Table 16: The mean log relative error of each gradient method after 1000 steps with tuned learning rates for branching process, coalescent process, and Gaussian datasets of size $K=10,100$ in $\mathbb{R}^{6}/\mathbb{R}\mathbf{1}$, $\mathbb{R}^{28}/\mathbb{R}\mathbf{1}$ when minimizing the $2$-Wasserstein projection objective function.
 The minimal mean errors for each dataset are in bold.
 Data CD TD SGD TSGD Adam Adamax TrAdamax
-6$\times$10 Branching Data -3.02 -4.09 -2.16 -2.11 -2.95 -3.00 -3.89
-6$\times$10 Coalescent Data -3.71 -3.92 -2.44 -3.13 -3.78 -3.68 -4.12
-6$\times$10 Gaussian Data -3.59 -3.83 -2.87 -2.75 -3.61 -3.57 -3.89
-6$\times$100 Branching Data -3.44 -4.24 -2.91 -1.60 -3.79 -3.79 -4.02
-6$\times$100 Coalescent Data -4.26 -4.55 -3.27 -2.58 -4.39 -4.45 -4.50
-6$\times$100 Gaussian Data -3.69 -4.17 -3.33 -2.73 -3.78 -3.78 -4.18
-28$\times$10 Branching Data -3.39 -4.31 -3.16 -2.03 -3.53 -3.55 -4.32
-28$\times$10 Coalescent Data -4.11 -3.79 -3.39 -2.42 -4.17 -4.12 -3.80
-28$\times$10 Gaussian Data -3.43 -3.91 -3.21 -2.15 -3.50 -3.44 -3.83
-28$\times$100 Branching Data -3.61 -4.55 -3.37 -2.65 -3.85 -3.85 -4.51
-28$\times$100 Coalescent Data -4.11 -4.57 -3.89 -3.32 -4.21 -4.19 -4.56
-28$\times$100 Gaussian Data -3.66 -4.52 -3.41 -2.61 -3.82 -3.80 -4.49
-Table 13: The mean log relative error of each gradient method after 1000 steps for each dataset when minimizing the $\infty$-Wasserstein projection objective function.
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Table 17: The mean log relative error of each gradient method after 1000 steps with tuned learning rates for branching process, coalescent process, and Gaussian datasets of size $K=10,100$ in $\mathbb{R}^{6}/\mathbb{R}\mathbf{1}$, $\mathbb{R}^{28}/\mathbb{R}\mathbf{1}$ when minimizing the $\infty$-Wasserstein projection objective function.
 The minimal mean errors for each dataset are in bold.
 Data CD TD SGD TSGD Adam Adamax TrAdamax
-6$\times$10 Branching Data -1.63 -4.32 -0.57 -0.91 -1.68 -1.69 -4.32
-6$\times$10 Coalescent Data -4.00 -4.47 -0.99 -1.43 -4.03 -4.04 -4.46
-6$\times$10 Gaussian Data -3.77 -4.42 -1.76 -3.33 -3.76 -3.75 -4.40
-6$\times$100 Branching Data -3.35 -4.48 -1.94 -0.95 -3.34 -3.36 -4.47
-6$\times$100 Coalescent Data -3.85 -4.51 -2.06 -2.21 -3.84 -3.85 -4.51
-6$\times$100 Gaussian Data -3.63 -4.40 -1.61 -1.22 -3.70 -3.63 -4.36
-28$\times$10 Branching Data -3.68 -3.53 -2.22 -1.17 -4.11 -3.93 -3.17
-28$\times$10 Coalescent Data -4.22 -3.95 -2.76 -1.92 -4.42 -4.36 -3.94
-28$\times$10 Gaussian Data -4.02 -3.14 -2.13 -0.95 -4.43 -4.30 -2.85
-28$\times$100 Branching Data -3.46 -3.58 -1.37 -1.19 -3.68 -3.59 -3.39
-28$\times$100 Coalescent Data -4.41 -3.92 -2.09 -2.00 -4.54 -4.51 -3.81
-28$\times$100 Gaussian Data -2.98 -3.30 -1.64 -1.61 -3.07 -3.04 -2.80
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]